### v2

# Security Code Review by Large Language Models

Abstract.
Security code review, as a time-consuming and labour-intensive process, typically requires integration with automated security defect detection tools to ensure code security. Despite the emergence of numerous security analysis tools, those tools face challenges in terms of their poor generalization, high false positive rates, and coarse detection granularity. A recent development with Large Language Models (LLMs) has made them a promising candidate to support security code review. To this end, we conducted the first empirical study to understand the capabilities of LLMs in security code review, delving into the performance, quality problems, and influential factors of LLMs to detect security defects in code reviews. Specifically, we compared the performance of 6 LLMs under five different prompts with the state-of-the-art static analysis tools to detect and analyze security defects. For the best-performing LLM, we conducted a linguistic analysis to explore quality problems in its responses, as well as a regression analysis to investigate the factors influencing its performance. The results are that: (1) existing pre-trained LLMs have limited capability in detecting security defects during code review but significantly outperform the state-of-the-art static analysis tools. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) GPT-4 makes few factual errors but frequently generates unnecessary content or responses that are not compliant with the task requirements given in the prompts. (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic and written by developers with less involvement in the project.
Large Language Model, Code Review, Security Analysis
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Security defects can have severe consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004). Security code review is proven to be an effective process in the early stage of software development to prevent the introduction of security defects into code repositories, thereby mitigating the subsequent fixing and maintenance costs. Consequently, many popular projects and organizations have integrated security code review into their development pipelines (Paul et al., 2021). However, the resource-intensive nature of security code review, requiring significant human effort and time to review and revise code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024). Therefore, it would be helpful to develop effective automated tools to assist with detecting security defects during code review. Despite various security code analysis tools being proposed, they each encounter practical challenges such as high false positive rates, input range limitations, and weak scalability. Although data-driven approaches for security defect detection, especially those based on deep learning (DL), can somewhat handle these challenges, they inevitably lack robustness and tend to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024).
Pre-trained LLMs have demonstrated promising capabilities across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test improvement (Kang et al., 2023), and specification generation (Xie et al., 2023). Consequently, they are also considered a potential candidate for security code review. However, LLMs have not yet been employed in this specific field. Although many studies have examined the use of LLMs to detect security defects, most have been limited to quantitative evaluations of their performance (Bakhshandeh et al., 2023; Zhang et al., 2023; Zhou et al., 2024b), lacking in-depth analysis of specific problems encountered during the detection process and factors affecting the performance of LLMs in this task. These insights are crucial for enhancing the effectiveness of LLMs in security code review.
To bridge the knowledge gap, this study comprehensively explored the potential of LLMs in security code review from both qualitative and quantitative perspectives, including their capabilities, existing quality problems and influential factors. Specifically, We utilized a dataset of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator). First, based on the prompting strategies formulated by Zhang et al. (Zhang et al., 2023), we identified 5 prompt templates to enhance LLMs. Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers. After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses. We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM. Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review. (4) GPT-4 is more likely to identify security defects in code files that have fewer tokens, contain functional logic and were written by developers with less involvement in the project. In summary, this study makes the following contributions:
(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review; (2) We firstly measured the impact of the randomness of LLMs on the consistency of security code review; (3) We identified quality problems in the responses generated by the top-performing LLM to present the challenges it faces in security code review; (4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
2. Related Work
Several studies have explored the performance of LLMs in detecting security defects. The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type. Zhou et al. utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgement under various prompts (Zhou et al., 2024b). They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art security defect detection approaches, while GPT-4 consistently outperformed the state-of-the-art. Purba et al. compared the performance of LLMs with popular static analysis tools in binary detection. Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools (Purba et al., 2023). Unlike previous studies, our work provides a more granular evaluation of the capability of LLMs in file-level security code reviews, by including additional insights such as defect localization, description, and possible solutions.
Previous studies also investigated whether LLMs could provide more information in security defect detection. Bakhshandeh et al. provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers(Bakhshandeh et al., 2023). Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools. Yin et al. evaluated LLMs’ performance across multiple tasks related to security defects, including defect detection, localization, and description (Yin and Ni, 2024). They found that the existing state-of-the-art approaches such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023) generally outperformed LLMs. Specifically, in the tasks of defect localization and description, LLMs exhibited limited overall accuracy and varied performance across different CWE types. Most of the aforementioned studies utilized datasets that were synthesized or collected from tool documentation and CVE databases, while our study adopted a dataset derived from real-world code reviews in four industrial open-source projects, thus providing a reflection of the practicability of LLMs in security code review.
Employing specific prompting frameworks to enhance the performances of LLMs in identifying security defects has also been explored. Yang et al. introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection (Yang et al., 2024). Their framework yielded superior performance compared to state-of-the-art prompting techniques. Nong et al. proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information, to evaluate the performance of LLMs in identifying security defects (including their classification and resolution) (Nong et al., 2024). The main differences between the two studies above and our study are that: (1) Yang et al. and Nong et al. employed CoT through few-shot learning, whereas we used multi-turn dialogue for CoT. (2) Nong et al. utilized synthetic datasets and CVE datasets, whereas we utilized a manually curated security code review dataset obtained from open-source projects to guarantee data quality and reflect real-world applications. (3) Nong et al.’s work covered five CWE categories, whereas the dataset in our work included 15 security defect types, which are considered the most frequently identified types in code reviews (Yu et al., 2023).
In contrast to all the aforementioned studies, besides quantitatively evaluating LLM performance in security defect detection, we conducted more in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance. Thus, we are able to gain a more comprehensive understanding of the potential and challenges faced by LLMs in security code review.
3. Methodology
3.1. Research Questions
The goal of this work is to comprehensively assess the capability of LLMs in assisting with security code review, by automatically detecting security defects in the given code. To achieve this goal, we divided our study into three Research Questions (RQs), as shown below. The research procedure of each RQ is illustrated in Fig. 1.
RQ1: How do LLMs perform in detecting security defects identified by reviewers in code reviews?
RQ1.1: How do LLMs perform in security code review under the basic prompt?
RQ1.2: Can prompting with auxiliary information improve the performance of LLMs?
RQ1.3: Can the CoT prompting approach improve the performance of LLMs?
RQ1.4: What is the impact of the randomness of LLMs on the consistency of security code review?
Motivation: Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a). To address these limitations, RQ1 utilizes a manually-curated dataset to ensure data quality and comprehensively evaluates the capability of LLMs in reviewing a complete code file from real-world code reviews for security defects. Additionally, due to the inherent randomness of LLMs, for the same prompt, the responses of LLMs may exhibit considerable divergence. However, this aspect of LLMs has not yet been studied when applied to security tasks. Hence, we conducted a quantitative analysis of the consistency of LLM responses to explore the impact of LLM randomness on the task of security code review.
RQ2: What are the quality problems in LLM-generated responses when applied to detect security defects?
Motivation: Kabir et al. noted that the responses of ChatGPT to programming questions may be verbose, inconsistent, and contain conceptual or logical errors (Kabir et al., 2024), which negatively impact the quality of the answers. However, for security code review, the quality problems present in responses generated by LLMs have not been thoroughly analyzed in previous studies. Hence, RQ2 conducts an in-depth linguistic analysis of LLM responses for quality problems, and examines the types and distribution of these problems to bridge the knowledge gap and indicate potential areas for improvement.
RQ3: Which factors influence the performance of LLMs in security code review?
Motivation: Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations. However, there is a lack of evidence substantiating such claims. RQ3 employs a regression analysis to validate whether these factors have a significant impact on the performance of LLMs. This RQ helps to pinpoint areas of concerns in applying LLMs to security code review and highlight possible directions for improvement.
Figure 1. An overview of the research procedure for investigating the three RQs
3.2. Data Collection
We leveraged the code review dataset constructed by Yu et al. (Yu et al., 2023) and examined the capability of LLMs in security code review. The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types, from four projects: OpenStack Nova and Neutron, and Qt Base and Creator. These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018). For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs. Instead, we included only the code from the single file, where a review comment was made by the reviewer to identify a security defect. The 614 security-related review comments were reorganized as follows:
First, we filtered out comments that were published at the patchset level, as such data cannot be located in a specific file to construct a prompt. Then, for each code file that these comments correspond to, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item. Next, we utilized Gerrit API to retrieve the content of each code file and the commit messages of the patchsets they belong to. During this procedure, if we cannot obtain the content of the code files through API, we ignored those files. Lastly, since 97% files in this dataset are Python and C/C++ files, we excluded those written in languages other than these two (e.g., .json and .sh).
The final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
3.3. Research Procedure of RQ1
3.3.1. LLM Selection
To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December, 2023. Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through API. For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs. Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings. The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform. During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways. The two models provided by the ChatGPT platform performed better, thus were employed in our study.
Table 1. Versions and hyper-parameters of 6 LLMs studied
Model Version Context Length Parameter
temperature top_p
GPT-3.5 May 2024 version - -
GPT-4 May 2024 version - -
GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0
Gemini Pro gemini-1.0-pro 32k 0.9 1.0
Llama 2 7B llama-2-7b-chat 4k 1.0 0.9
Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
3.3.2. Prompt Design
Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023). To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al. (Zhang et al., 2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt. During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps. To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts. The detailed construction strategies of the five prompts are provided in (Anonymous, 2024).
  1. (1)
    Prompt 1 ($\textbf{P}_{b}$): the basic prompt. We directly instruct the LLM to review the provided code for security defects and output their description, code line numbers, and fixing solution. If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
  2. (2)
    Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information. $\textbf{P}_{r}$ is designed based on $\textbf{P}_{b}$ but includes the name of the source project as auxiliary information. The LLM is asked to act as a code reviewer for the project of the provided code in the prompt.
  3. (3)
    Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction. $\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects, without providing specific CWE details.
  4. (4)
    Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction. $\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)). It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries. The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$)..
  5. (5)
    Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$. The task of security code review is split into two steps. In $\textbf{P}_{cot-1}$, given that we only provide a single code file from a complete project in the prompt, we include the corresponding commit message of the file and instruct the LLM to generate code context according to the commit message. Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
3.3.3. Evaluation Metrics
Aiming to measure the performance of LLMs, we adopted the evaluation methodology proposed by Mahajan et al. (Mahajan et al., 2020) to rate the responses that LLMs generate with some adjustments, as detailed in Table 2. First, we excluded the ‘Unavailable’ category because LLMs do not have a case of no response. Next, the definitions of each category were adjusted to match the task of security code review. The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$). Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers, while a lower M-Score signifies less misleading results produced during security defect detection. Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness, and calculated the average of the three performance scores as the final result.
Table 2. Adjusted IHMU-category for LLM responses
Classification Definition
Instrumental (I) The response explicitly indicates the existence of the security defect identified by the reviewer and provides a fully accurate description.
Helpful
The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
Uncertain
The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects. Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
3.3.4. Baselines
We attempted different tools and learning techniques for detecting security defects to select appropriate baselines for our work. To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects. This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022). Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Cppcheck Solutions AB, 2024) for C/C++ and Bandit (ban, 2024) for Python. Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files. These two tools can directly analyze code files written in dynamically-typed languages like python, but require code files in statically-typed languages like C/C++ to be compiled first. Therefore, we could not apply them to our C/C++ dataset comprised of uncompiled and isolated C/C++ code files. Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets. In order to obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities. Other tools utilized their default rule sets for analysis.
3.3.5. Data Labelling
Prompts were constructed using five templates and then fed into six LLMs to collect their responses. To rate the responses, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2. To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses. Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately. Discrepancies in labelling results were discussed with the second author to reach a consensus. The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89. Then the first and third authors proceeded to label all the rest responses and discussed any ambiguous cases with the second author to finalize the categorization. Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset and found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1). Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
3.3.6. Consistency Calculation
To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments. Specifically, for a given code file under a specific prompt template, an LLM generated three responses. We then calculated the entropy for the four response categories in the three responses and determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment. A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
3.4. Research Procedure of RQ2
We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze quality problems present in these responses. We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
Based on the problem categories formulated by Kabir et al. (Kabir et al., 2024) and problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations. Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types. Notably, we excluded the Consistency theme defined in Kabir et al. (Kabir et al., 2024) from our study. The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1. Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors. In each round, 10 responses were randomly selected for the two authors to analyze their quality problems. The results were then compared and discussed, with partial adjustments made to the codebook. Since a single response could contain multiple problems, problem analysis is a multi-level and multi-label classification task where the labels are not mutually exclusive. Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement. Then the remaining data analysis was completed by the two authors and verified by the second author. Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis. For each problem type, we computed its average frequency in three repetitive experiments as the final occurrence rate of such problem in GPT-4 responses.
3.5. Research Procedure of RQ3
To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model. From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M. In terms of the helpfulness of the response, by definition, I is more important than H, and so forth. Accounting for the ordinal nature of the response variable and the randomness of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008). This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together. The specific construction process is outlined in the following subsections.
3.5.1. Attributes Construction
To comprehensively collect factors influencing the detection of security defects by LLMs, excluding the prompt and the model itself, we undertook the following steps:
First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work. Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntoshet al.’s work (McIntosh et al., 2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens. Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al.’s work  (Chen et al., 2023), transforming them into factors for our experiment. Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations. In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change. We used Python scripts complemented with manual analysis to query, compute, and label these factors. Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values. After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments and obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination. This comprised 405 responses from each experiment, along with their factor values, ready for model fitting.
Table 3. Factors that may influence the performance of LLM in security code review
Factor Source Definition Rationale
Token
Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file. Computed using the TikToken (tik, 2024) tokenizer. Larger code files may be more difficult for LLMs to analyze.
FileType
Created Type of the code file, i.e., Source or Auxiliary. Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp). The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
SecurityDefectType
Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al. (Yu et al., 2023). ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023). Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
Commit
Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change. Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
Complexity
Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976). Computed using the Lizard (liz, 2024) analyzer. Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
Community
Created Name of the community source of the code file, i.e., OpenStack or Qt. OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components. The LLMs’ performance may differ on data derived from these two communities.
AuthorExperience
Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to. Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
AnnotationRatio
Created The ratio of manual annotations in the code file. If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
AnnotationHasCode
Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code. Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
AnnotationisSecurityRelated
Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects). Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.
3.5.2. Regression Model Construction
Correlation and Redundancy Analysis: If explanatory variables are highly correlated with each other, it can cause interference in model construction and analysis. As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor. The correlation coefficients obtained are illustrated in Fig. 2. We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003). From Fig. 2, where each factor is denoted by an abbreviation, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors. However, two factors without high correlation can still be duplicates, which can distort the relationship between explanatory and response variables in model fitting. We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012). Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
Figure 2. The Phi_K correlation coefficients of explanatory variables.
Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman rank correlations (${\rho}^{2}$) between them. As recommended by Harrell  (Harrell, 2001), explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5. According to the computation results shown in Fig. 3, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘TK’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom. For categorical variables, they should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
Figure 3. Dotplot of the Spearman ${\rho}^{2}$ of explanatory variable and response variable in three repetitive experiments
Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015). To balance model complexity and fit, we tested models with and without allocating 3 degrees of freedom to the ‘TK’ factor. The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Anonymous, 2024).
4. Results and Analysis
4.1. RQ1 Performance of LLMs
Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts and compared the performance of LLMs with baseline tools in the C/C++ and Python dataset, respectively. If two out of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the other ones. Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied. To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances. As a result, two sets of scores are provided: 1) the response count and performance scores of each LLM-prompt combination on the complete dataset (Table 4), and 2) the performance scores of each LLM-prompt combination on the selected 57 instances (Table 5).
From Table 4, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance. However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best. However, Table 5 demonstrated that in the selected 57 cases, Llama 2 70B scored obviously lower than GPT-4. We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in Table 4. On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination. As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code review are still limited. Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools.
Table 4. Performance scores and entropy for each LLM-Prompt combination on the python and c/c++ dataset
Python C/C++
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 Resp. Count 123 123 123 123 122 GPT-3.5 Resp. Count 141 141 141 139 137
I-Score 1.08% 1.35% 0.27% 0.27% 1.37% I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
IH-Score 2.71% 3.80% 1.35% 0.27% 5.19% IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
M-Score 89.37% 84.28% 92.41% 97.29% 69.12% M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
Entropy 0.2572 0.3692 0.2070 0.0746 0.6842 Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
GPT-4 Resp. Count 175 175 175 174 174 GPT-4 Resp. Count 247 247 247 247 247
I-Score 4.76% 3.43% 3.24% 5.17*3.06% I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
IH-Score 9.34% 9.52% 8.38% 14.56*8.62% IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
M-Score 61.31% 60.95% 46.47% 44.06% 61.69% M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
Entropy 0.7426 0.7283 0.8600 0.9416 0.7450 Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
GPT-4 Turbo Resp. Count 201 201 201 201 201 GPT-4 Turbo Resp. Count 269 269 269 269 269
I-Score 0.66% 0.33% 0.66 0.83 0.17 I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
IH-Score 0.66 0.83 2.32 1.99 0.33 IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
M-Score 97.18% 96.35% 85.74% 92.87% 98.01% M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
Entropy 0.1757 0.0746 0.3370 0.6125 0.0404* Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
Gemini Pro Resp. Count 162 161 160 160 159 Gemini Pro Resp. Count 223 221 221 221 220
I-Score 0.00% 0.00% 0.63% 1.25% 0.21% I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
IH-Score 0.41% 0.00% 0.84% 3.54% 0.21% IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
M-Score 90.54% 97.09% 84.04% 55.21% 98.53% M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
Entropy 0.0777 0.0901 0.2424 0.1873 0.0548 Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
Llama 2 7B Resp. Count 62 62 61 61 50 Llama 2 7B Resp. Count 57 57 57 51 33
I-Score 0.00% 0.54% 1.62% 2.19% 1.33 I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
IH-Score 1.61% 1.61% 5.40% 8.74% 4.67% IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
M-Score 87.10% 83.33% 56.78% 40.98% 33.33*M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
Entropy 0.2625 0.3070 0.4766 0.4351 0.5521 Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
Llama 2 70B Resp. Count 62 62 62 61 47 Llama 2 70B Resp. Count 57 57 57 51 37
I-Score 0.54% 1.07% 2.15% 3.83% 1.42% I-Score 1.17% 1.17% 1.75% 6.54*1.80%
IH-Score 4.84% 4.30% 6.45% 11.48% 3.55% IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
M-Score 79.57% 84.41% 58.60% 43.17% 36.17% M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
Entropy 0.4470 0.3474 0.3514 0.3832 0.5762 Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
Baselines SAST SonarQube CodeQL Bandit Semgrep Baselines SAST CppCheck Semgrep
Resp. Count 258 258 258 258 Resp. Count 276 276
I-Score 0.00% 0.39% 0.00% 0.00% I-Score 0.00% 0.00%
IH-Score 0.00% 1.16% 1.16% 0.00% IH-Score 0.72% 0.36%
M-Score 91.47% 82.56% 78.29% 90.31% M-Score 98.19% 96.74%
  • The best values for each score are highlighted in bold and marked with *.
Table 5. Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
GPT-3.5 I-Score 3.51% 3.51% 0.58% 0.58% 2.34%
IH-Score 7.60% 8.77% 1.17% 0.58% 4.09%
M-Score 81.87% 79.53% 94.74% 95.32% 77.19%
GPT-4 I-Score 8.77% 8.77% 5.85% 12.87*9.94
IH-Score 20.46% 19.30% 14.04% 23.98*19.88
M-Score 43.86% 42.11% 45.61% 40.35% 41.52%
GPT-4 Turbo I-Score 0.58% 1.75% 2.34% 2.34% 0.00%
IH-Score 0.58% 3.51% 5.26% 4.09% 1.75%
M-Score 97.08% 91.23% 75.44% 88.30% 92.98%
Gemini Pro I-Score 0.58% 0.00% 0.00% 0.58% 0.00%
IH-Score 1.17% 0.00% 0.58% 2.92% 0.00%
M-Score 96.49% 100.00% 88.30% 76.61% 100.00%
Llama 2 7B I-Score 0.58% 1.17% 1.17% 0.58% 1.75%
IH-Score 1.75% 2.34% 4.09% 2.34% 7.60%
M-Score 87.13% 84.21% 60.82% 53.80% 25.73%
Llama 2 70B I-Score 2.34% 1.17% 2.92% 5.26% 4.09%
IH-Score 4.68% 2.34% 7.02% 17.54% 4.68%
M-Score 70.18% 80.12% 42.69% 20.47*44.44%
4.1.1. RQ1.1
When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%). Following closely behind is Llama 2 70B and GPT-3.5. GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit. Similar results are observed in Table 5. Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%. It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset. Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
4.1.2. RQ1.2
As shown in Table 4, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small. Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively. Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement. A possible reason is that the name of components or packages utilized may already imply the source project of the code file, thereby diminishing the impact of project information provided in $\textbf{P}_{r}$. Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific. We speculate that some LLMs may already understand the task described in the prompt quite well, thus introducing a persona could add unnecessary complexity of the prompt and lead to distractions.
We can see from Table 4 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets. The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset. However, all other LLMs achieved higher performance scores across both datasets. Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$. Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review. It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects, but it does not guarantee the accuracy of the detection results. Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4. Hence, how to leverage CWE information in the prompt is a question worthy of consideration.
4.1.3. RQ1.3
Table 4 illustrates that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B. However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs. As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets. Such instability may be attributed to differences in training data among different LLMs. Additionally, we first instructed the LLM to generate missing code context based on the corresponding commit message, where the consistency between this generated context and the actual code context cannot be guaranteed. Through analyzing specific cases, we discovered that some LLMs automatically fixed security defects while generating code context, leading to confusing and erroneous detection results in the second step of detecting security defects in the provided code.
4.1.4. RQ1.4
According to the mean entropies of various LLM-prompt combinations in Table 4, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency. Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404. On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195. Given the poor security defect detection capabilities of these two LLM-prompt combinations, their high consistency may be due to their tendency to not detect any security defects in the code. GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review, but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively. These results suggest that LLMs with stronger security defect detection capabilities may also exhibit greater randomness.
4.2. RQ2 Quality Problems in Responses
The average distribution of 12 problem types across 4 themes in responses to three repetitive experiments is illustrated in Fig. 4, where the same color represents types within the same theme, with darker shades indicating the themes.
Figure 4. Average proportion of each problem type present in responses generated by GPT-4
From the perspective of the four themes, a vast majority of responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below, the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
Example 1 - Incorrect concept
GPT-4: ...**CWE-682 (Incorrect Calculation):**
**Issue:** The method _inventory_has_changed uses a comparison logic that might not be entirely accurate in all scenarios...
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to the situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing overly general solutions. An example of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
Example 2 - Vague statement
GPT-4: ...
**CWE-703(Improper Check or Handling of Exceptional Conditions)**
**General Concern**: Throughout the code, there are instances where error conditions might not be handled comprehensively...
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appear in the responses generated by GPT-4. Such information includes the reasons for the absence of security defect in the current code, as well as general suggestions on security measures and review processes. As illustrated in the example below, the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
Example 3 - Unnecessary
GPT-4: ...**CWE-284 (Improper Access Control)**:
- There doesn’t seem to be any direct issue related to improper access control in the given code...
Compliance: Three types of non-compliance were identified with regard to whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and solutions of detected security defects requested in the prompt. From the example below, we can see that while the potential memory management issues caused by temporary files and directories are described and a solution is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
Example 4 - Incomplete answer
GPT-4: ### 3. CWE-664
**Potential Issue:** The creation of temporary files and directories (e.g., in function createQmlrcFile‘) ... can potentially lead to resource management issues.
**Solution:** Implement thorough checks and error handling...
4.3. RQ3 Factors Influencing LLMs
4.3.1. Model Stability
To assess the adequacy of our regression model’s fit, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation present in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
4.3.2. Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact. As shown in Table 6, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (¡.0001). Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$¡.0001) and ($\chi^{2}$=18.28, $p$¡.0001). Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively. However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold. Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
Table 6. Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(¿Chi)
TK 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
CMT Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
FT Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
ASR -1:0 0.97 2 9.95 0.0069**
1:0 2.69
AE 0.87 1 4.89 0.0271*
CT 1.01 1 3.06 0.0801
SDT Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
Crash:Race Condition 0.86
Deadlock:Race Condition 0.92
DoS:Race Condition 0.87
Encryption:Race Condition 1.18
Improper Access:Race Condition 0.68
Integer Overflow:Race Condition 0.86
Resource Leak:Race Condition 1.41
Use After Free:Race Condition 0.37
Extra:Race Condition 1.43
CPT 1.03 1 1.74 0.1868
AR 0.56 1 1.55 0.2126
AHC 1.11 1 1.09 0.2973
  • Statistical significance of explanatory power according to Wald $\chi^{2}$ test:
    ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$
4.3.3. Relationship between Factors and Responses
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 6). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files that contain functional logic, have fewer tokens, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
5. Implications
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (Steenhoek et al., 2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing the accuracy of detection. However, in real-world security reviews, although with the experience and domain knowledge, the reviewer is not guaranteed to consider all potential security defect types in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to effectively focus on the target defect, thereby yielding adverse effects.
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, it appears that annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations there are in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend developers write code comments in a standardized way to clarify the purpose of each comment, thus better support LLMs in code review tasks.
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amount of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs to overlook certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and LLama 2. Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
6. Threats to Validity
Internal Validity: One threat could be data leakage. The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs. Since the training data time of GPT-3.5 and GPT-4 is not clearly disclosed, we cannot avoid this overlap by filtering the data for a specific time period. Moreover, according to the findings by Cao et al. (Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cut-off date. Another threat is the design of prompts. To evoke the best model performance, we design our prompt according to the best practices. We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness. To mitigate this threat, we conducted three repetitive experiments for RQ1. For RQ2, we calculated the average frequency of each problem across these three experiments. In RQ3, we combined the responses from the three experiments to fit the model.
External validity: In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale). In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies. We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
Construct validity: With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure. These models are also not made publically available via API, thereby hindering the reproducibility of this study. Therefore, we explicitly indicate the versions of these two models used in the experiments, and provided all the responses generated by these models (Anonymous, 2024). In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements. Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
7. Conclusions
In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems present in responses generated by this LLM and performed regression analysis to explore factors influencing its performance. Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to the state-of-the-art static analysis tools. However, their application remains limited. (3) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) During code review, GPT-4 is more effective in identifying security defects in code that has fewer tokens, containing functional logic and written by developers with less involvement in the project. (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.

### v3

@@
-Security Code Review by Large Language Models
+An Insight into Security Code Review with LLMs: Capabilities, Obstacles and Influential Factors
 Abstract.
-Security code review, as a time-consuming and labour-intensive process, typically requires integration with automated security defect detection tools to ensure code security.
-Despite the emergence of numerous security analysis tools, those tools face challenges in terms of their poor generalization, high false positive rates, and coarse detection granularity.
-A recent development with Large Language Models (LLMs) has made them a promising candidate to support security code review.
-To this end, we conducted the first empirical study to understand the capabilities of LLMs in security code review, delving into the performance, quality problems, and influential factors of LLMs to detect security defects in code reviews.
-Specifically, we compared the performance of 6 LLMs under five different prompts with the state-of-the-art static analysis tools to detect and analyze security defects.
-For the best-performing LLM, we conducted a linguistic analysis to explore quality problems in its responses, as well as a regression analysis to investigate the factors influencing its performance.
-The results are that: (1) existing pre-trained LLMs have limited capability in detecting security defects during code review but significantly outperform the state-of-the-art static analysis tools.
+Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools.
+However, existing security analysis tools struggle with poor generalization, high false positive rates, and coarse detection granularity.
+Large Language Models (LLMs) have been considered promising candidates for addressing those challenges.
+In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review.
+Specifically, we evaluated the performance of six LLMs under five different prompts and compared them with state-of-the-art static analysis tools.
+We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance.
+Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review butå significantly outperform the state-of-the-art static analysis tools.
 (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference.
-(3) GPT-4 makes few factual errors but frequently generates unnecessary content or responses that are not compliant with the task requirements given in the prompts.
-(4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic and written by developers with less involvement in the project.
+(3) GPT-4 frequently generates responses that are verbose or not compliant with the task requirements given in the prompts.
+(4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
 Large Language Model, Code Review, Security Analysis
-copyright: acmcopyrightjournalyear: 2024doi: 10.1145/xxxxxxx.xxxxxxxisbn: 978-1-4503-XXXX-X/18/06
+copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software post-development issues
 1. Introduction
-Security defects can have severe consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004).
-Security code review is proven to be an effective process in the early stage of software development to prevent the introduction of security defects into code repositories, thereby mitigating the subsequent fixing and maintenance costs.
-Consequently, many popular projects and organizations have integrated security code review into their development pipelines (Paul et al., 2021).
-However, the resource-intensive nature of security code review, requiring significant human effort and time to review and revise code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024).
-Therefore, it would be helpful to develop effective automated tools to assist with detecting security defects during code review.
-Despite various security code analysis tools being proposed, they each encounter practical challenges such as high false positive rates, input range limitations, and weak scalability.
-Although data-driven approaches for security defect detection, especially those based on deep learning (DL), can somewhat handle these challenges, they inevitably lack robustness and tend to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024).
-Pre-trained LLMs have demonstrated promising capabilities across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test improvement (Kang et al., 2023), and specification generation (Xie et al., 2023).
-Consequently, they are also considered a potential candidate for security code review.
-However, LLMs have not yet been employed in this specific field.
-Although many studies have examined the use of LLMs to detect security defects, most have been limited to quantitative evaluations of their performance (Bakhshandeh et al., 2023; Zhang et al., 2023; Zhou et al., 2024b), lacking in-depth analysis of specific problems encountered during the detection process and factors affecting the performance of LLMs in this task.
-These insights are crucial for enhancing the effectiveness of LLMs in security code review.
-To bridge the knowledge gap, this study comprehensively explored the potential of LLMs in security code review from both qualitative and quantitative perspectives, including their capabilities, existing quality problems and influential factors.
-Specifically, We utilized a dataset of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator).
-First, based on the prompting strategies formulated by Zhang et al.
-(Zhang et al., 2023), we identified 5 prompt templates to enhance LLMs.
+Security defects in software programs can have severe consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004).
+The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002).
+Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (GitLab, 2022).
+Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality.
+During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations.
+Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (Weir et al., 2017), from being merged into the source code repository.
+In this study, we focus on security analysis in the context of code review, referring to it as security code review (Edmundson et al., 2013).
+Security code review is increasingly integrated into the development pipelines by project teams (Paul et al., 2021).
+Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024).
+Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021).
+Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013) and weak scalability (Nong et al., 2021).
+Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021).
+However, due to the lack of sizable and realistic training datasets (Nong et al., 2022), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024a).
+Nong et al.
+found that even using the training dataset augmented by the latest vulnerability injection technique (Nong et al., 2024b), the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively.
+Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
+Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023).
+Such models show remarkable capability to understand and generate explanatory natural language responses, thus considered a potential candidate for security code review.
+We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way.
+Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently.
+However, the application of LLMs in the specific field of security code review remains largely unexplored.
+Many studies have focused on leveraging LLMs for vulnerability detection.
+Zhou et al.
+evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts (Zhou et al., 2024b), while Purba et al.
+compared LLMs with static analysis tools in detecting software vulnerabilities (Purba et al., 2023).
+These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks, which was to judge whether there were security defects in the code.
+They did not require LLMs to provide more detailed information such as the location and type of vulnerability.
+Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al.
+(Yang et al., 2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al.
+(Nong et al., 2024a).
+However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
+Motivated by the above-mentioned limitations, our aim is to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories.
+Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description and suggested fix.
+The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator).
+We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review.
+First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al.
+(Zhang et al., 2023)).
 Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers.
@@
 (3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review.
-(4) GPT-4 is more likely to identify security defects in code files that have fewer tokens, contain functional logic and were written by developers with less involvement in the project.
+(4) GPT-4 is more likely to identify security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project.
 In summary, this study makes the following contributions:
-(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review; (2) We firstly measured the impact of the randomness of LLMs on the consistency of security code review; (3) We identified quality problems in the responses generated by the top-performing LLM to present the challenges it faces in security code review; (4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
+  • We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review.
+  • We measured the impact of the randomness of LLMs on the consistency of security code review.
+  • We identified quality problems in the responses generated by the top-performing LLM to present its challenges in security code review.
+  • We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
+Paper Organization.
+Section 2 surveys the related work.
+Section 3 describes the methodology employed in this study.
+Section 4 presents the results, followed by a discussion of the implications of our study results in Section 5.
+Section 6 clarifies the threats to validity.
+Finally, Section 7 concludes this work and outlines future directions.
+The replication package of this study has been provided online (Yu et al., 2024).
 2. Related Work
+Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, in this section, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or repair.
+Charoenwet et al.
+conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of vulnerability-contributing commits (VCCs) (Charoenwet et al., [n.
+d.]).
+By combining several tools, warnings could be produced for 78% of VCCs.
+However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review.
+Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
 Several studies have explored the performance of LLMs in detecting security defects.
@@
 Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools (Purba et al., 2023).
-Unlike previous studies, our work provides a more granular evaluation of the capability of LLMs in file-level security code reviews, by including additional insights such as defect localization, description, and possible solutions.
+Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description and suggestions for defect remediation.
 Previous studies also investigated whether LLMs could provide more information in security defect detection.
@@
 Yin et al.
-evaluated LLMs’ performance across multiple tasks related to security defects, including defect detection, localization, and description (Yin and Ni, 2024).
-They found that the existing state-of-the-art approaches such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023) generally outperformed LLMs.
-Specifically, in the tasks of defect localization and description, LLMs exhibited limited overall accuracy and varied performance across different CWE types.
-Most of the aforementioned studies utilized datasets that were synthesized or collected from tool documentation and CVE databases, while our study adopted a dataset derived from real-world code reviews in four industrial open-source projects, thus providing a reflection of the practicability of LLMs in security code review.
+evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description (Yin and Ni, 2024), found that the existing state-of-the-art approaches such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023) generally outperformed LLMs.
+Specifically, in defect localization and description tasks, LLMs exhibited limited overall accuracy and varied performance across different CWE types.
+These studies have all been limited to quantitatively evaluating LLM performance.
+In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to gain a comprehensive understanding of the potential and challenges LLMs face in security code review.
 Employing specific prompting frameworks to enhance the performances of LLMs in identifying security defects has also been explored.
@@
 Nong et al.
-proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information, to evaluate the performance of LLMs in identifying security defects (including their classification and resolution) (Nong et al., 2024).
-The main differences between the two studies above and our study are that: (1) Yang et al.
-and Nong et al.
-employed CoT through few-shot learning, whereas we used multi-turn dialogue for CoT.
-(2) Nong et al.
-utilized synthetic datasets and CVE datasets, whereas we utilized a manually curated security code review dataset obtained from open-source projects to guarantee data quality and reflect real-world applications.
-(3) Nong et al.’s work covered five CWE categories, whereas the dataset in our work included 15 security defect types, which are considered the most frequently identified types in code reviews (Yu et al., 2023).
-In contrast to all the aforementioned studies, besides quantitatively evaluating LLM performance in security defect detection, we conducted more in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance.
-Thus, we are able to gain a more comprehensive understanding of the potential and challenges faced by LLMs in security code review.
+proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution) (Nong et al., 2024a).
+The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours.
+These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements.
+Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
+It is worth noting that the aforementioned studies focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs.
+Inspired by these related works, we aim to leverage the potential of LLMs in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.
 3. Methodology
 3.1. Research Questions
-The goal of this work is to comprehensively assess the capability of LLMs in assisting with security code review, by automatically detecting security defects in the given code.
+The goal of this work is to comprehensively assess the capability of LLMs to assist with security code review by automatically detecting security defects in the given code.
 To achieve this goal, we divided our study into three Research Questions (RQs), as shown below.
-The research procedure of each RQ is illustrated in Fig.
+The research procedure for each RQ is illustrated in Fig.
 1.
-RQ1: How do LLMs perform in detecting security defects identified by reviewers in code reviews?
-RQ1.1: How do LLMs perform in security code review under the basic prompt?
+RQ1: How effectively do LLMs detect security defects during code review?
+RQ1.1: How effectively do LLMs perform security code reviews under the basic prompt?
 RQ1.2: Can prompting with auxiliary information improve the performance of LLMs?
-RQ1.3: Can the CoT prompting approach improve the performance of LLMs?
+RQ1.3: Can the Chain-of-Thought (CoT) prompting approach improve the performance of LLMs?
 RQ1.4: What is the impact of the randomness of LLMs on the consistency of security code review?
-Motivation: Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a).
-To address these limitations, RQ1 utilizes a manually-curated dataset to ensure data quality and comprehensively evaluates the capability of LLMs in reviewing a complete code file from real-world code reviews for security defects.
-Additionally, due to the inherent randomness of LLMs, for the same prompt, the responses of LLMs may exhibit considerable divergence.
+Motivation.
+Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a).
+To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality.
+It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects.
+Additionally, due to the inherent randomness of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt.
 However, this aspect of LLMs has not yet been studied when applied to security tasks.
 Hence, we conducted a quantitative analysis of the consistency of LLM responses to explore the impact of LLM randomness on the task of security code review.
-RQ2: What are the quality problems in LLM-generated responses when applied to detect security defects?
-Motivation: Kabir et al.
-noted that the responses of ChatGPT to programming questions may be verbose, inconsistent, and contain conceptual or logical errors (Kabir et al., 2024), which negatively impact the quality of the answers.
+RQ2: What are the quality problems in LLM-generated responses during security code review?
+Motivation.
+Kabir et al.
+noted that the responses of ChatGPT to programming questions may be verbose, inconsistent, and contain conceptual or logical errors (Kabir et al., 2024), negatively impacting the quality of the answers.
 However, for security code review, the quality problems present in responses generated by LLMs have not been thoroughly analyzed in previous studies.
-Hence, RQ2 conducts an in-depth linguistic analysis of LLM responses for quality problems, and examines the types and distribution of these problems to bridge the knowledge gap and indicate potential areas for improvement.
+Hence, RQ2 conducts an in-depth linguistic analysis of LLM responses for quality problems and examines the types and distribution of these problems to bridge the knowledge gap and indicate potential areas for improvement.
 RQ3: Which factors influence the performance of LLMs in security code review?
-Motivation: Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations.
+Motivation.
+Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations.
 However, there is a lack of evidence substantiating such claims.
-RQ3 employs a regression analysis to validate whether these factors have a significant impact on the performance of LLMs.
-This RQ helps to pinpoint areas of concerns in applying LLMs to security code review and highlight possible directions for improvement.
+RQ3 employs a regression analysis to validate whether these factors significantly impact the performance of LLMs.
+This RQ helps pinpoint areas of concern in applying LLMs to security code review and highlights possible improvement directions.
 Figure 1. An overview of the research procedure for investigating the three RQs
@@
 (Yu et al., 2023) and examined the capability of LLMs in security code review.
-The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types, from four projects: OpenStack Nova and Neutron, and Qt Base and Creator.
+The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (primarily written in Python), and Qt Base and Creator (primarily written in C++).
 These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018).
 For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs.
-Instead, we included only the code from the single file, where a review comment was made by the reviewer to identify a security defect.
+Instead, we included only the code from the single file, where the reviewer commented to identify a security defect.
 The 614 security-related review comments were reorganized as follows:
-First, we filtered out comments that were published at the patchset level, as such data cannot be located in a specific file to construct a prompt.
-Then, for each code file that these comments correspond to, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item.
-Next, we utilized Gerrit API to retrieve the content of each code file and the commit messages of the patchsets they belong to.
-During this procedure, if we cannot obtain the content of the code files through API, we ignored those files.
-Lastly, since 97% files in this dataset are Python and C/C++ files, we excluded those written in languages other than these two (e.g., .json and .sh).
-The final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
+First, we filtered out comments published at the patchset level, as such data cannot be located in a specific file to construct a prompt.
+Then, for each code file to which these comments correspond, we consolidated all security-related comments in the file and the corresponding security defect records into a single data item.
+Next, we utilized the Gerrit API to retrieve the content of each code file and the commit messages of the patchsets to which they belong.
+During this procedure, if we cannot obtain the content of the code files through the API, we ignore those files.
+Lastly, since 97% files in this dataset are Python, C and C++ files, we decided to remove any other files (all are configuration and scripting files) as those only formed $\approx 3\%$ of the files and thus had a negligible impact on our results.
+Our final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
 3.3. Research Procedure of RQ1
-3.3.1. LLM Selection
-To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December, 2023.
-Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through API.
-For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs.
-Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings.
-The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
-Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform.
-During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways.
-The two models provided by the ChatGPT platform performed better, thus were employed in our study.
-Table 1. Versions and hyper-parameters of 6 LLMs studied
+Table 1. Versions and hyper-parameters of 6 LLMs studied.
 Model Version Context Length Parameter
@@
 Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
+3.3.1. LLM Selection
+To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023.
+Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API.
+For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs.
+Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings.
+The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
+Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform.
+During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways.
+The two models provided by the ChatGPT platform performed better and thus were employed in our study.
 3.3.2. Prompt Design
@@
 To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts.
-The detailed construction strategies of the five prompts are provided in (Anonymous, 2024).
+Fig.
+2 illustrates the construction strategies for the five prompts.
   1. (1)
     Prompt 1 ($\textbf{P}_{b}$): the basic prompt.
-We directly instruct the LLM to review the provided code for security defects and output their description, code line numbers, and fixing solution.
+We instruct the LLM to review the provided code for security defects and output their description, code line numbers, and suggested fix.
 If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
@@
 $\textbf{P}_{r}$ is designed based on $\textbf{P}_{b}$ but includes the name of the source project as auxiliary information.
-The LLM is asked to act as a code reviewer for the project of the provided code in the prompt.
+The LLM is asked to act as a code reviewer for the project using the provided code in the prompt.
   3. (3)
     Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction.
-$\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects, without providing specific CWE details.
+$\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects without providing specific CWE details.
   4. (4)
@@
     Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$.
+Since we only included the code of a single file in the prompt instead of an entire patchset, we designed $\textbf{P}_{cot}$ to analyze the impact of missing contextual information on the performance of LLMs.
 The task of security code review is split into two steps.
@@
 Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
-3.3.3. Evaluation Metrics
-Aiming to measure the performance of LLMs, we adopted the evaluation methodology proposed by Mahajan et al.
-(Mahajan et al., 2020) to rate the responses that LLMs generate with some adjustments, as detailed in Table 2.
-First, we excluded the ‘Unavailable’ category because LLMs do not have a case of no response.
-Next, the definitions of each category were adjusted to match the task of security code review.
-The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$).
-Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers, while a lower M-Score signifies less misleading results produced during security defect detection.
-Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness, and calculated the average of the three performance scores as the final result.
-Table 2. Adjusted IHMU-category for LLM responses
+Figure 2. Construction templates for the five prompts
+Table 2. Adjusted IHMU-category for LLM responses.
 Classification Definition
 Instrumental (I) The response explicitly indicates the existence of the security defect identified by the reviewer and provides a fully accurate description.
-Helpful
-The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
+Helpful (H) The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
 Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
-Uncertain
-The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects.
+Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects.
 Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
+3.3.3. Evaluation Metrics
+During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements.
+We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al.
+(Mahajan et al., 2020) for measuring responses in Stack Overflow.
+We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al.
+because LLMs do not have a case of no response.
+Next, the definitions of each category were adjusted to match the task of security code review.
+Specifically, each response was categorized into one of the four rating categories detailed in Table 2.
+The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$).
+Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers.
+At the same time, a lower M-Score signifies less misleading results produced during security defect detection.
+Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness.
+We calculated the average of the three performance scores as the final result.
 3.3.4. Baselines
-We attempted different tools and learning techniques for detecting security defects to select appropriate baselines for our work.
+We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.
 To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects.
@@
 Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files.
-These two tools can directly analyze code files written in dynamically-typed languages like python, but require code files in statically-typed languages like C/C++ to be compiled first.
-Therefore, we could not apply them to our C/C++ dataset comprised of uncompiled and isolated C/C++ code files.
+These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first.
+Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files.
 Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets.
-In order to obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities.
+To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities.
 Other tools utilized their default rule sets for analysis.
@@
 Prompts were constructed using five templates and then fed into six LLMs to collect their responses.
-To rate the responses, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2.
+Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2.
 To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses.
@@
 The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89.
-Then the first and third authors proceeded to label all the rest responses and discussed any ambiguous cases with the second author to finalize the categorization.
-Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset and found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1).
+Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization.
+Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset.
+We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 4.1).
 Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
@@
 To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments.
-Specifically, for a given code file under a specific prompt template, an LLM generated three responses.
-We then calculated the entropy for the four response categories in the three responses and determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment.
+Specifically, an LLM generated three responses for a given code file under a specific prompt template.
+We then calculated the entropy for the four response categories in the three responses.
+We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment.
 A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
 3.4. Research Procedure of RQ2
-We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze quality problems present in these responses.
+We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses.
 We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
 Based on the problem categories formulated by Kabir et al.
-(Kabir et al., 2024) and problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations.
+(Kabir et al., 2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations.
 Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types.
-Notably, we excluded the Consistency theme defined in Kabir et al.
+In particular, we excluded the Consistency theme defined in Kabir et al.
 (Kabir et al., 2024) from our study.
@@
 Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors.
-In each round, 10 responses were randomly selected for the two authors to analyze their quality problems.
+In each round, ten responses were randomly selected for the two authors to analyze their quality problems.
 The results were then compared and discussed, with partial adjustments made to the codebook.
-Since a single response could contain multiple problems, problem analysis is a multi-level and multi-label classification task where the labels are not mutually exclusive.
+Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive.
 Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement.
@@
 Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis.
-For each problem type, we computed its average frequency in three repetitive experiments as the final occurrence rate of such problem in GPT-4 responses.
+For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
 3.5. Research Procedure of RQ3
@@
 This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together.
-The specific construction process is outlined in the following subsections.
+The specific construction process is described in the following subsections.
+Table 3. Factors that may influence the performance of LLM in security code review
+Factor Source* Definition Rationale
+Token Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file.
+Computed using the TikToken (tik, 2024) tokenizer.
+Larger code files may be more difficult for LLMs to analyze.
+FileType Created Type of the code file, i.e., Source or Auxiliary.
+Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp).
+The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
+SecurityDefectType Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al.
+(Yu et al., 2023).
+ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023).
+Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
+Commit Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change.
+Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
+Complexity Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976).
+Computed using the Lizard (liz, 2024) analyzer.
+Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
+Community Created Name of the community source of the code file, i.e., OpenStack or Qt.
+OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components.
+The LLMs’ performance may differ on data derived from these two communities.
+AuthorExperience Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to.
+Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
+AnnotationRatio Created The ratio of manual annotations in the code file.
+If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
+AnnotationHasCode Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code.
+Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
+AnnotationisSecurityRelated Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects).
+Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.
+    The sources of factors correspond to the descriptions in bold in Section 3.5.1.
 3.5.1. Attributes Construction
@@
 Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values.
-After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments and obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination.
-This comprised 405 responses from each experiment, along with their factor values, ready for model fitting.
-Table 3. Factors that may influence the performance of LLM in security code review
-Factor Source Definition Rationale
-Token
-Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file.
-Computed using the TikToken (tik, 2024) tokenizer.
-Larger code files may be more difficult for LLMs to analyze.
-FileType
-Created Type of the code file, i.e., Source or Auxiliary.
-Source indicates files that directly contain the logic of the application (i.e., .cpp, .c and .py) while Auxiliary files are utilized for declarations (i.e., .h and .hpp).
-The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
-SecurityDefectType
-Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al.
-(Yu et al., 2023).
-ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023).
-Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
-Commit
-Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change.
-Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
-Complexity
-Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976).
-Computed using the Lizard (liz, 2024) analyzer.
-Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
-Community
-Created Name of the community source of the code file, i.e., OpenStack or Qt.
-OpenStack and Qt differ significantly in terms of their structure, technology stack, project function, and main components.
-The LLMs’ performance may differ on data derived from these two communities.
-AuthorExperience
-Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to.
-Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
-AnnotationRatio
-Created The ratio of manual annotations in the code file.
-If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
-AnnotationHasCode
-Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code.
-Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
-AnnotationisSecurityRelated
-Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects).
-Keywords in the description related to target security defect in annotations can act as cues for LLMs, while the descriptions of non-target security defects often involve the historical fixing or security mechanisms, which may cause interference to LLMs.
+After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments.
+We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination.
+This comprised 405 responses from each experiment and their factor values ready for model fitting.
 3.5.2. Regression Model Construction
-Correlation and Redundancy Analysis: If explanatory variables are highly correlated with each other, it can cause interference in model construction and analysis.
+Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis.
 As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor.
 The correlation coefficients obtained are illustrated in Fig.
-2.
+3.
 We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003).
 From Fig.
-2, where each factor is denoted by an abbreviation, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors.
-However, two factors without high correlation can still be duplicates, which can distort the relationship between explanatory and response variables in model fitting.
+3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors.
+However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting.
 We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012).
 Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
-Figure 2. The Phi_K correlation coefficients of explanatory variables.
-Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman rank correlations (${\rho}^{2}$) between them.
-As recommended by Harrell  (Harrell, 2001), explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5.
+Figure 3. The Phi_K correlation coefficients of explanatory variables
+Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman’s rank correlations (${\rho}^{2}$) between them.
+As Harrell  (Harrell, 2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5.
 According to the computation results shown in Fig.
-3, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘TK’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom.
-For categorical variables, they should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
-Figure 3. Dotplot of the Spearman ${\rho}^{2}$ of explanatory variable and response variable in three repetitive experiments
+4, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom.
+Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
+Figure 4. Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
 Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015).
-To balance model complexity and fit, we tested models with and without allocating 3 degrees of freedom to the ‘TK’ factor.
+We tested models with and without allocating 3 degrees of freedom to the ‘Token’ factor to balance model complexity and fit.
 The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
-All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Anonymous, 2024).
-4. Results and Analysis
-4.1. RQ1 Performance of LLMs
-Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts and compared the performance of LLMs with baseline tools in the C/C++ and Python dataset, respectively.
-If two out of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the other ones.
+All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2024).
+4. Results
+4.1. Performance of LLMs (RQ1)
+Based on the evaluation metrics defined in Section 3.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts.
+We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively.
+If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the others.
 Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied.
 To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances.
-As a result, two sets of scores are provided: 1) the response count and performance scores of each LLM-prompt combination on the complete dataset (Table 4), and 2) the performance scores of each LLM-prompt combination on the selected 57 instances (Table 5).
-From Table 4, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance.
+As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 4 and Table 5), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 6).
+The best values for each score are highlighted in bold and marked with asterisks in the tables.
+From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance.
 However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best.
-However, Table 5 demonstrated that in the selected 57 cases, Llama 2 70B scored obviously lower than GPT-4.
-We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in Table 4.
+However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4.
+We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets.
 On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination.
-As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code review are still limited.
+As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited.
 Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools.
-Table 4. Performance scores and entropy for each LLM-Prompt combination on the python and c/c++ dataset
-Python C/C++
-Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
-GPT-3.5 Resp. Count 123 123 123 123 122 GPT-3.5 Resp. Count 141 141 141 139 137
-I-Score 1.08% 1.35% 0.27% 0.27% 1.37% I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
-IH-Score 2.71% 3.80% 1.35% 0.27% 5.19% IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
-M-Score 89.37% 84.28% 92.41% 97.29% 69.12% M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
-Entropy 0.2572 0.3692 0.2070 0.0746 0.6842 Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
-GPT-4 Resp. Count 175 175 175 174 174 GPT-4 Resp. Count 247 247 247 247 247
-I-Score 4.76% 3.43% 3.24% 5.17*3.06% I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
-IH-Score 9.34% 9.52% 8.38% 14.56*8.62% IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
-M-Score 61.31% 60.95% 46.47% 44.06% 61.69% M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
-Entropy 0.7426 0.7283 0.8600 0.9416 0.7450 Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
-GPT-4 Turbo Resp. Count 201 201 201 201 201 GPT-4 Turbo Resp. Count 269 269 269 269 269
-I-Score 0.66% 0.33% 0.66 0.83 0.17 I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
-IH-Score 0.66 0.83 2.32 1.99 0.33 IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
-M-Score 97.18% 96.35% 85.74% 92.87% 98.01% M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
-Entropy 0.1757 0.0746 0.3370 0.6125 0.0404* Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
-Gemini Pro Resp. Count 162 161 160 160 159 Gemini Pro Resp. Count 223 221 221 221 220
-I-Score 0.00% 0.00% 0.63% 1.25% 0.21% I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
-IH-Score 0.41% 0.00% 0.84% 3.54% 0.21% IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
-M-Score 90.54% 97.09% 84.04% 55.21% 98.53% M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
-Entropy 0.0777 0.0901 0.2424 0.1873 0.0548 Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
-Llama 2 7B Resp. Count 62 62 61 61 50 Llama 2 7B Resp. Count 57 57 57 51 33
-I-Score 0.00% 0.54% 1.62% 2.19% 1.33 I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
-IH-Score 1.61% 1.61% 5.40% 8.74% 4.67% IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
-M-Score 87.10% 83.33% 56.78% 40.98% 33.33*M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
-Entropy 0.2625 0.3070 0.4766 0.4351 0.5521 Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
-Llama 2 70B Resp. Count 62 62 62 61 47 Llama 2 70B Resp. Count 57 57 57 51 37
-I-Score 0.54% 1.07% 2.15% 3.83% 1.42% I-Score 1.17% 1.17% 1.75% 6.54*1.80%
-IH-Score 4.84% 4.30% 6.45% 11.48% 3.55% IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
-M-Score 79.57% 84.41% 58.60% 43.17% 36.17% M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
-Entropy 0.4470 0.3474 0.3514 0.3832 0.5762 Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
-Baselines SAST SonarQube CodeQL Bandit Semgrep Baselines SAST CppCheck Semgrep
-Resp. Count 258 258 258 258 Resp. Count 276 276
-I-Score 0.00% 0.39% 0.00% 0.00% I-Score 0.00% 0.00%
-IH-Score 0.00% 1.16% 1.16% 0.00% IH-Score 0.72% 0.36%
-M-Score 91.47% 82.56% 78.29% 90.31% M-Score 98.19% 96.74%
-  • The best values for each score are highlighted in bold and marked with *.
+This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions.
+As such features can only be accurately modelled using dynamic analysis, these security defects are challenging for static analysis tools to detect (Charoenwet et al., [n.
+d.]).
+Table 4. Performance scores and entropy for each LLM-Prompt combination on the Python dataset
+Python
+Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
+GPT-3.5 Resp. Count 123 123 123 123 122
+I-Score 1.08% 1.35% 0.27% 0.27% 1.37%
+IH-Score 2.71% 3.80% 1.35% 0.27% 5.19%
+M-Score 89.37% 84.28% 92.41% 97.29% 69.12%
+Entropy 0.2572 0.3692 0.2070 0.0746 0.6842
+GPT-4 Resp. Count 175 175 175 174 174
+I-Score 4.76% 3.43% 3.24% 5.17*3.06%
+IH-Score 9.34% 9.52% 8.38% 14.56*8.62%
+M-Score 61.31% 60.95% 46.47% 44.06% 61.69%
+Entropy 0.7426 0.7283 0.8600 0.9416 0.7450
+GPT-4 Turbo Resp. Count 201 201 201 201 201
+I-Score 0.66% 0.33% 0.66 0.83 0.17
+IH-Score 0.66 0.83 2.32 1.99 0.33
+M-Score 97.18% 96.35% 85.74% 92.87% 98.01%
+Entropy 0.1757 0.0746 0.3370 0.6125 0.0404*
+Gemini Pro Resp. Count 162 161 160 160 159
+I-Score 0.00% 0.00% 0.63% 1.25% 0.21%
+IH-Score 0.41% 0.00% 0.84% 3.54% 0.21%
+M-Score 90.54% 97.09% 84.04% 55.21% 98.53%
+Entropy 0.0777 0.0901 0.2424 0.1873 0.0548
+Llama 2 7B Resp. Count 62 62 61 61 50
+I-Score 0.00% 0.54% 1.62% 2.19% 1.33
+IH-Score 1.61% 1.61% 5.40% 8.74% 4.67%
+M-Score 87.10% 83.33% 56.78% 40.98% 33.33*%
+Entropy 0.2625 0.3070 0.4766 0.4351 0.5521
+Llama 2 70B Resp. Count 62 62 62 61 47
+I-Score 0.54% 1.07% 2.15% 3.83% 1.42%
+IH-Score 4.84% 4.30% 6.45% 11.48% 3.55%
+M-Score 79.57% 84.41% 58.60% 43.17% 36.17%
+Entropy 0.4470 0.3474 0.3514 0.3832 0.5762
+Baselines SAST SonarQube CodeQL Bandit Semgrep
+Resp. Count 258 258 258 258
+I-Score 0.00% 0.39% 0.00% 0.00%
+IH-Score 0.00% 1.16% 1.16% 0.00%
+M-Score 91.47% 82.56% 78.29% 90.31%
+Table 5. Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
+C/C++
+Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
+GPT-3.5 Resp. Count 141 141 141 139 137
+I-Score 1.89% 2.13% 0.00% 0.24% 0.73%
+IH-Score 5.67% 3.31% 0.47% 0.96% 1.46%
+M-Score 76.60% 86.05% 99.29% 95.44% 84.43%
+Entropy 0.4860 0.2439 0.0195* 0.1153 0.2931
+GPT-4 Resp. Count 247 247 247 247 247
+I-Score 4.45% 3.91% 3.91% 5.40% 5.26%
+IH-Score 7.96% 8.01% 9.31% 11.74% 11.47%
+M-Score 55.47% 64.24% 44.53% 39.00% 43.59%
+Entropy 0.6215 0.5329 0.6505 0.6955 0.7330
+GPT-4 Turbo Resp. Count 269 269 269 269 269
+I-Score 0.25% 0.12% 0.25% 0.50% 0.25%
+IH-Score 0.25% 0.37% 1.24% 1.61% 1.12%
+M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
+Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
+Gemini Pro Resp. Count 223 221 221 221 220
+I-Score 0.00% 0.15% 0.30% 1.66% 0.00%
+IH-Score 0.45% 0.15% 0.90% 3.77% 0.45%
+M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
+Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
+Llama 2 7B Resp. Count 57 57 57 51 33
+I-Score 0.58% 0.00% 0.00% 2.61% 1.01%
+IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
+M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
+Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
+Llama 2 70B Resp. Count 57 57 57 51 37
+I-Score 1.17% 1.17% 1.75% 6.54*1.80%
+IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
+M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
+Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
+Baselines SAST CppCheck Semgrep
+Resp. Count 276 276
+I-Score 0.00% 0.00%
+IH-Score 0.72% 0.36%
+M-Score 98.19% 96.74%
 Table 6. Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
@@
 4.1.1. RQ1.1
-When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%).
+When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4 and Table 5), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%).
 Following closely behind is Llama 2 70B and GPT-3.5.
@@
 4.1.2. RQ1.2
-As shown in Table 4, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small.
+As shown in Table 4 and Table 5, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small.
 Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively.
@@
 Furthermore, the improvement of emphasizing the task through adding a persona in the prompt is LLM-specific.
-We speculate that some LLMs may already understand the task described in the prompt quite well, thus introducing a persona could add unnecessary complexity of the prompt and lead to distractions.
-We can see from Table 4 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets.
+We speculate that some LLMs may already understand the task described in the prompt quite well, and thus, introducing a persona could add unnecessary complexity to the prompt and lead to distractions.
+We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets.
 The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset.
@@
 Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review.
-It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects, but it does not guarantee the accuracy of the detection results.
+It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects.
+Still, it does not guarantee the accuracy of the detection results.
 Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4.
-Hence, how to leverage CWE information in the prompt is a question worthy of consideration.
+Hence, how to leverage CWE information in the prompt is worthy of consideration.
 4.1.3. RQ1.3
-Table 4 illustrates that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B.
+Table 4 and Table 5 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets.
 However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs.
 As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets.
-Such instability may be attributed to differences in training data among different LLMs.
-Additionally, we first instructed the LLM to generate missing code context based on the corresponding commit message, where the consistency between this generated context and the actual code context cannot be guaranteed.
-Through analyzing specific cases, we discovered that some LLMs automatically fixed security defects while generating code context, leading to confusing and erroneous detection results in the second step of detecting security defects in the provided code.
+It can be seen that the impact of supplementing missing context by CoT prompting on the performance of LLMs is multifaceted and unstable.
+A possible explanation is that different LLMs vary in their ability to process and understand input.
+Moreover, we instructed the LLM to generate missing code context based on the corresponding commit message in $\textbf{P}_{cot}$.
+On the one hand, the generated context may contain the critical information needed to analyze the target security defect, which can help the LLM in security code review.
+On the other hand, the consistency between this generated context and the actual code context cannot be guaranteed.
+We observed that in some cases, the code context generated under Prompt $\textbf{P}_{cot-1}$ was inconsistent with the actual code context, resulting in the original security defect no longer being identified as a defect within the generated context.
+Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
 4.1.4. RQ1.4
-According to the mean entropies of various LLM-prompt combinations in Table 4, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency.
+According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency.
 Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404.
 On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195.
-Given the poor security defect detection capabilities of these two LLM-prompt combinations, their high consistency may be due to their tendency to not detect any security defects in the code.
-GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review, but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively.
-These results suggest that LLMs with stronger security defect detection capabilities may also exhibit greater randomness.
-4.2. RQ2 Quality Problems in Responses
-The average distribution of 12 problem types across 4 themes in responses to three repetitive experiments is illustrated in Fig.
-4, where the same color represents types within the same theme, with darker shades indicating the themes.
+Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code.
+GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively.
+These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater randomness.
+4.2. Quality Problems in Responses (RQ2)
+The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig.
+5, where the same color represents types within the same theme, with darker shades indicating the themes.
 Figure 5. Average proportion of each problem type present in responses generated by GPT-4
-From the perspective of the four themes, a vast majority of responses contain problems related to Conciseness, accounting for 90.18%.
+From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%.
 Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met.
@@
 Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type.
-As shown in the example below, the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
+As shown in the example below (Example 1), the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
 Example 1 - Incorrect concept
@@
 Among these, Vague statement is the most prevalent one.
-It refers to the situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing overly general solutions.
-An example of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
+It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing a overly general fix suggestion.
+An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
 Example 2 - Vague statement
@@
 Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses.
-Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appear in the responses generated by GPT-4.
-Such information includes the reasons for the absence of security defect in the current code, as well as general suggestions on security measures and review processes.
-As illustrated in the example below, the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason.
+Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4.
+Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes.
+As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason.
 However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
@@
 - There doesn’t seem to be any direct issue related to improper access control in the given code...
-Compliance: Three types of non-compliance were identified with regard to whether the responses satisfied the requirements in the prompts.
+Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts.
 Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks.
 Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found.
-Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and solutions of detected security defects requested in the prompt.
-From the example below, we can see that while the potential memory management issues caused by temporary files and directories are described and a solution is provided, no specific line numbers are given.
+Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt.
+From the example below (Example 4), we can see that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given.
 Hence, it is categorized as an Incomplete answer.
@@
 **Solution:** Implement thorough checks and error handling...
-4.3. RQ3 Factors Influencing LLMs
+While Kabir et al.
+(Kabir et al., 2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours.
+Specifically, Kabir et al.
+requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues.
+In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%.
+This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific areas of application.
+4.3. Factors Influencing LLMs (RQ3)
 4.3.1. Model Stability
-To assess the adequacy of our regression model’s fit, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991).
+To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991).
 This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022).
-A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation present in the dataset.
+A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset.
 Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model.
 Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
-4.3.2. Explanatory Power of Each Factor
-The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable.
-A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model.
-We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact.
-As shown in Table 6, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (¡.0001).
-Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$¡.0001) and ($\chi^{2}$=18.28, $p$¡.0001).
-Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively.
-However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold.
-Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom.
-As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
 Table 7. Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
-Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(¿Chi)
-TK 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
-CMT Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
-FT Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
-ASR -1:0 0.97 2 9.95 0.0069**
+Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(>Chi)
+Token 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
+Community Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
+FileType Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
+AnnotationisSecurityRelated -1:0 0.97 2 9.95 0.0069**
 1:0 2.69
-AE 0.87 1 4.89 0.0271*
-CT 1.01 1 3.06 0.0801
-SDT Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
+AuthorExperience 0.87 1 4.89 0.0271*
+Commit 1.01 1 3.06 0.0801
+SecurityDefectType Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
 Crash:Race Condition 0.86
@@
 Extra:Race Condition 1.43
-CPT 1.03 1 1.74 0.1868
-AR 0.56 1 1.55 0.2126
-AHC 1.11 1 1.09 0.2973
+Complexity 1.03 1 1.74 0.1868
+AnnotationRatio 0.56 1 1.55 0.2126
+AnnotationHasCode 1.11 1 1.09 0.2973
   • Statistical significance of explanatory power according to Wald $\chi^{2}$ test:
     ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$
-4.3.3. Relationship between Factors and Responses
+4.3.2. Explanatory Power of Each Factor
+The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable.
+A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model.
+We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact.
+As shown in Table 7, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (<.0001).
+Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$<.0001) and ($\chi^{2}$=18.28, $p$<.0001).
+Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively.
+However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold.
+Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom.
+As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
+4.3.3. Relationships between Factors and Responses
 The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016).
@@
 The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively.
-This suggests that the LLM is more likely to detect security defects in code files that contain functional logic, have fewer tokens, and were written by developers with less involvement in the project.
+This suggests that the LLM is more likely to detect security defects in code files that have fewer tokens, contain functional logic, and were written by developers with less involvement in the project.
 Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt.
-In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations the LLM is more likely to identify the target security defect.
+In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect.
 Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97.
 This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
-5. Implications
+5. Discussion
+A multi-layered review strategy is suggested to apply LLMs to security code review.
+As the results of RQ1 indicate, the existing general-purpose LLMs still fall significantly short of reaching the effectiveness of manual security code review.
+However, LLMs can serve as auxiliary tools to assist reviewers.
+The findings of RQ3 suggest that LLMs are more adept at identifying relatively simple security defects in shorter, logic-centric code written by developers with less project experience.
+We recommend a multi-layered review strategy to fully leverage the capability of LLMs - reviewers can first use LLMs to conduct an initial automatic review of those short and simple commits for security defects, then review both the submitted code and the LLM’s detection results to improve efficiency.
+This would allow reviewers to concentrate on an in-depth review of those complex commits, including lengthy code written by more experienced developers, thereby reducing the risk of missing sophisticated security defects.
+Additionally, where LLMs currently show limitations, these complex commits should be more heavily incorporated into datasets used for fine-tuning LLMs to improve their effectiveness in security code review.
 Providing a task scope is crucial for enhancing the performance of LLMs.
@@
 (Steenhoek et al., 2024).
-Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing the accuracy of detection.
-However, in real-world security reviews, although with the experience and domain knowledge, the reviewer is not guaranteed to consider all potential security defect types in the code.
+Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy.
+However, in real-world security reviews, although with the experience and domain knowledge, The reviewers lack prior knowledge of all potential types of security defects that may be present in the code.
 Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary.
@@
 In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs.
-This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to effectively focus on the target defect, thereby yielding adverse effects.
+This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects.
 Using LLMs in security code review requires developers to standardize their code commenting practices.
-According to RQ3, it appears that annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review.
+According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review.
 This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects.
 Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative.
-The more annotations there are in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers.
-Hence, we recommend developers write code comments in a standardized way to clarify the purpose of each comment, thus better support LLMs in code review tasks.
+The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers.
+Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better-supporting LLMs in code review tasks.
 With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs.
-The emergence of LLMs that support large amount of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks.
+The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks.
 However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects.
-This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs to overlook certain details, thereby failing to detect security defects introduced by small code snippets.
+This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets.
 Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2.
 Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
+A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results.
+The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses.
+This may be because the outputs generated by LLMs with higher randomness are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower randomness.
+Both detection capability and the stability of detection results are crucial for the practical application of LLMs in security code review.
+However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results.
+Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the randomness inherent in LLMs represents a valuable direction for future improvements.
 6. Threats to Validity
-Internal Validity: One threat could be data leakage.
+Internal Validity.
+One threat could be data leakage.
 The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs.
-Since the training data time of GPT-3.5 and GPT-4 is not clearly disclosed, we cannot avoid this overlap by filtering the data for a specific time period.
+Since the training data cutoff dates for GPT-3.5 and GPT-4 are not disclosed and the dataset we used (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time.
 Moreover, according to the findings by Cao et al.
-(Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cut-off date.
+(Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cutoff date.
 Another threat is the design of prompts.
-To evoke the best model performance, we design our prompt according to the best practices.
+We design our prompt according to best practices to evoke the best model performance.
 We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration.
+In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information.
+To specifically analyze its impact, we designed prompt $\textbf{P}_{cot}$, in which we instructed LLMs to generate code context according to the commit message before detecting security defects.
 Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness.
@@
 In RQ3, we combined the responses from the three experiments to fit the model.
-External validity: In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale).
+Conclusion Validity.
+In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale).
 In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies.
 We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
-Construct validity: With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure.
+External Validity.
+With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure.
 These models are also not made publically available via API, thereby hindering the reproducibility of this study.
-Therefore, we explicitly indicate the versions of these two models used in the experiments, and provided all the responses generated by these models (Anonymous, 2024).
+Therefore, we explicitly indicate the versions of these two models used in the experiments and provide all the responses generated by these models (Yu et al., 2024).
 In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements.
 Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
-7. Conclusions
+7. Conclusions and Future Work
 In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review.
 We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews.
-For the best-performing LLM, we conducted a linguistic analysis of the quality problems present in responses generated by this LLM and performed regression analysis to explore factors influencing its performance.
-Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to the state-of-the-art static analysis tools.
+For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM.
+We performed a regression analysis to explore factors influencing its performance.
+Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to state-of-the-art static analysis tools.
 However, their application remains limited.
 (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference.
-(3) During code review, GPT-4 is more effective in identifying security defects in code that has fewer tokens, containing functional logic and written by developers with less involvement in the project.
+(3) During code review, GPT-4 is more effective in identifying security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project.
 (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.
+We plan to extend this work in the next step by incorporating code reviews from proprietary software projects into our dataset and attempting enhancement strategies such as few-shot learning and fine-tuning.
+In addition to applying LLMs to help reviewers detect security defects during code review, we also plan to analyze the ability of LLMs to fix security defects in code so that developers can comprehensively evaluate the potential of LLMs in the task of secure code review.
+Acknowledgements.
+This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.

### v4

@@
-An Insight into Security Code Review with LLMs: Capabilities, Obstacles and Influential Factors
-Abstract.
+Conceptualization, Methodology, Investigation, Data curation, Formal analysis, Writing - Original draft preparation
+Abstract
 Security code review is a time-consuming and labor-intensive process typically requiring integration with automated security defect detection tools.
@@
 We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance.
-Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review butå significantly outperform the state-of-the-art static analysis tools.
+Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review but significantly outperform the state-of-the-art static analysis tools.
 (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference.
-(3) GPT-4 frequently generates responses that are verbose or not compliant with the task requirements given in the prompts.
+(3) GPT-4 frequently generates verbose or non-compliant responses with the task requirements given in the prompts.
 (4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
+keywords:
 Large Language Model, Code Review, Security Analysis
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software post-development issues
-1. Introduction
-Security defects in software programs can have severe consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004).
+1 Introduction
+Security defects are potential security risks or weaknesses in the system introduced during implementation (Paul et al., 2021).
+Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004).
 The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002).
@@
 Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021).
-Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013) and weak scalability (Nong et al., 2021).
+Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013), and weak scalability (Nong et al., 2021).
 Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021).
@@
 Nong et al.
-found that even using the training dataset augmented by the latest vulnerability injection technique (Nong et al., 2024b), the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively.
+(2024b) found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively.
 Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
 Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023).
-Such models show remarkable capability to understand and generate explanatory natural language responses, thus considered a potential candidate for security code review.
+Such models can understand and generate explanatory natural language responses, and are thus considered a potential candidate for security code review.
 We expect that LLM can act as a code reviewer, not only identifying security defects in code files, but also explaining the specific defect scenarios in detail to developers in a human-readable way.
 Applying LLMs to support security code review can help detect security defects earlier and improve productivity by helping code reviewers check their code faster and more efficiently.
-However, the application of LLMs in the specific field of security code review remains largely unexplored.
+However, applying LLMs in the specific field of security code review remains largely unexplored.
 Many studies have focused on leveraging LLMs for vulnerability detection.
 Zhou et al.
-evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts (Zhou et al., 2024b), while Purba et al.
-compared LLMs with static analysis tools in detecting software vulnerabilities (Purba et al., 2023).
-These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks, which was to judge whether there were security defects in the code.
-They did not require LLMs to provide more detailed information such as the location and type of vulnerability.
+(2024b) evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while Purba et al.
+(2023) compared LLMs with static analysis tools in detecting software vulnerabilities.
+These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks to judge whether there were security defects in the code.
+They did not require LLMs to provide more detailed information, such as the location and type of vulnerability.
 Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al.
-(Yang et al., 2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al.
-(Nong et al., 2024a).
+(2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al.
+(2024a).
 However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
-Motivated by the above-mentioned limitations, our aim is to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories.
-Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description and suggested fix.
+Motivated by the limitations mentioned above, we aim to bridge the knowledge gap by conducting an empirical study to comprehensively explore the potential of LLMs in fine-grained security code review on a practical dataset constructed from real-world code repositories.
+Specifically, in this study, LLMs were requested to provide detailed information on the identified security defect, including its line number, type, specific description, and suggested fix.
 The dataset we utilized consists of 534 code review files obtained from four open-source projects (namely, OpenStack Nova and Neutron, and Qt Base and Creator).
@@
 First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al.
-(Zhang et al., 2023)).
+(2023)).
 Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers.
@@
 In summary, this study makes the following contributions:
-  • We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review.
-  • We measured the impact of the randomness of LLMs on the consistency of security code review.
-  • We identified quality problems in the responses generated by the top-performing LLM to present its challenges in security code review.
-  • We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
+(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review.
+(2) We measured the impact of the non-determinism of LLMs on the consistency of security code review.
+(3) We identified quality problems and quantified hallucinations in the responses generated by the top-performing LLM, aiming to highlight the challenges faced by LLMs in security code review.
+(4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
 Paper Organization.
@@
 Finally, Section 7 concludes this work and outlines future directions.
-The replication package of this study has been provided online (Yu et al., 2024).
-2. Related Work
-Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, in this section, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or repair.
-Charoenwet et al.
-conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of vulnerability-contributing commits (VCCs) (Charoenwet et al., [n.
-d.]).
-By combining several tools, warnings could be produced for 78% of VCCs.
-However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review.
-Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
-Several studies have explored the performance of LLMs in detecting security defects.
-The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type.
-Zhou et al.
-utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgement under various prompts (Zhou et al., 2024b).
-They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art security defect detection approaches, while GPT-4 consistently outperformed the state-of-the-art.
-Purba et al.
-compared the performance of LLMs with popular static analysis tools in binary detection.
-Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools (Purba et al., 2023).
-Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description and suggestions for defect remediation.
-Previous studies also investigated whether LLMs could provide more information in security defect detection.
-Bakhshandeh et al.
-provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers(Bakhshandeh et al., 2023).
-Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools.
-Yin et al.
-evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description (Yin and Ni, 2024), found that the existing state-of-the-art approaches such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023) generally outperformed LLMs.
-Specifically, in defect localization and description tasks, LLMs exhibited limited overall accuracy and varied performance across different CWE types.
-These studies have all been limited to quantitatively evaluating LLM performance.
-In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to gain a comprehensive understanding of the potential and challenges LLMs face in security code review.
-Employing specific prompting frameworks to enhance the performances of LLMs in identifying security defects has also been explored.
-Yang et al.
-introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection (Yang et al., 2024).
-Their framework yielded superior performance compared to state-of-the-art prompting techniques.
-Nong et al.
-proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution) (Nong et al., 2024a).
-The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours.
-These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements.
-Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
-It is worth noting that the aforementioned studies focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs.
-Inspired by these related works, we aim to leverage the potential of LLMs in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.
-3. Methodology
-3.1. Research Questions
-The goal of this work is to comprehensively assess the capability of LLMs to assist with security code review by automatically detecting security defects in the given code.
-To achieve this goal, we divided our study into three Research Questions (RQs), as shown below.
+The replication package of this study has been provided online (Yu et al., 2025).
+2 Methodology
+2.1 Research Questions
+This work aims to comprehensively assess the capability of LLMs to assist with security code review by automatically detecting security defects in the given code.
+We divided our study into three Research Questions (RQs), as shown below.
 The research procedure for each RQ is illustrated in Fig.
@@
 RQ1.3: Can the Chain-of-Thought (CoT) prompting approach improve the performance of LLMs?
-RQ1.4: What is the impact of the randomness of LLMs on the consistency of security code review?
+RQ1.4: What is the impact of the non-determinism of LLMs on the consistency of security code review?
 Motivation.
@@
 It comprehensively evaluates the capability of LLMs to review a complete code file from real-world code reviews for security defects.
-Additionally, due to the inherent randomness of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt.
-However, this aspect of LLMs has not yet been studied when applied to security tasks.
-Hence, we conducted a quantitative analysis of the consistency of LLM responses to explore the impact of LLM randomness on the task of security code review.
-RQ2: What are the quality problems in LLM-generated responses during security code review?
+Additionally, due to the inherent non-determinism of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt across different runs.
+Non-determinism is mainly caused by decoding configurations (Song et al., 2024), with other hypotheses such as floating point, unreliable GPU calculations (Ouyang et al., 2025), and its sparse MoE architecture failing to enforce per-sequence determinism (La Malfa et al., 2024; Puigcerver et al., 2023).
+Considering that the non-determinism of LLMs has not yet been studied when applied to security tasks, we conducted a quantitative analysis of the consistency of LLM responses to explore its impact on the task of security code review.
+RQ2: What quality problems exist in LLM-generated responses during security code review?
+RQ2.1: What is the distribution of quality problems in these responses?
+RQ2.2: Which of these quality problems are hallucination-related, and what are their proportions?
 Motivation.
 Kabir et al.
-noted that the responses of ChatGPT to programming questions may be verbose, inconsistent, and contain conceptual or logical errors (Kabir et al., 2024), negatively impacting the quality of the answers.
-However, for security code review, the quality problems present in responses generated by LLMs have not been thoroughly analyzed in previous studies.
-Hence, RQ2 conducts an in-depth linguistic analysis of LLM responses for quality problems and examines the types and distribution of these problems to bridge the knowledge gap and indicate potential areas for improvement.
+(2024) noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers.
+Among these problems, hallucination in particular has been extensively studied  (Liu et al., 2024; Tonmoy et al., 2024).
+Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (Huang et al., 2025).
+If we applied LLMs to real-world software development tasks, particularly during security code review, such hallucination may mislead practitioners, increase review costs, and even compromise system security.
+Previous studies have not yet thoroughly examined quality problems present in LLM-generated responses during security code review.
+RQ2 intends to conduct an in-depth linguistic analysis to investigate possible hallucination issues.
+First, we investigated the distribution of quality problems in LLMs’ responses, such as Vague statement, Incorrect concept, and Incomplete answer.
+Then we identified hallucination-related issues among these problems and quantified their proportion to bridge the knowledge gap and highlight potential areas for improvement.
 RQ3: Which factors influence the performance of LLMs in security code review?
@@
 This RQ helps pinpoint areas of concern in applying LLMs to security code review and highlights possible improvement directions.
-Figure 1. An overview of the research procedure for investigating the three RQs
-3.2. Data Collection
+Figure 1: An overview of the research procedure for investigating the three RQs
+2.2 Data Collection
 We leveraged the code review dataset constructed by Yu et al.
-(Yu et al., 2023) and examined the capability of LLMs in security code review.
-The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (primarily written in Python), and Qt Base and Creator (primarily written in C++).
+(2023) and examined the capability of LLMs in security code review.
+The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (almost entirely written in Python), and Qt Base and Creator (mostly in C++ with some portions in C, Objective-C++, and Python).
 These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018).
@@
 During this procedure, if we cannot obtain the content of the code files through the API, we ignore those files.
-Lastly, since 97% files in this dataset are Python, C and C++ files, we decided to remove any other files (all are configuration and scripting files) as those only formed $\approx 3\%$ of the files and thus had a negligible impact on our results.
+Lastly, since 97% of the files in this dataset are Python and C/C++ files, we decided to remove any other files (all are configuration and scripting files) as those only formed $\approx 3\%$ of the files and thus had a negligible impact on our results.
 Our final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
-3.3. Research Procedure of RQ1
-Table 1. Versions and hyper-parameters of 6 LLMs studied.
-Model Version Context Length Parameter
+2.3 Research Procedure of RQ1
+Table 1: Versions and hyper-parameters of 6 LLMs studied.
+Model Version Ctx Len Parameter
 temperature top_p
@@
 Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
-3.3.1. LLM Selection
+2.3.1 LLM Selection
 To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023.
 Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API.
-For open-source LLMs, we deployed the smallest and largest chat versions of models in Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs.
-Throughout the experiments, we kept all hyper-parameters of these LLMs at their default settings.
-The versions, token constraints, and parameter settings of LLMs used in our study are summarized in Table 1.
+For open-source LLMs, we deployed the smallest and largest chat versions of models in the Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of VRAM.
+To realistically reflect on the practical capabilities of LLMs in security defect detection, we used the default hyperparameter settings throughout our experiments.
+Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (Atil et al., 2024; Blackwell et al., 2024).
+In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (Song et al., 2024) and may not guarantee determinism (Ouyang et al., 2025; Astekin et al., 2024).
+We retained the default temperature settings and quantified the non-determinism of LLMs as one aspect of performance evaluation in RQ1.4, rather than artificially suppressing such non-determinism via hyperparameter tuning.
+The versions, context lengths (Ctx Len), and parameter settings of LLMs used in our study are summarized in Table 1.
 Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform.
@@
 The two models provided by the ChatGPT platform performed better and thus were employed in our study.
-3.3.2. Prompt Design
+2.3.2 Prompt Design
 Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023).
 To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al.
-(Zhang et al., 2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt.
+(2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt.
 During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps.
-To fully harness the capabilities of LLMs in detecting security defects, we finally selected five best-performing prompts.
+To fully harness the capabilities of LLMs in detecting security defects, we finally selected the five best-performing prompts.
 Fig.
 2 illustrates the construction strategies for the five prompts.
-  1. (1)
   • Prompt 1 ($\textbf{P}_{b}$): the basic prompt.
@@
 If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
-  2. (2)
   • Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information.
@@
 The LLM is asked to act as a code reviewer for the project using the provided code in the prompt.
-  3. (3)
   • Prompt 3 ($\textbf{P}_{c}$): $\textbf{P}_{b}$ + General CWE Instruction.
 $\textbf{P}_{c}$ directly instructs the LLM to use CWE as a reference for identifying security defects without providing specific CWE details.
-  4. (4)
   • Prompt 4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction.
@@
 It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries.
-The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$)..
-  5. (5)
+The first level of CWEs in VIEW-1000 encompasses a broad range of commonly encountered security defects, thus applied in Prompt 4 ($\textbf{P}_{cid}$).
+We acknowledge that providing a list of CWEs may limit the scope of detection set of CWEs and miss security defects that do not fall into the provided CWE categories.
+However, including all CWEs is infeasible due to the limited context window of the LLM and the potential loss of focus.
   • Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$.
@@
 Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
-Figure 2. Construction templates for the five prompts
-Table 2. Adjusted IHMU-category for LLM responses.
-Classification Definition
-Instrumental (I) The response explicitly indicates the existence of the security defect identified by the reviewer and provides a fully accurate description.
-Helpful (H) The response raises concerns related to the security defect identified by the reviewer, but their descriptions may not be entirely accurate or specific enough.
+Figure 2: Construction templates for the five prompts
+Table 2: Adjusted IHMU-category for LLM responses.
+Classification Definition Instance
+Instrumental (I) The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario.
+Reviewer: (Line 76)I think this is race-y.
+Let’s say we have a three coroutines scheduled to call function.
+First one gets to L74 and function yields.
+Second coroutine gets in before threshold and gets to L68, sets is_deferred and sleeps on L70.
+First coroutine is finished and unsets on L76.
+Third coroutine enters still before threshold so L62 doesn’t influence it as planned and gets to L70, sleeps similar time as second coroutine.
+Then both, second and third coroutines call function in very short time difference.
+LLM: **Potential Time of Check to Time of Use (TOCTOU) race condition in ‘Throttler.call()‘** (Lines 58-77):
+- The is_deferred flag is used to determine if a function should be deferred.
+However, multiple threads can check the is_deferred flag almost simultaneously and hence more than one thread might execute the contained code.
+- **Solution**: A synchronization mechanism (like a lock) should be employed to ensure thread safety.
+Explanation: LLM indicated the security defect identified by the reviewer, with the correct type, accurate code line range and detailed description, thus categorized as Instrumental.
+Helpful (H) The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate.
+However, the response still provides useful evidence pointing toward the existence of the target defect.
+Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g.
+password) that mask_password looks for, you’d be replacing that with the masked token instead of masking the password or whatever it is.
+But I guess that’s unlikely.
+LLM:**Potential Information Disclosure**(Line 141):
+- Logging connection information can potentially expose sensitive details. Even though ‘mask_password‘ is being used, it’s essential to validate that no other sensitive information is leaked.
+- **Solution**: Limit the details being logged and ensure that all sensitive data is properly obfuscated before logging.
+Explanation: The reviewer indicated that the mask_password function might incorrectly replace different types of sensitive information.
+While the LLM correctly identified the relevant, it only provided general emphasis on checks for potential sensitive information exposure, failing to recognize this specific risk scenario.
+Consequently, this instance was categorized as Helpful rather than Instrumental.
 Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
+Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole.
+In this specific case a malicious user could create a ’raw’ image which actually contained a malicious qcow2 header specifying a backing file of the the compute host’s root disk.
+This would result in copying the compute host’s root disk to the specified rbd volume and exposing it to the user.
+LLM: Based on the provided code snippet, no security defects are detected.
+For a more comprehensive analysis, a full code audit in the context of the entire application and its environment would be recommended.
+Explanation: LLM did not detect any security defects, but the reviewer actually identified a security hole in the disk format detection. Therefore, this instance was considered Misleading.
 Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects.
 Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
-3.3.3. Evaluation Metrics
+Reviewer: (Line 469) So not even timeout cancels the future?
+Will this cause that threads will hang around forever leaking OS resources?
+This feels dangerous.
+LLM: **Potential Insecure Direct Object References (Lines 89-563):**
+- The ‘RequestContext‘ class and its methods provide direct access to project and user identifiers, and various methods in this file allow direct manipulation of database and message queue connections.
+If access controls are not properly enforced or if input validation is weak, this could lead to unauthorized access or manipulation of data.
+Explanation: The reviewer identified that the thread handling may lead to OS resource leaks.
+However, LLM did not mention any about resource leak, instead highlighting a potential improper access issue.
+Since access mechanisms might be uniformly managed by a middleware, and considering the unavailability of detailed implementation information, we cannot determine whether the issue pointed out by LLM exists.
+Thus this instance was categorized as Uncertain.
+2.3.3 Evaluation Metrics
 During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements.
 We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al.
-(Mahajan et al., 2020) for measuring responses in Stack Overflow.
+(2020) for measuring responses in Stack Overflow.
 We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al.
-because LLMs do not have a case of no response.
+(2020) because LLMs do not have a case of no response.
 Next, the definitions of each category were adjusted to match the task of security code review.
-Specifically, each response was categorized into one of the four rating categories detailed in Table 2.
+Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain) detailed in Table 2.
+To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description.
 The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$).
 Higher I-Score and IH-Score indicate a stronger ability of the model to detect security defects identified by reviewers.
-At the same time, a lower M-Score signifies less misleading results produced during security defect detection.
-Given that LLMs exhibit considerable randomness, we repeated the experiment for each LLM-prompt combination three times to mitigate the impact of such randomness.
-We calculated the average of the three performance scores as the final result.
-3.3.4. Baselines
-We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.
-To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects.
+At the same time, a lower M-Score signifies fewer misleading results produced during security defect detection.
+Given that LLMs exhibit considerable non-determinism, we repeated the experiment for each LLM-prompt combination three times and calculated the average of the three performance scores as the final result, so as to alleviate this threat and ensure a reliable evaluation of LLMs’ capabilities in security defect detection.
+2.3.4 Baselines
+We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects.
 This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022).
-Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Cppcheck Solutions AB, 2024) for C/C++ and Bandit (ban, 2024) for Python.
-Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SonarSource SA, 2024) to analyze Python files.
+Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Marjamäki et al., 2024) for C/C++ and Bandit (PyCQA, 2024) for Python.
+Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SA, 2024) to analyze Python files.
 These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first.
 Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files.
-Additionally, since Semgrep (sem, 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets.
+Additionally, since Semgrep (Dennison et al., 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets.
 To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities.
 Other tools utilized their default rule sets for analysis.
-3.3.5. Data Labelling
+2.3.5 Data Labelling
 Prompts were constructed using five templates and then fed into six LLMs to collect their responses.
@@
 Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization.
-Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ dataset.
+Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ datasets.
 We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 3.1).
 Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
-3.3.6. Consistency Calculation
-To measure the randomness of LLMs, we adopted the metrics in Chang et al.’s work (Chang et al., 2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments.
+2.3.6 Consistency Calculation
+To measure the non-determinism of LLMs, we adopted the metrics in Chang et al.
+(2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments.
 Specifically, an LLM generated three responses for a given code file under a specific prompt template.
@@
 A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
-3.4. Research Procedure of RQ2
+2.4 Research Procedure of RQ2
 We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses.
-We employed the open coding and constant comparative method (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
+We employed the open coding and constant comparison (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
 Based on the problem categories formulated by Kabir et al.
-(Kabir et al., 2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations.
+(2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations.
 Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types.
 In particular, we excluded the Consistency theme defined in Kabir et al.
-(Kabir et al., 2024) from our study.
+(2024) from our study.
 The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1.
@@
 Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement.
-Then the remaining data analysis was completed by the two authors and verified by the second author.
+Then the two authors completed the remaining data analysis and verified by the second author.
 Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis.
 For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
-3.5. Research Procedure of RQ3
+We consider the finalized codebook to cover all types of quality problems present in the LLM-generated responses within our dataset.
+To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by Huang et al.
+(2025), which categorizes hallucinations into two main types: Factual and Faithfulness.
+Factual Hallucination refers to outputs that contradict real-world facts or lack factual grounding, including Factual Contradiction and Factual Fabrication.
+Faithfulness hallucination, on the other hand, refers to deviations of the output from the instruction, current context, and reasoning process, including Instruction Inconsistency, Context Inconsistency, and Logical Inconsistency.
+We mapped the quality problems defined in our codebook to these hallucination subtypes to calculate their distribution across the dataset.
+Detailed mapping and results are provided in Section  3.2.
+2.5 Research Procedure of RQ3
 To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model.
@@
 In terms of the helpfulness of the response, by definition, I is more important than H, and so forth.
-Accounting for the ordinal nature of the response variable and the randomness of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008).
+Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008).
 This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together.
 The specific construction process is described in the following subsections.
-Table 3. Factors that may influence the performance of LLM in security code review
+Table 3: Factors that may influence the performance of LLM in security code review
 Factor Source* Definition Rationale
-Token Adjusted from  (Paul et al., 2021; McIntosh et al., 2016) Number of tokens of the code file.
-Computed using the TikToken (tik, 2024) tokenizer.
+Token Adjusted from  Paul et al.
+(2021); McIntosh et al.
+(2016) Number of tokens of the code file.
+Computed using the TikToken OpenAI (2024) tokenizer.
 Larger code files may be more difficult for LLMs to analyze.
@@
 The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
-SecurityDefectType Transformed from  (Chen et al., 2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al.
-(Yu et al., 2023).
-ChatGPT is proven to display distinct detection performances across various vulnerabilities (Chen et al., 2023; Zhou et al., 2023).
+SecurityDefectType Transformed from  Chen et al.
+(2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al.
+Yu et al.
+(2023).
+ChatGPT is proven to display distinct detection performances across various vulnerabilities Chen et al.
+(2023); Zhou et al.
+(2023).
 Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
-Commit Collected from  (Paul et al., 2021) Number of historical commits for the code file in its code change.
+Commit Collected from  Paul et al.
+(2021) Number of historical commits for the code file in its code change.
 Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
-Complexity Collected from  (Paul et al., 2021; McIntosh et al., 2016) McCabe’s Cyclomatic Complexity (McCabe, 1976).
-Computed using the Lizard (liz, 2024) analyzer.
-Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases (Paul et al., 2021).
+Complexity Collected from  Paul et al.
+(2021); McIntosh et al.
+(2016) McCabe’s Cyclomatic Complexity McCabe (1976).
+Computed using the Lizard Terry (2024) analyzer.
+Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases Paul et al.
+(2021).
 Community Created Name of the community source of the code file, i.e., OpenStack or Qt.
@@
 The LLMs’ performance may differ on data derived from these two communities.
-AuthorExperience Collected from  (Paul et al., 2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to.
+AuthorExperience Collected from  Paul et al.
+(2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to.
 Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
@@
 If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
-AnnotationHasCode Transformed from  (Chen et al., 2023) Whether the annotations in the code file include code.
-Code in annotations was abandoned by developers but often interpreted as actual code by LLMs (Chen et al., 2023), leading to incorrect security defect detection.
+AnnotationHasCode Transformed from  Chen et al.
+(2023) Whether the annotations in the code file include code.
+Code in annotations was abandoned by developers but often interpreted as actual code by LLMs Chen et al.
+(2023), leading to incorrect security defect detection.
 AnnotationisSecurityRelated Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects).
@@
     The sources of factors correspond to the descriptions in bold in Section 2.5.1.
-3.5.1. Attributes Construction
+2.5.1 Attributes Construction
 To comprehensively collect factors influencing the detection of security defects by LLMs, excluding the prompt and the model itself, we undertook the following steps:
 First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work.
-Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntoshet al.’s work (McIntosh et al., 2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens.
-Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al.’s work  (Chen et al., 2023), transforming them into factors for our experiment.
+Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntosh et al.
+(2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens.
+Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al.
+(2023), transforming them into factors for our experiment.
 Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations.
@@
 Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values.
-After filtering out code files with multiple security defects, we aggregated the dataset from three repetitive experiments.
+We aggregated the dataset from three repetitive experiments after filtering out code files with multiple security defects.
 We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination.
-This comprised 405 responses from each experiment and their factor values ready for model fitting.
-3.5.2. Regression Model Construction
+This comprised 405 responses from each experiment, and their factor values were ready for model fitting.
+2.5.2 Regression Model Construction
 Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis.
@@
 We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012).
-Utilizing the redun function in the rms package, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
-Figure 3. The Phi_K correlation coefficients of explanatory variables
-Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating the Spearman’s rank correlations (${\rho}^{2}$) between them.
-As Harrell  (Harrell, 2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned more degrees of freedom within the range of 3 to 5.
+Utilizing the redun function in the rms package [^2] https://cran.r-project.org/web/packages/rms/index.html, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
+Figure 3: The Phi_K correlation coefficients of explanatory variables
+Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating Spearman’s rank correlations (${\rho}^{2}$) between them.
+As  Harrell (2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.
+Paul et al.
+(2021) suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting.
 According to the computation results shown in Fig.
-4, for interval variables, we considered using the rcs function from the rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom.
+4, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom.
 Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
-Figure 4. Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
+Figure 4: Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
 Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015).
@@
 The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
-All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2024).
-4. Results
-4.1. Performance of LLMs (RQ1)
+All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2025).
+3 Results
+3.1 Performance of LLMs (RQ1)
 Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts.
 We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively.
-If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperforms the others.
+If two of the three metrics were superior (i.e., higher I-Score or IH-Score, lower M-Score), we deemed that the current combination outperformed the others.
 Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied.
@@
 This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions.
-As such features can only be accurately modelled using dynamic analysis, these security defects are challenging for static analysis tools to detect (Charoenwet et al., [n.
-d.]).
-Table 4. Performance scores and entropy for each LLM-Prompt combination on the Python dataset
+As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (Charoenwet et al., 2024).
+Table 4: Performance scores and entropy for each LLM-Prompt combination on the Python dataset
 Python
@@
 M-Score 91.47% 82.56% 78.29% 90.31%
-Table 5. Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
+Table 5: Performance scores and entropy for each LLM-Prompt combination on the C/C++ dataset
 C/C++
@@
 M-Score 98.19% 96.74%
-Table 6. Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
+Table 6: Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
 Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
@@
 M-Score 70.18% 80.12% 42.69% 20.47*44.44%
-4.1.1. RQ1.1
+3.1.1 RQ1.1
 When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4 and Table 5), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%).
@@
 Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
-4.1.2. RQ1.2
+3.1.2 RQ1.2
 As shown in Table 4 and Table 5, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small.
@@
 Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review.
-It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflect that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects.
+It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects.
 Still, it does not guarantee the accuracy of the detection results.
 Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4.
-Hence, how to leverage CWE information in the prompt is worthy of consideration.
-4.1.3. RQ1.3
+Hence, leveraging CWE information in the prompt is worthy of consideration.
+3.1.3 RQ1.3
 Table 4 and Table 5 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets.
@@
 Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
-4.1.4. RQ1.4
+3.1.4 RQ1.4
 According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency.
@@
 GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively.
-These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater randomness.
-4.2. Quality Problems in Responses (RQ2)
+These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater non-determinism.
+3.2 Quality Problems in Responses (RQ2)
+3.2.1 RQ2.1
+Figure 5: Average proportion of each problem type present in responses generated by GPT-4
 The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig.
 5, where the same color represents types within the same theme, with darker shades indicating the themes.
-Figure 5. Average proportion of each problem type present in responses generated by GPT-4
 From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%.
@@
 Among these, Vague statement is the most prevalent one.
-It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing a overly general fix suggestion.
+It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing an overly general fix suggestion.
 An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
@@
 As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason.
-However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
+However, this is not the content we required GPT-4 to generate in the prompt, hence it is deemed to be Unnecessary information.
 Example 3 - Unnecessary
@@
 Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts.
-Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks.
+Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform a security check in its response or misinterpreted requirements in the prompts for other tasks.
 Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found.
 Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt.
-From the example below (Example 4), we can see that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given.
+The example below (Example 4) shows that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given.
 Hence, it is categorized as an Incomplete answer.
@@
 While Kabir et al.
-(Kabir et al., 2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours.
+(2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours.
 Specifically, Kabir et al.
-requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues.
+(2024) requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues.
 In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%.
-This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific areas of application.
-4.3. Factors Influencing LLMs (RQ3)
-4.3.1. Model Stability
+This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific application areas.
+3.2.2 RQ2.2
+We selected all items related to hallucination from the quality problem types in RQ1.1 and mapped them to the predefined set of hallucination subtypes.
+The details of mapping, corresponding explanations, and proportion results are shown in Table 7.
+Since a single response may exhibit multiple types of hallucination simultaneously, the sum of the proportions for each subtype exceeds the proportion of responses containing at least one hallucination.
+Overall, 20.19 % of the responses contain hallucinations.
+Among these, Factual Contradiction is the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency and Factual Fabrication, which appear in 6.25% and 5.62% of responses, respectively.
+To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (Tonmoy et al., 2024; Lewis et al., 2020), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (Kumar et al., 2025).
+Table 7: Mapping of hallucination and quality problem types with corresponding proportions.
+Halluc. Type* Subtype QP Type* Explanation Prop.*
+Factuality Hallucination Factual Contradiction Incorrect Concept In Incorrect Concept cases, the model’s output contradicts the authoritative definition of the corresponding CWE, thus categorized as Factual Contradiction.
+10.13%
+Factual Fabrication Groundless Speculation Groundless speculation refers to subjective assumptions made without contextual evidence or knowledge support.
+As it involves fabricated factual content, it is categorized as Factual Fabrication.
+5.62%
+Faithfulness Hallucination Instruction Inconsistency Detection Failure Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction.
+Therefore, it is categorized as Instruction Inconsistency.
+0.15%
+Context Inconsistency Incorrect Fact Incorrect Fact denotes cases where the line numbers, identifiers, or code content generated by the model do not align with the provided code context, thus categorized as Context Inconsistency.
+6.25%
+Logical Inconsistency Logical Contradiction Logical Contradiction refers to cases where the model’s reasoning about a security defect is internally contradictory, thus classified as Logical Inconsistency.
+1.11%
+Any Hallucination 20.19%
+    Abbreviations: Halluc. = Hallucination, QP = Quality Problem, Prop. = Proportion.
+3.3 Factors Influencing LLMs (RQ3)
+3.3.1 Model Stability
 To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991).
 This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022).
-A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset.
+A possible explanation is that although we have considered the non-determinism of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset.
 Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model.
 Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
-Table 7. Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
+Table 8: Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
 Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(>Chi)
@@
     ${}^{*}p<0.05$; ${}^{**}p<0.01$; ${}^{***}p<0.001$
-4.3.2. Explanatory Power of Each Factor
+3.3.2 Explanatory Power of Each Factor
 The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable.
@@
 As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
-4.3.3. Relationships between Factors and Responses
+3.3.3 Relationships between Factors and Responses
 The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016).
@@
 The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively.
-This suggests that the LLM is more likely to detect security defects in code files that have fewer tokens, contain functional logic, and were written by developers with less involvement in the project.
+This suggests that the LLM is more likely to detect security defects in code files with fewer tokens, that contain functional logic, and were written by developers with less involvement in the project.
 Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt.
@@
 Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97.
-This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
-5. Discussion
+This suggests that annotations may disrupt the LLM and are less likely to detect the target security defect.
+4 Discussion
 A multi-layered review strategy is suggested to apply LLMs to security code review.
@@
 Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al.
-(Steenhoek et al., 2024).
+(2024).
 Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy.
-However, in real-world security reviews, although with the experience and domain knowledge, The reviewers lack prior knowledge of all potential types of security defects that may be present in the code.
+However, in real-world security reviews, although with experience and domain knowledge, the reviewers lack prior knowledge of all potential types of security defects that may be present in the code.
 Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary.
 It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses.
-In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs.
+In Bakhshandeh et al.
+(2023), providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs.
 This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects.
+Due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible.
+The lack of code context made it difficult for the LLM to fully understand the dependencies and interaction logic between files, which in turn affected its performance in identifying security defects, such as Race Condition and Insecure Library Version.
+We tested various approaches, including one-shot and few-shot prompting strategies, as well as providing the complete patchset within the prompt.
+These strategies required splitting the code into sections for input, which we found significantly degraded the performance of LLMs compared to the basic prompt $\textbf{P}_{b}$.
+In order to provide as much effective information as possible within the limited context window, we designed prompts that combined multiple elements, including project metadata, CWEs, code diffs, and the commit messages.
+We also standardized the output format to help the model better understand the task and introduced CoT reasoning to guide the model in inferring missing context.
+Ultimately, we identified and selected four prompts that outperformed others, which, together with the basic prompt, were selected as templates for our formal experiments.
+The results demonstrate that including a list of CWEs significantly enhances LLM performance, making it our study’s most effective prompting strategy.
+As the context window of LLMs continues to expand and prompt engineering techniques evolve, we plan to explore additional strategies in the future, such as incorporating RAG for external information and employing automated prompt optimization to dynamically adjust prompts during reasoning, thereby further improving the capability of LLMs in security defect detection.
 Using LLMs in security code review requires developers to standardize their code commenting practices.
@@
 The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers.
-Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better-supporting LLMs in code review tasks.
+Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better supporting LLMs in code review tasks.
 With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs.
@@
 This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets.
-Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2.
-Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
+Additionally, according to RQ1, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2.
+Therefore, ensuring the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
 A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results.
 The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses.
-This may be because the outputs generated by LLMs with higher randomness are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower randomness.
-Both detection capability and the stability of detection results are crucial for the practical application of LLMs in security code review.
+This may be because the outputs generated by LLMs with higher non-determinism are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower non-determinism.
+Both detection capability and the stability of detection results are crucial for applying LLMs in security code review.
 However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results.
-Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the randomness inherent in LLMs represents a valuable direction for future improvements.
-6. Threats to Validity
-Internal Validity.
-One threat could be data leakage.
-The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs.
-Since the training data cutoff dates for GPT-3.5 and GPT-4 are not disclosed and the dataset we used (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time.
-Moreover, according to the findings by Cao et al.
-(Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cutoff date.
+Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the non-determinism inherent in LLMs represents a valuable direction for future improvements.
+5 Threats to Validity
+Internal Validity refers to whether an observed association between two variables can be attributed to a causal link between them.
+One of the internal validity threats could be data leakage.
+The data used in our study was obtained from two large-scale open-source projects, which may overlap with the data used for training LLMs, a common problem with LLM-based studies.
+Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time.
+Moreover, according to Cao et al.
+(2024), contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date.
 Another threat is the design of prompts.
 We design our prompt according to best practices to evoke the best model performance.
-We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration.
+We selected words that are as unbiased, non-suggestive, and unambiguous as possible.
+We iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration.
 In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information.
-To specifically analyze its impact, we designed prompt $\textbf{P}_{cot}$, in which we instructed LLMs to generate code context according to the commit message before detecting security defects.
-Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness.
-To mitigate this threat, we conducted three repetitive experiments for RQ1.
+To specifically analyze its impact, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to generate code context according to the commit message before detecting security defects.
+External Validity concerns whether a causal link generalizes across contexts.
+In this study, the experiments were conducted on 6 LLMs based on 4 open-source projects across programming languages including Python and C/C++.
+We acknowledge that the dataset we utilized may not fully represent industrial projects across various programming languages, and the performance of the selected LLMs may not represent the capabilities of all existing LLMs.
+We implemented the following strategies to ensure the generalizability of the experimental results.
+For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular LLMs.
+For dataset construction, we collected code review data from real-world open-source projects in the OpenStack and Qt communities.
+These projects differ in architectures, programming languages, and community structures, providing a comprehensive and valuable dataset for empirical analysis.
+Construct Validity assesses whether a specific set of metrics corresponds to what they are intended to measure.
+To measure the performance of LLMs, we classified the natural language responses of LLMs into four predefined categories, with human reviews of security defects in code as the ground truth.
+There is a potential natural language understanding bias which affects classification accuracy.
+To reduce the bias, we established granular definitions for each category in three dimensions: defect type, location, and description, supplemented by examples for each category to ensure that the metrics are accurate, appropriate, and consistent throughout the evaluation process.
+In addition, the data labeling in RQ1 and the analysis of quality problems in RQ2 were all carried out manually.
+In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (Glaser, 1965).
+In the data labeling employed for RQ1, the first and third authors conducted a pilot data labeling and completed all the labeling work after reaching a consensus on the criteria through discussion.
+In the data analysis of RQ2, the first and seventh authors, conducted two rounds of pilot data analysis to determine the definition of each quality problem and achieved a high level of consistency in the results.
+All discrepancies in the manual work of our study were reviewed and resolved through negotiation and discussion among at least three authors, thus ensuring the reliability of the findings.
+Conclusion Validity refers to whether the conclusions drawn about relationships in the data are statistically sound and reasonable.
+Due to the non-determinism of LLMs, the same input may yield different outputs across different runs.
+Therefore, a single round of experiments is insufficient to accurately reflect the performance of LLMs.
+To enhance the reliability and representativeness of our evaluation, we repeated the RQ1 experiment three times.
 For RQ2, we calculated the average frequency of each problem across these three experiments.
-In RQ3, we combined the responses from the three experiments to fit the model.
-Conclusion Validity.
-In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale).
-In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies.
-We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
-External Validity.
-With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure.
-These models are also not made publically available via API, thereby hindering the reproducibility of this study.
-Therefore, we explicitly indicate the versions of these two models used in the experiments and provide all the responses generated by these models (Yu et al., 2024).
-In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements.
-Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
-7. Conclusions and Future Work
-In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review.
+For RQ3, we combined the responses from the three experiments to fit the model.
+We also assessed their non-determinism as part of the overall performance.
+Specifically, in RQ1.4, we calculated the consistency of LLMs’ responses across three runs to compare the non-determinism of different LLMs under different prompts.
+6 Related Work
+6.1 Security Defect Detection in Code Review
+Several studies have analyzed security defect detection during code reviews.
+Paul et al.
+(2021) built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects.
+The study analyzed 18 code review attributes, including review duration, reviewer experience, and author experience, to explore factors influencing the identification of security defects.
+It was found that the duration of code reviews and the frequency of peer reviews positively affected the detection of security defects, while the number of directories under review and the number of prior commits had a negative impact.
+Alfadel et al.
+(2023) examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews.
+Their work indicated that only a small fraction of PRs raise security problems, and most of the identified security problems were resolved during the code review.
+6.2 Code Characteristics of Security Defects
+Security defects mainly stem from improper system design or low-quality code, caused by a lack of security expertise or violations of secure coding practices during implementation.
+Many studies have focused on analyzing code for vulnerabilities and security defects in hopes of helping developers understand how security problems originate in a system, so that defects are detected early.
+Alves et al.
+(2016) constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others.
+This study demonstrated that most of these metrics are effective in distinguishing vulnerable from non-vulnerable code, thereby providing valuable support for security practitioners.
+Meneely et al.
+(2013) conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects.
+The findings suggest that longer lines of code and less experienced committers are more likely to introduce security defects.
+Bosu et al.
+(2014) identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with Meneely et al.
+(2013) that longer lines of code and less experienced committers are more likely to introduce security defects.
+Additionally, they indicated that modified code files are more likely to contain security defects than new files.
+6.3 Automated Tools of Security Defect Detection
+Various tools have been developed to automatically detect security defects.
+Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or program repair.
+Charoenwet et al.
+(2024) conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs.
+By combining several tools, warnings could be produced for 78% of the VCCs.
+However, at least 76% of these warnings are irrelevant to the vulnerability in VCCs, highlighting the challenges that SAST tools face in providing detailed information about the vulnerability during security code review.
+Inspired by this work, our work investigated the application of LLMs in security code review, aiming to overcome the limitations of SAST tools to some extent.
+Several studies have explored the performance of LLMs in detecting security defects.
+The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type.
+Zhou et al.
+(2024b) utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts.
+They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art detection approaches, while GPT-4 consistently outperformed the state-of-the-art.
+Purba et al.
+(2023) compared the performance of LLMs with popular static analysis tools in binary detection.
+Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools.
+Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description, and suggestions for defect remediation.
+Previous studies also investigated whether LLMs could provide more information in security defect detection.
+Bakhshandeh et al.
+(2023) provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers.
+Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools.
+Yin and Ni (2024) evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023), generally outperformed LLMs.
+Specifically, LLMs exhibited limited overall accuracy and varied performance across different CWE types in defect localization and description tasks.
+These studies have all been limited to quantitatively evaluating LLM performance.
+In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to understand the potential and challenges LLMs face in security code review.
+Employing specific prompting frameworks to enhance the performance of LLMs in identifying security defects has also been explored.
+Yang et al.
+(2024) introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection.
+Their framework yielded superior performance compared to state-of-the-art prompting techniques.
+Nong et al.
+(2024a) proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution).
+The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours.
+These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements.
+Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
+It is worth noting that the studies mentioned above focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs.
+Inspired by these related works, we aim to leverage LLMs’ potential in security analysis to identify the obstacles, application strategies, and improvement directions for integrating LLMs into security code review.
+7 Conclusions and Future Work
+In our study, we comprehensively analyzed the practicality of current popular LLMs in the context of security code review.
 We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews.
@@
 In addition to applying LLMs to help reviewers detect security defects during code review, we also plan to analyze the ability of LLMs to fix security defects in code so that developers can comprehensively evaluate the potential of LLMs in the task of secure code review.
-Acknowledgements.
-This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.
+Data availability
+We have shared the link to our dataset in the reference (Yu et al., 2025).
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.
+[…]

### v5

@@
 In this study, we conducted an empirical study to explore the potential of LLMs in detecting security defects during code review.
-Specifically, we evaluated the performance of six LLMs under five different prompts and compared them with state-of-the-art static analysis tools.
-We also performed linguistic and regression analyses for the best-performing LLM to identify quality problems in its responses and factors influencing its performance.
-Our findings show that: (1) existing pre-trained LLMs have limited capability in security code review but significantly outperform the state-of-the-art static analysis tools.
-(2) GPT-4 performs best among all LLMs when provided with a CWE list for reference.
-(3) GPT-4 frequently generates verbose or non-compliant responses with the task requirements given in the prompts.
-(4) GPT-4 is more adept at identifying security defects in code files with fewer tokens, containing functional logic, or written by developers with less involvement in the project.
+Specifically, we evaluated the performance of seven LLMs under five different prompts and compared them with state-of-the-art static analysis tools.
+We also performed linguistic and regression analyses for the two top-performing LLMs to identify quality problems in their responses and factors influencing their performance.
+Our findings show that: (1) In security code review, LLMs significantly outperform state-of-the-art static analysis tools, and the reasoning-optimized LLM performs better than general-purpose LLMs.
+(2) DeepSeek-R1 achieves the highest performance, followed by GPT-4.
+The optimal prompt for DeepSeek-R1 incorporates both the commit message and chain-of-thought (CoT) guidance, while for GPT-4, the prompt with a Common Weakness Enumeration (CWE) list works best.
+(3) GPT-4 frequently produces vague expressions and exhibits difficulties in accurately following instructions in the prompts, while DeepSeek-R1 more commonly generates inaccurate code details in its outputs.
+(4) LLMs are more adept at identifying security defects in code files that have fewer tokens and security-relevant annotations.
 keywords:
@@
 1 Introduction
-Security defects are potential security risks or weaknesses in the system introduced during implementation (Paul et al., 2021).
-Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (Telang and Wattal, 2007; Cavusoglu et al., 2004).
-The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (Planning, 2002).
-Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (GitLab, 2022).
+Security defects are potential security risks or weaknesses in the system introduced during implementation (paul2021security).
+Once attackers maliciously exploit a security defect, it could lead to serious consequences, including data breaches, financial losses, and service disruptions (telang2007empirical; cavusoglu2004effect).
+The longer a security defect stays in the program, the higher its associated fixing and maintenance costs are (planning2002economic).
+Thus, many organizations are shifting security analysis to earlier stages of software development, typically at code review time (gitlab2022url).
 Code review is a human-intensive process in which the reviewer examines the code submitted by the developer to detect bugs, verify the implementation of the specification, ensure compliance with guidelines, and ensure quality.
 During this process, the reviewer raises issues in the code, discusses them with the developer, and provides recommendations.
-Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (Weir et al., 2017), from being merged into the source code repository.
-In this study, we focus on security analysis in the context of code review, referring to it as security code review (Edmundson et al., 2013).
-Security code review is increasingly integrated into the development pipelines by project teams (Paul et al., 2021).
-Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (Guo et al., 2024).
-Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review (Gholami and Amri, 2021).
-Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (Nong et al., 2021), high false positive rates (Johnson et al., 2013; Sui et al., 2020), input range limitations (Haller et al., 2013), and weak scalability (Nong et al., 2021).
-Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (Chakraborty et al., 2021).
-However, due to the lack of sizable and realistic training datasets (Nong et al., 2022), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (Nong et al., 2024a).
-Nong et al.
-(2024b) found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively.
+Adopting security analysis in code review can incorporate diverse viewpoints from both reviewers and developers, and effectively prevent security defects created by programmers working alone, as they naturally have a limited perspective of potential security risks (weir2017d), from being merged into the source code repository.
+In this study, we focus on security analysis in the context of code review, referring to it as security code review (edmundson2013empirical).
+Security code review is increasingly integrated into the development pipelines by project teams (paul2021security).
+Its resource-intensive nature, which requires significant human effort and time to review and revise the code, poses a notable challenge, particularly in popular large-scale open-source projects with numerous contributions (guo2024exploring).
+Therefore, it has been desired to develop effective automated tools to assist code reviewers in security code review.
+Despite various static/dynamic/hybrid program analysis tools being proposed, they each encounter practical challenges, such as excessive imprecision (nong2021evaluating), high false positive rates (johnson2013don; sui2020recall), input range limitations (haller2013dowsing), and weak scalability (nong2021evaluating).
+Several data-driven approaches based on machine learning (ML)/deep learning (DL) have been developed for vulnerability detection and repair, which can somewhat handle the challenges encountered by traditional program analysis tools (chakraborty2021deep).
+However, due to the lack of sizable and realistic training datasets (nong2022generating), those approaches inevitably lacked robustness and tended to fall short when implemented in unfamiliar, real-world projects (nong2024chain).
+nong2024vgx found that even using the training dataset augmented by the latest vulnerability injection technique, the highest F1 score and the top-1 accuracy that state-of-the-art DL-based vulnerability detection and repair tools can achieve on the dataset constructed from real-world projects are still limited, at 20.01% and 21.05% respectively.
 Therefore, it remains essential to further explore techniques with better generalization and robustness to address vulnerabilities in real-world projects.
-Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (Jiang et al., 2023), test generation (Kang et al., 2023; Schäfer et al., 2023), and specification generation (Xie et al., 2023).
+Recently, pre-trained LLMs have demonstrated promising performances across a broad spectrum of software engineering tasks, such as program repair (jiang2023impact), test generation (kang2023large; schafer2023empirical), and specification generation (xie2023impact).
 Such models can understand and generate explanatory natural language responses, and are thus considered a potential candidate for security code review.
@@
 Many studies have focused on leveraging LLMs for vulnerability detection.
-Zhou et al.
-(2024b) evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while Purba et al.
-(2023) compared LLMs with static analysis tools in detecting software vulnerabilities.
+zhou2024largelanguage evaluated the capabilities of GPT-3.5 and GPT-4 in vulnerability detection under few-shot learning prompts, while purba2023software compared LLMs with static analysis tools in detecting software vulnerabilities.
 These studies mainly conducted a coarse-grained assessment of LLMs in binary classification tasks to judge whether there were security defects in the code.
 They did not require LLMs to provide more detailed information, such as the location and type of vulnerability.
-Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by Yang et al.
-(2024) and the vulnerability-semantic guided prompting approach formulated by Nong et al.
-(2024a).
+Researchers have also developed prompting strategies to enhance the effectiveness of LLMs in vulnerability detection, such as the DL-based prompting framework proposed by yang2024dlap and the vulnerability-semantic guided prompting approach formulated by nong2024chain.
 However, the datasets utilized in these studies are synthetic and CVE datasets, which do not fully represent real-world codebases.
@@
 We evaluated the performance of LLMs, analyzed existing quality problems, and examined factors influencing their performance to offer insights into the real-world applicability and limitations of LLM in security code review.
-First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by Zhang et al.
-(2023)).
-Then, we compared the performance of 6 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers.
-After selecting the responses generated by the best-performing LLM-prompt combination, we first manually extracted quality problems from these responses.
-We then used these responses to construct a cumulative link mixed model to explore the impact of 10 factors on the performance of the LLM.
-Our findings reveal that: (1) current popular general-purpose LLMs exhibit limited capability in detecting security defects identified by reviewers in code reviews.
-(2) GPT-4 performs best among all LLMs when provided with a CWE list for reference.
-(3) Inconcise and non-compliant responses to prompts are common problems encountered by GPT-4 in security code review.
-(4) GPT-4 is more likely to identify security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project.
-In summary, this study makes the following contributions:
-(1) We conducted a fine-grained evaluation of the capability of the currently popular LLMs in security code review.
+First, we identified five prompt templates to enhance LLMs (based on the prompting strategies formulated by zhang2024prompt).
+Then, we compared the performance of 7 open-source and closed-source LLMs with static analysis tools across these prompt templates in detecting security defects identified by reviewers.
+After collecting the responses generated by the two best-performing LLMs with their respective optimal prompts, we manually extracted quality problems present in these responses.
+We also used these responses to construct two separate cumulative link mixed models to explore the impact of 11 factors on the performance of each LLM.
+Our findings reveal that: (1) LLMs outperform state-of-the-art static analysis tools in security code review.
+Among LLMs, the reasoning-optimized model surpasses general-purpose models.
+(2) DeepSeek-R1 is the best-performing LLM, followed by GPT-4.
+For DeepSeek-R1, the prompt including the commit message and CoT instruction is the most effective, while for GPT-4, the prompt with a CWE list works best.
+(3) GPT-4 struggles with vague expressions and poor instruction-following, while DeepSeek-R1 is more prone to producing incorrect code details in its responses.
+(4) LLMs detect security defects more effectively in code files with fewer tokens and security-related annotations.
+In summary, this study makes the following contributions: (1) We conducted a fine-grained evaluation of the capability of popular and representative LLMs in security code review.
 (2) We measured the impact of the non-determinism of LLMs on the consistency of security code review.
-(3) We identified quality problems and quantified hallucinations in the responses generated by the top-performing LLM, aiming to highlight the challenges faced by LLMs in security code review.
-(4) We carried out the first analysis of the factors that may influence the performance of the LLM in security code review.
+(3) We identified quality problems and quantified hallucinations in the responses generated by the two top-performing LLMs, aiming to highlight the challenges faced by LLMs in security code review.
+(4) We carried out the first analysis of the factors that might influence the performance of LLMs in security code review.
 Paper Organization.
@@
 Finally, Section 7 concludes this work and outlines future directions.
-The replication package of this study has been provided online (Yu et al., 2025).
+The replication package of this study has been provided online (replpack).
 2 Methodology
@@
 The research procedure for each RQ is illustrated in Fig.
-1.
+2.
 RQ1: How effectively do LLMs detect security defects during code review?
@@
 Motivation.
-Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (Zhou et al., 2024a).
+Many studies have focused on the performance of LLMs in detecting security defects, but they often suffer from low-quality datasets collected automatically and a narrow detection scope primarily confined to function-level code (zhou2025large).
 To address these limitations, RQ1 utilizes a manually curated dataset to ensure data quality.
@@
 Additionally, due to the inherent non-determinism of LLMs, the responses of LLMs may exhibit considerable divergence for the same prompt across different runs.
-Non-determinism is mainly caused by decoding configurations (Song et al., 2024), with other hypotheses such as floating point, unreliable GPU calculations (Ouyang et al., 2025), and its sparse MoE architecture failing to enforce per-sequence determinism (La Malfa et al., 2024; Puigcerver et al., 2023).
+Non-determinism is mainly caused by decoding configurations (song2024good), with other hypotheses such as floating point, unreliable GPU calculations (ouyang2025empirical), and its sparse MoE architecture failing to enforce per-sequence determinism (la2024language; puigcerver2023sparse).
 Considering that the non-determinism of LLMs has not yet been studied when applied to security tasks, we conducted a quantitative analysis of the consistency of LLM responses to explore its impact on the task of security code review.
@@
 RQ2.1: What is the distribution of quality problems in these responses?
-RQ2.2: Which of these quality problems are hallucination-related, and what are their proportions?
+RQ2.2: Which of these problems are hallucination-related, and what are their proportions?
 Motivation.
-Kabir et al.
-(2024) noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers.
-Among these problems, hallucination in particular has been extensively studied  (Liu et al., 2024; Tonmoy et al., 2024).
-Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (Huang et al., 2025).
+kabir2023answers noted that LLMs’ responses may be verbose, inconsistent, and contain conceptual or logical errors, negatively impacting the quality of the answers.
+Among these problems, hallucination in particular has been extensively studied  (liu2024exploring; tonmoy2024comprehensive).
+Hallucination refers to content generated by LLMs that appears plausible but lacks factual grounding (huang2025survey).
 If we applied LLMs to real-world software development tasks, particularly during security code review, such hallucination may mislead practitioners, increase review costs, and even compromise system security.
@@
 Motivation.
-Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (Chen et al., 2023; Kabir et al., 2024; Zhou et al., 2023), primarily through case studies or observations.
+Previous studies have identified potential factors influencing the effectiveness of LLMs in detecting security defects (chen2023chatgpt; kabir2023answers; zhout2023devil), primarily through case studies or observations.
 However, there is a lack of evidence substantiating such claims.
@@
 This RQ helps pinpoint areas of concern in applying LLMs to security code review and highlights possible improvement directions.
+Figure 1: Distribution of LoC of the code files with security defects
 Figure 2: An overview of the research procedure for investigating the three RQs
+Figure 3: Distribution of security defect types on the Python and C/C++ dataset
 2.2 Data Collection
-We leveraged the code review dataset constructed by Yu et al.
-(2023) and examined the capability of LLMs in security code review.
+We leveraged the code review dataset constructed by yu2023security and examined the capability of LLMs in security code review.
 The dataset contains 614 review comments, each identifying a specific security defect across 15 predefined types from four projects: OpenStack Nova and Neutron (almost entirely written in Python), and Qt Base and Creator (mostly in C++ with some portions in C, Objective-C++, and Python).
-These projects are the most active within OpenStack and Qt (Hirao et al., 2020) and have been widely used in code review-related studies (Han et al., 2022; Hamasaki et al., 2013; Spadini et al., 2018).
+These projects are the most active within OpenStack and Qt (hirao2020code) and have been widely used in code review-related studies (han2022code; hamasaki2013does; spadini2018testing).
 For each security defect, we could not include the code from the entire patchset in our prompt, which is always too long and exceeds the input token limits of LLMs.
@@
 Our final dataset consists of 534 code files (258 Python and 276 C/C++ files), each including the source code, the commit message of the associated patchset, and detailed information of the target security defects as identified by the reviewers.
+We analyzed the distribution of security defect types in Python and C/C++ code files (see Fig.
+3).
+In Python code files, the Python language employs automatic memory management and garbage collection.
+Therefore, Memory Leak and Buffer Overflow were relatively rare, while Race Condition was the most common type, accounting for 65.1%.
+In contrast, C/C++ requires programmers to manage memory and pointers manually, thus the C/C++ dataset has a more balanced distribution of security defects.
+In C/C++ code files, Crash was the most prevalent (32%), followed by Race Condition (17.4%), Resource Leak (16.3%) and Integer Overflow (14.9%).
+We also counted the lines of code (LoC) of the collected code files in our dataset.
+Fig.
+1 shows the distribution of LoC for different languages (Python and C/C++, respectively) and the distribution of LoC for the entire dataset.
+We found that the median LoC is 815 and the average reaches 2,107, which substantially exceeds the scale of function-level datasets like Big-Vul (avg.
+30 LoC) and SVEN (avg.
+168 LoC) adopted in related studies (yin2024multitask; steenhoek2024err).
 2.3 Research Procedure of RQ1
 Table 1: Versions and hyper-parameters of 6 LLMs studied.
-Model Version Ctx Len Parameter
-temperature top_p
+Model Version Context Window Parameter
+temperature top_p Other parameters
 GPT-3.5 May 2024 version - -
 GPT-4 May 2024 version - -
-GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0
-Gemini Pro gemini-1.0-pro 32k 0.9 1.0
-Llama 2 7B llama-2-7b-chat 4k 1.0 0.9
-Llama 2 70B llama-2-70b-chat 4k 1.0 0.9
+GPT-4 Turbo gpt-4-1106-preview 128k 1.0 1.0 max_tokens: 4096; frequency_penalty: 0.0;
+Gemini Pro gemini-1.0-pro 32k 0.9 1.0 maxOutputTokens: 2048;
+Llama 2 7B llama-2-7b-chat 4k 1.0 0.9 max_seq_len: 4096;
+Llama 2 70B llama-2-70b-chat 4k 1.0 0.9 max_seq_len: 4096;
+DeepSeek-R1 deepseek-r1-250528 128k 1.0 0.7 max_tokens: 4096; frequency_penalty:0.0;
 2.3.1 LLM Selection
-To fully evaluate the performance of LLMs in security defect detection, we selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023.
+To fully evaluate the performance of LLMs in security defect detection, we firstly selected four representative LLMs from the proprietary GPT and Gemini families, along with two from the open-source Llama 2 family, which are the most popularly used LLMs when we started our experiments in December 2023.
 Specifically, for proprietary LLMs, we chose GPT-3.5 and GPT-4 provided by the ChatGPT platform (May 2024 version), and GPT-4 Turbo and Gemini Pro accessed through the API.
 For open-source LLMs, we deployed the smallest and largest chat versions of models in the Llama 2 series–namely, Llama 2 7B and Llama 2 70B–on a server with eight NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of VRAM.
+Given the rapid development of LLMs and the emergence of reasoning-optimized models, we additionally employed DeepSeek-R1 and accessed its updated version released on May 28, 2025 through the API.
 To realistically reflect on the practical capabilities of LLMs in security defect detection, we used the default hyperparameter settings throughout our experiments.
-Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (Atil et al., 2024; Blackwell et al., 2024).
-In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (Song et al., 2024) and may not guarantee determinism (Ouyang et al., 2025; Astekin et al., 2024).
+Some studies mitigate the non-determinism of LLMs by setting the temperature to zero (atil2024llm; blackwell2024towards).
+In contrast, we refrained from doing so, as this may adversely affect the performance of the LLMs (song2024good) and may not guarantee determinism (ouyang2025empirical; astekin2024exploratory).
 We retained the default temperature settings and quantified the non-determinism of LLMs as one aspect of performance evaluation in RQ1.4, rather than artificially suppressing such non-determinism via hyperparameter tuning.
-The versions, context lengths (Ctx Len), and parameter settings of LLMs used in our study are summarized in Table 1.
+The versions, context windows, and parameter settings of LLMs used in our study are summarized in Table 1.
 Notably, GPT-3.5 and GPT-4 can be accessed through both API and the ChatGPT platform.
-During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed by these two ways.
-The two models provided by the ChatGPT platform performed better and thus were employed in our study.
+During our preliminary exploration stage, we compared the performances of GPT-4 and GPT-3.5 accessed through these two ways.
+The two models provided by the ChatGPT platform performed better and were thus employed in our study.
+As the default parameters of the employed GPT-4 and GPT-3.5 were not publicly disclosed, the corresponding entries in Table 1 are indicated with ‘-’.
 2.3.2 Prompt Design
-Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (Liu et al., 2023).
-To design prompts, we followed the best practices published by OpenAI (OpenAI, 2023) and the three prompt design strategies in Zhang et al.
-(2023): basic prompt, enhanced prompt with auxiliary information, and CoT prompt.
+Prompt engineering is a process of designing prompts to optimize model performance on downstream tasks (liu2023pre).
+To design prompts, we followed the best practices published by OpenAI (GPT2023bestpractices) and the three prompt design strategies in zhang2024prompt: basic prompt, enhanced prompt with auxiliary information, and CoT prompt.
 During our preliminary exploration stage, we tried prompts with various auxiliary information and different CoT intermediate reasoning steps.
@@
 We instruct the LLM to review the provided code for security defects and output their description, code line numbers, and suggested fix.
-If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.
+If no defects are found, the LLM should output a fixed sentence, ‘No security defects are detected in the code.’.
+It is worth noting that we instruct the model to generate a solution for each identified security defect, since an incorrectly detected security defect usually cannot lead to a logically consistent fix in the source code.
+This process naturally drives the model to evaluate the validity of its detection and acts as an implicit self-reflection mechanism.
   • Prompt 2 ($\textbf{P}_{r}$): $\textbf{P}_{b}$ + Project Information.
@@
   • Prompt4 ($\textbf{P}_{cid}$): $\textbf{P}_{b}$ + Specific CWE Instruction.
-$\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (Pan et al., 2023)).
+$\textbf{P}_{cid}$ provides a specific list of the first level CWEs in VIEW-1000 (Research Concepts) [^1] VIEW-1000 is a weaknesses classification framework designed for academic research, widely adopted in prior work (e.g.,  (pan2023fine)).
 It organizes CWEs into a hierarchical structure based on parent-child relationships among CWE entries.
@@
 However, including all CWEs is infeasible due to the limited context window of the LLM and the potential loss of focus.
-  • Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$+$\textbf{P}_{cot-2}$.
+  • Prompt 5 ($\textbf{P}_{cot}$): $\textbf{P}_{cot-1}$$\textbf{P}_{cot-2}$.
 Since we only included the code of a single file in the prompt instead of an entire patchset, we designed $\textbf{P}_{cot}$ to analyze the impact of missing contextual information on the performance of LLMs.
@@
 Then $\textbf{P}_{cot-2}$ guides the LLM to review the provided code for security defects with reference to the generated code context.
+However, in the two-stage $\textbf{P}_{cot}$, the performance of LLM can be influenced by three factors: the commit message, CoT guidance, and the code context generated by LLM.
+Among these, the fabricated code context may contradict the actual code, leading to incorrect judgments.
+To isolate the impact of the LLM-generated context, we proposed a guardrail version of the $\textbf{P}_{cot}$ prompt and conducted experiments exclusively on DeepSeek-R1 under this additional prompt.
+  • Prompt 6 ($\textbf{P}_{cot-guardrail}$): This prompt retains the commit message and CoT instruction but does not use the code context generated by LLM, allowing us to evaluate the effect of fabricated contexts.
+The detailed template is shown in Fig 5.
 Figure 4: Construction templates for the five prompts
-Table 2: Adjusted IHMU-category for LLM responses.
-Classification Definition Instance
-Instrumental (I) The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario.
-Reviewer: (Line 76)I think this is race-y.
+Figure 5: Construction template for the guardrail version of $\textbf{P}_{cot}$
+2.3.3 Evaluation Metrics
+During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements.
+We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by mahajan2020recommending for measuring responses in Stack Overflow.
+We adapted the evaluation metrics formulated in their work (mahajan2020recommending) as follows: First, we excluded the ‘Unavailable’ category defined by mahajan2020recommending because LLMs do not have a case of no response.
+Next, the definitions of each category were adjusted to match the task of security code review.
+Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain).
+To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description.
+The detailed definition and corresponding examples of these four rating categories are as follows:
+Instrumental (I): The type of the security defect identified in the response is exactly the same as the target security defect, the location is strictly within the code line range of the target security defect, and the defect description can provide effective information for understanding the scenario.
+Example 1 - Instrumental Reviewer: (Line 76)I think this is race-y.
 Let’s say we have a three coroutines scheduled to call function.
-First one gets to L74 and function yields.
-Second coroutine gets in before threshold and gets to L68, sets is_deferred and sleeps on L70.
-First coroutine is finished and unsets on L76.
-Third coroutine enters still before threshold so L62 doesn’t influence it as planned and gets to L70, sleeps similar time as second coroutine.
+the first may yield at L74, the second enters before the threshold and sleeps at L70 while setting is_deferred, then the first finishes and unsets at L76, allowing the third to enter before the threshold as well.
 Then both, second and third coroutines call function in very short time difference.
-LLM: **Potential Time of Check to Time of Use (TOCTOU) race condition in ‘Throttler.call()‘** (Lines 58-77):
+LLM: Race condition in ‘Throttler.call()‘(Lines 58-77):
 - The is_deferred flag is used to determine if a function should be deferred.
 However, multiple threads can check the is_deferred flag almost simultaneously and hence more than one thread might execute the contained code.
-- **Solution**: A synchronization mechanism (like a lock) should be employed to ensure thread safety.
-Explanation: LLM indicated the security defect identified by the reviewer, with the correct type, accurate code line range and detailed description, thus categorized as Instrumental.
-Helpful (H) The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate.
+Explanation: LLM indicated the race condition defect identified by the reviewer, with the accurate code line range and detailed description, thus categorized as Instrumental.
+Helpful (H): The type or location of the identified security defect differ to some extent from the target security defect, or the description may be too broad or partially inaccurate.
 However, the response still provides useful evidence pointing toward the existence of the target defect.
-Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g.
+Example 2 - Helpful Reviewer: (Line 141) If connect_info ever includes a secret of a type other than token (e.g.
 password) that mask_password looks for, you’d be replacing that with the masked token instead of masking the password or whatever it is.
 But I guess that’s unlikely.
-LLM:**Potential Information Disclosure**(Line 141):
+LLM:Potential Information Disclosure(Line 141):
 - Logging connection information can potentially expose sensitive details. Even though ‘mask_password‘ is being used, it’s essential to validate that no other sensitive information is leaked.
-- **Solution**: Limit the details being logged and ensure that all sensitive data is properly obfuscated before logging.
-Explanation: The reviewer indicated that the mask_password function might incorrectly replace different types of sensitive information.
+Explanation: The reviewer noted that the mask_password function might incorrectly replace different types of sensitive information.
 While the LLM correctly identified the relevant, it only provided general emphasis on checks for potential sensitive information exposure, failing to recognize this specific risk scenario.
-Consequently, this instance was categorized as Helpful rather than Instrumental.
-Misleading (M) The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
-Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole.
+So this instance was categorized as Helpful rather than Instrumental.
+Misleading (M): The response does not provide any helpful information, e.g., claiming no defects were found or failing to perform the detection task.
+Example 3 - Misleading Reviewer: (Line 329) This does implicit format detection on the imported disk, which is a well-known security hole.
 In this specific case a malicious user could create a ’raw’ image which actually contained a malicious qcow2 header specifying a backing file of the the compute host’s root disk.
@@
 For a more comprehensive analysis, a full code audit in the context of the entire application and its environment would be recommended.
-Explanation: LLM did not detect any security defects, but the reviewer actually identified a security hole in the disk format detection. Therefore, this instance was considered Misleading.
-Uncertain (U) The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects.
+Explanation: The reviewer pointed out a sensitive data exposure at line 329, whereas the LLM did not detect any security defects. Therefore, it was classified as Misleading.
+Uncertain (U): The response does not mention any defects identified by the reviewer but instead points out the existence of other security defects.
 Due to the lack of expertise of the researchers about the code and its context, the actual existence of these security defects cannot be confirmed.
-Reviewer: (Line 469) So not even timeout cancels the future?
-Will this cause that threads will hang around forever leaking OS resources?
-This feels dangerous.
+Example 4 - Uncertain Reviewer: (Line 469) So not even timeout cancels the future? Will this cause that threads will hang around forever leaking OS resources? This feels dangerous.
 LLM: **Potential Insecure Direct Object References (Lines 89-563):**
@@
 Thus this instance was categorized as Uncertain.
-2.3.3 Evaluation Metrics
-During the experiments, the responses of LLMs were provided in the form of natural language text, making it necessary to manually inspect these responses to ensure the accuracy of the performance measurements.
-We considered that LLM-generated responses could be evaluated in a manner similar to the approach adopted by Mahajan et al.
-(2020) for measuring responses in Stack Overflow.
-We adapted the evaluation metrics formulated in their work (Mahajan et al., 2020) as follows: First, we excluded the ‘Unavailable’ category defined by Mahajan et al.
-(2020) because LLMs do not have a case of no response.
-Next, the definitions of each category were adjusted to match the task of security code review.
-Specifically, each response was categorized into one of the four rating categories (Instrumental, Helpful, Misleading, Uncertain) detailed in Table 2.
-To ensure clearer differentiation, for the ‘Helpful’ and ‘Instrumental’ categories, we specified their definitions based on their performance across three key attributes: defect type, location, and description.
 The corresponding evaluation metrics are I-Score ($\frac{I}{I+H+U+M}\times 100\%$), IH-Score ($\frac{I+H}{I+H+U+M}\times 100\%$), and M-Score ($\frac{M}{I+H+U+M}\times 100\%$).
@@
 Given that LLMs exhibit considerable non-determinism, we repeated the experiment for each LLM-prompt combination three times and calculated the average of the three performance scores as the final result, so as to alleviate this threat and ensure a reliable evaluation of LLMs’ capabilities in security defect detection.
+It is important to note that the cost and latency of LLMs in security code review are significant concerns for practitioners.
+However, in this study, these two points were not specifically discussed.
+Due to the various invocation methods of the LLMs involved (including calls through the ChatGPT platform, local deployment, and API calls), it is difficult to provide a consistent cost estimate across different LLMs.
+Moreover, the latency of the models is significantly affected by their deployment environments, such as computational resources, inference engines, and parallelization strategies, making the latency of different LLMs in our experiments incomparable.
 2.3.4 Baselines
 We tried different tools and learning techniques for detecting security defects to select appropriate baselines for our work.To compare with LLMs, we focused on techniques providing fine-grained detection, including classification, localization, and description of defects.
-This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (Ren et al., 2024), GRACE (Lu et al., 2024) and LineVul (Fu and Tantithamthavorn, 2022).
-Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (Marjamäki et al., 2024) for C/C++ and Bandit (PyCQA, 2024) for Python.
-Furthermore, we also used CodeQL (GitHub, 2021) and SonarQube (SA, 2024) to analyze Python files.
+This led us to exclude techniques with insufficient detection granularity, such as ProRLearn (ren2024prorlearn), GRACE (lu2024grace) and LineVul (fu2022linevul).
+Given that we evaluated the performance of LLMs on the C/C++ and Python datasets separately, for language-specific static analysis tools, we selected CppCheck (cppcheck) for C/C++ and Bandit (bandit) for Python.
+Furthermore, we also used CodeQL (codeql) and SonarQube (sonarqube) to analyze Python files.
 These two tools can directly analyze code files written in dynamically-typed languages like Python but require code files in statically-typed languages like C/C++ to be compiled first.
 Therefore, we could not apply them to our C/C++ dataset, which is comprised of uncompiled and isolated C/C++ code files.
-Additionally, since Semgrep (Dennison et al., 2024) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets.
+Additionally, since Semgrep (semgrep) supports multiple programming languages and can analyze individual code files, we adopted it for both C/C++ and Python datasets.
 To obtain comprehensive analysis results, CodeQL adopted the Python-security-and-quality.qls test suite, covering a wide range of issues from basic code structure and naming conventions to advanced security and performance vulnerabilities.
@@
 2.3.5 Data Labelling
-Prompts were constructed using five templates and then fed into six LLMs to collect their responses.
-Since the responses of LLMs were provided in the form of natural language text, to rate the responses accurately, we manually inspected the content of LLM-generated responses, their corresponding source code, and security defects identified by reviewers to categorize them into the four categories formulated in Table 2.
-To mitigate bias, pilot data labelling was conducted, in which we randomly selected 100 out of 534 code files from our dataset, constructed five prompts for each file and fed these prompts into LLMs to collect responses.
-Using the definitions of categories in Table 2 as the labelling criteria, the first and third authors labelled the collected responses separately.
-Discrepancies in labelling results were discussed with the second author to reach a consensus.
-The inter-rater reliability was calculated using Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.89.
-Then, the first and third authors labelled all the rest of the responses and discussed any ambiguous cases with the second author to finalize the categorization.
-Based on the labelling results, we calculated the performance scores of each LLM-prompt combination in the Python and C/C++ datasets.
-We found that the combination of GPT-4 and prompt $\textbf{P}_{cid}$ outperformed the others (as detailed in Section 3.1).
-Therefore, we aggregated the responses generated by GPT-4 with $\textbf{P}_{cid}$ across the Python and C/C++ datasets, resulting in a total of 421 responses (some code files exceeded the token limitation of GPT-4) for each of the three repetitive experiments, to further analyze RQ2 and RQ3.
+[…]
 2.3.6 Consistency Calculation
-To measure the non-determinism of LLMs, we adopted the metrics in Chang et al.
-(2024) and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments.
+To measure the non-determinism of LLMs, we adopted the metrics in chang2024beyond and utilized entropy to measure the consistency of responses generated by LLMs across three repetitive experiments.
 Specifically, an LLM generated three responses for a given code file under a specific prompt template.
@@
 We determined the overall consistency by averaging the entropy across all code files using the following formula: $\frac{1}{R}\sum_{r\in R}\text{entropy}(p^{r}_{1},\dots,p^{r}_{4})$, where R is the response count in each round of the experiment.
-A higher average entropy suggests the LLM has higher inconsistency in repetitions of experiments.
+A higher mean entropy suggests the LLM has higher inconsistency in repetitions of experiments.
+To further quantify the uncertainty of the entropy estimates and ensure the reliability of our findings, bootstrap confidence intervals were calculated for each LLM–prompt combination.
+Following the guidance of  efron1994introduction, we resampled the original dataset with replacement to match its original size and calculated the mean entropy for each resample.
+As suggested by  davison1997bootstrap, if confidence levels 0.95 and 0.99 are to be used, then it is advisable to have [the number of resamples]=999 or more, if practically feasible.
+Therefore, consistent with the settings adopted in  ahmed2025can, we performed 1,000 bootstrap replications to obtain the 95% confidence intervals for the entropy estimates.
 2.4 Research Procedure of RQ2
-We manually inspected the responses of GPT-4 with $\textbf{P}_{cid}$ across the three repetitive experiments to analyze the quality problems present in these responses.
-We employed the open coding and constant comparison (Glaser, 1965) with a predefined category by the MAXQDA tool for qualitatively analyzing the responses.
-Based on the problem categories formulated by Kabir et al.
-(2024) and the problems of LLM-generated text in various domains identified by previous studies (Borji, 2023; Mitrović et al., 2023; Kalla and Kuraku, 2023), two of the authors firstly inspected a few GPT-4 responses independently and recorded their observations.
-Then they collaboratively reviewed the inspected responses and discussed to establish an initial codebook, resulting in four themes: Correctness, Understandability, Conciseness, and Compliance, each with a series of problem types.
-In particular, we excluded the Consistency theme defined in Kabir et al.
-(2024) from our study.
-The reason is that the consistency between the LLM’s responses and the reviewers’ assessments (i.e., the ground truth) has been investigated in RQ1.
-Following the codebook, we conducted a pilot data analysis across two iterations to improve the agreement between the two authors.
-In each round, ten responses were randomly selected for the two authors to analyze their quality problems.
-The results were then compared and discussed, with partial adjustments made to the codebook.
-Since a single response could contain multiple problems, two authors completed the remaining data analysis task where the labels are not mutually exclusive.
-Therefore, we calculated the agreement level using Krippendorff’s alpha (Krippendorff, 2004) with Jaccard distance, which was improved from 0.40 in the first round to 0.80 in the second, indicating a high level of agreement.
-Then the two authors completed the remaining data analysis and verified by the second author.
-Throughout the entire procedure, we employed a negotiated agreement approach (Campbell et al., 2013): any conflicts were consulted and addressed by three authors, ensuring the reliability of data analysis.
-For each type of problem, we computed its average frequency in three repetitive experiments as the final occurrence rate of this problem in GPT-4 responses.
+[…]
+[…]
 We consider the finalized codebook to cover all types of quality problems present in the LLM-generated responses within our dataset.
-To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by Huang et al.
-(2025), which categorizes hallucinations into two main types: Factual and Faithfulness.
+To further analyze the proportion of hallucinations in LLMs’ responses, we adopted the granular taxonomy proposed by huang2025survey, which categorizes hallucinations into two main types: Factual and Faithfulness.
 Factual Hallucination refers to outputs that contradict real-world facts or lack factual grounding, including Factual Contradiction and Factual Fabrication.
@@
 2.5 Research Procedure of RQ3
-To answer RQ3, we adopted the approach formulated by Harrell (Harrell, 2001) to construct a regression model.
-From Table 2, we can see that the response variables in our dataset are the ratings of I, H, U, and M.
+To answer RQ3, we adopted the approach formulated by Harrell (harrell2001regression) to construct a regression model.
+As mentioned in Section 2.3.3, the response variables in our dataset are the ratings of I, H, U, and M.
 In terms of the helpfulness of the response, by definition, I is more important than H, and so forth.
-Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (Powers and Xie, 2008).
+Accounting for the ordinal nature of the response variable and the non-determinism of the LLM in three rounds of experiments, we ultimately aggregated the data from all three experiments to fit a cumulative link mixed model (powers2008statistical).
 This model supports ordered categorical response variables and allows us to include random effects to group responses in each round together.
@@
 Factor Source* Definition Rationale
-Token Adjusted from  Paul et al.
-(2021); McIntosh et al.
-(2016) Number of tokens of the code file.
-Computed using the TikToken OpenAI (2024) tokenizer.
+Token Adjusted from  paul2021security; mcintosh2016empirical Number of tokens of the code file.
+Computed using the TikToken tiktoken tokenizer.
 Larger code files may be more difficult for LLMs to analyze.
+Position Collected from  Sovrano_2025 Position of the security defect within the code file.
+Determined by calculating the proportion of tokens preceding the defect relative to the total token count, using the TikToken tiktoken tokenizer.
+LLMs tend to significantly underperform when detecting security defects located toward the end of files (Sovrano_2025; rafi2024impact).
 FileType Created Type of the code file, i.e., Source or Auxiliary.
@@
 The understanding capability of LLMs and the scenarios of security defects may vary across different types of files.
-SecurityDefectType Transformed from  Chen et al.
-(2023) Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by Yu et al.
-Yu et al.
-(2023).
-ChatGPT is proven to display distinct detection performances across various vulnerabilities Chen et al.
-(2023); Zhou et al.
-(2023).
+SecurityDefectType Transformed from  chen2023chatgpt Type of the security defect that the reviewers identified in this code file, i.e., one of the 15 types of security defect formulated by yu2023security.
+ChatGPT is proven to display distinct detection performances across various vulnerabilities chen2023chatgpt; zhout2023devil.
 Thus, we assume that LLMs may be more adept at detecting certain specific types of security defects in code.
-Commit Collected from  Paul et al.
-(2021) Number of historical commits for the code file in its code change.
+Commit Collected from  paul2021security Number of historical commits for the code file in its code change.
 Excessive history commits may lead developers to fatigue or overconfidence, resulting in their ignorance of security considerations in simpler scenarios.
-Complexity Collected from  Paul et al.
-(2021); McIntosh et al.
-(2016) McCabe’s Cyclomatic Complexity McCabe (1976).
-Computed using the Lizard Terry (2024) analyzer.
-Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases Paul et al.
-(2021).
+Complexity Collected from  paul2021security; mcintosh2016empirical McCabe’s Cyclomatic Complexity mccabe1976complexity.
+Computed using the Lizard lizard analyzer.
+Code comprehensibility may become more challenging for LLMs as its cyclomatic complexity increases paul2021security.
 Community Created Name of the community source of the code file, i.e., OpenStack or Qt.
@@
 The LLMs’ performance may differ on data derived from these two communities.
-AuthorExperience Collected from  Paul et al.
-(2021) Number of closed code commits the author has submitted prior to the commit this code file corresponding to.
+AuthorExperience Collected from  paul2021security Number of closed code commits the author has submitted prior to the commit this code file corresponding to.
 Experienced authors may write more rigorous and standardized code, while the security vulnerabilities in the code are also relatively more subtle and difficult to detect.
@@
 If the token number of all annotations in the file f is $\textbf{a}_{t}$ and the total token number of this file is $\textbf{f}_{t}$, then $\textbf{AR}=\textbf{a}_{t}/\textbf{f}_{t}$ The capability of LLMs to comprehend manual annotations and code may vary.
-AnnotationHasCode Transformed from  Chen et al.
-(2023) Whether the annotations in the code file include code.
-Code in annotations was abandoned by developers but often interpreted as actual code by LLMs Chen et al.
-(2023), leading to incorrect security defect detection.
+AnnotationHasCode Transformed from  chen2023chatgpt Whether the annotations in the code file include code.
+Code in annotations was abandoned by developers but often interpreted as actual code by LLMs chen2023chatgpt, leading to incorrect security defect detection.
 AnnotationisSecurityRelated Created The relevance of annotations to the security defects identified by reviewers in this code file, i.e., 1 (annotations mentioned target security defects identified by reviewers), 0 (Annotations were unrelated to security) and -1 (Annotations mentioned non-target security defects).
@@
 To comprehensively collect factors influencing the detection of security defects by LLMs, excluding the prompt and the model itself, we undertook the following steps:
-First, we collected the attributes from previous works (Paul et al., 2021; McIntosh et al., 2016; Thongtanunam et al., 2017) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work.
-Given that LLMs understand and generate text token by token, we adjusted the factor adopted from McIntosh et al.
-(2016) — ‘Size’, to ‘Token’ (see the first row of Table 3), as larger code files typically contain more tokens.
-Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in Chen et al.
-(2023), transforming them into factors for our experiment.
+First, we collected the attributes from previous works (paul2021security; mcintosh2016empirical; thongtanunam2017review) that were utilized in analyzing relationships related to code review and applicable to the code file data in our work.
+Given that LLMs understand and generate text token by token, we adjusted the factor adopted from mcintosh2016empirical — ‘Size’, to ‘Token’ (see the first row of Table 2), as larger code files typically contain more tokens.
+Next, we extracted characteristics of data that were explicitly indicated to affect or correlate with the capabilities of LLMs in chen2023chatgpt; Sovrano_2025, transforming them into factors for our experiment.
 Lastly, after thoroughly observing and analyzing the distribution of responses, we further created and adopted several potentially influential factors based on our observations.
-In the end, we obtained a list of 10 factors in Table 3, along with the source, definition, and rationale for each factor.
-Since acquiring a factor like ‘AuthorExperience’ entails querying the code review history of projects, we utilized Gerrit API to retrieve detailed information for all patchsets under each code change.
-We used Python scripts complemented with manual analysis to query, compute, and label these factors.
-Besides, since we adopted ‘SecurityDefectType’ in our factor list, we only considered data where the code file contains a single security defect to avoid the case in which ‘SecurityDefectType’ could have multiple values.
-We aggregated the dataset from three repetitive experiments after filtering out code files with multiple security defects.
-We obtained a total of 1,215 responses generated by the GPT-4 and $\textbf{P}_{cid}$ combination.
-This comprised 405 responses from each experiment, and their factor values were ready for model fitting.
+In the end, we obtained a list of 11 factors in Table 2, along with the source, definition, and rationale for each factor.
+[…]
+Figure 6: The Phi_K correlation coefficients of explanatory variables
 2.5.2 Regression Model Construction
 Correlation and Redundancy Analysis: If explanatory variables are highly correlated, it can cause interference in model construction and analysis.
-As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (Baak et al., 2020), to measure the correlation between each factor.
+As the predefined explanatory variables in our study include both categorical and continuous variables, we utilized a refined Pearson’s hypothesis test of independence — Phi_K (baak2020new), to measure the correlation between each factor.
 The correlation coefficients obtained are illustrated in Fig.
-3.
-We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (Hinkle et al., 2003).
+6.
+We chose $|\rho|>0.7$ as the threshold because it is recommended as the threshold for strong correlation (hinkle2003applied).
 From Fig.
-3, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors.
+6, where an abbreviation denotes each factor, it can be seen that the correlation coefficients between each pair of factors are all below 0.7, suggesting that there is no strong correlation between these factors.
 However, two factors without high correlation can still be duplicates, distorting the relationship between explanatory and response variables in model fitting.
-We encoded categorical variables as dummy variables for redundancy analysis (Legendre and Legendre, 2012).
+We encoded categorical variables as dummy variables for redundancy analysis (legendre2012numerical).
 Utilizing the redun function in the rms package [^2] https://cran.r-project.org/web/packages/rms/index.html, we assessed potential redundant variables with its default threshold of $R^{2}\geq 0.9$, ultimately finding none.
-Figure 3: The Phi_K correlation coefficients of explanatory variables
 Degrees of Freedom Allocation: To assign degrees of freedom to each explanatory variable, we measured the potential of nonlinear relationships between each factor and the response variable by calculating Spearman’s rank correlations (${\rho}^{2}$) between them.
-As  Harrell (2001) recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.
-Paul et al.
-(2021) suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting.
+As  harrell2001regression recommended, explanatory variables with larger ${\rho}^{2}$ can be assigned higher degrees of freedom.
+paul2021security suggested limiting the maximum allocated degrees of freedom for an explanatory variable to five or fewer, thus reducing the risk of overfitting.
 According to the computation results shown in Fig.
-4, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom.
+7, for interval variables, we considered using the rcs function from the R’s rms package to allocate 3 degrees of freedom for the factor ‘Token’ with the highest ${\rho}^{2}$, while the other factors are allocated 1 degree of freedom.
 Category variables should be converted into factor variables, and their corresponding degrees of freedom are automatically allocated during model fitting.
 Figure 7: Dotplot of the Spearman ${\rho}^{2}$ between explanatory and response variables across three repetitive experiments
-Model Fitting: Next, we fitted the 1,215 responses and their rating categories onto a cumulative link mixed model using the clmm function from the ordinal R package (Christensen and Christensen, 2015).
-We tested models with and without allocating 3 degrees of freedom to the ‘Token’ factor to balance model complexity and fit.
-The model that did not allocate these degrees of freedom, which had a lower Akaike Information Criterion (AIC), was then selected as our final fitting result (Burnham and Anderson, 2004).
-All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (Yu et al., 2025).
+[…]
+All the results and scripts utilized in data labelling, extraction, and regression analysis are provided in our replication package (replpack).
 3 Results
 3.1 Performance of LLMs (RQ1)
-Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across five different prompts.
+Based on the evaluation metrics defined in Section 2.3.3, we evaluated the capability of six popular LLMs in security code review across different prompts.
 We compared the performance of LLMs with baseline tools in the C/C++ and Python datasets, respectively.
@@
 Due to differences in the token limit of each LLM and the inherent token count of each prompt template, the number of responses successfully generated by each LLM-prompt combination also varied.
-To ensure rigorous comparisons, we selected the intersection of source code files in our complete dataset so that all LLM-prompt combinations could generate responses, resulting in 57 instances.
+To avoid bias due to different samples and to ensure rigorous comparisons, we also performed analysis on the intersection of source code files that all LLM-prompt combinations could generate responses, which yielded 57 instances.
 As a result, two sets of scores are provided: (1) the response count and performance scores of each LLM-prompt combination on the complete Python and C/C++ dataset (see Table 3 and Table 4), and (2) the performance scores of each LLM-prompt combination on the selected 57 instances (see Table 5).
 The best values for each score are highlighted in bold and marked with asterisks in the tables.
-From Table 4 and Table 5, it is evident that, on the Python dataset, the combination of GPT-4 with $\textbf{P}_{cid}$ achieved the best performance.
-However, on the C/C++ dataset, the combination of Llama 2 70B with $\textbf{P}_{cid}$ performs the best.
-However, Table 6 demonstrated that in the selected 57 cases, Llama 2 70B scored lower than GPT-4.
-We speculate that this discrepancy could be attributed to the smaller volume of responses obtained by Llama 2 70B, leading to inflated performance scores of Llama 2 70B in the complete datasets.
-On the whole, we consider GPT-4 paired with $\textbf{P}_{cid}$ as the top-performing LLM-prompt combination.
-As of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited.
+It is evident that, DeepSeek-R1 outperformed all other models, followed by GPT-4.
+Specifically, Table 3 shows that on the Python dataset, the combination of DeepSeek-R1 with $\textbf{P}_{cid}$ achieved the best performance.
+According to Table 4, on the C/C++ dataset, the combination of DeepSeek-R1 with $\textbf{P}_{cot-guardrail}$ performed the best.
+Furthermore, Table 5 demonstrates that in the selected 57 cases, DeepSeek-R1 with $\textbf{P}_{b}$ achieved better results than all other LLM-prompt combinations.
+Although GPT-4 did not perform as well as DeepSeek-R1, it still obviously surpassed the remaining five LLMs.
+For the top two models, DeepSeek-R1 and GPT-4, we calculated the performance scores of these two models across different prompt templates on the full dataset (including both Python and C/C++ datasets), as shown in Table 6.
+We found that DeepSeek-R1 performed best with $\textbf{P}_{cot-guardrail}$, whereas GPT-4 achieved its highest scores with $\textbf{P}_{cid}$.
+On the whole, as of the time of our experiment (May 2024), the capabilities of currently popular LLMs in conducting security code reviews are still limited.
 Notably, with enhanced prompts, the best performance of each LLM in security code review is significantly superior to that of static analysis tools.
-This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (Russo and Sabelfeld, 2010); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions.
-As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (Charoenwet et al., 2024).
+This could be attributed to two reasons: (1) for context-sensitive tools such as CodeQL, static analysis can be impacted by the lack of code context in our dataset (russo2010dynamic); (2) there is a relatively high frequency of security defects related to multi-threading and asynchronous programming, such as race conditions.
+As such, features can only be accurately modelled using dynamic analysis, and these security defects are challenging for static analysis tools to detect (charoenwet2024empirical).
 Table 3: Performance scores and entropy for each LLM-Prompt combination on the Python dataset
 Python
-Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
+Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot-guardrail}$
 GPT-3.5 Resp. Count 123 123 123 123 122
@@
 GPT-4 Resp. Count 175 175 175 174 174
-I-Score 4.76% 3.43% 3.24% 5.17*3.06%
-IH-Score 9.34% 9.52% 8.38% 14.56*8.62%
+I-Score 4.76% 3.43% 3.24% 5.17% 3.06%
+IH-Score 9.34% 9.52% 8.38% 14.56*% 8.62%
 M-Score 61.31% 60.95% 46.47% 44.06% 61.69%
@@
 GPT-4 Turbo Resp. Count 201 201 201 201 201
-I-Score 0.66% 0.33% 0.66 0.83 0.17
+I-Score 0.66% 1.65% 0.66 0.83 0.17
 IH-Score 0.66 0.83 2.32 1.99 0.33
 M-Score 97.18% 96.35% 85.74% 92.87% 98.01%
-Entropy 0.1757 0.0746 0.3370 0.6125 0.0404*
+Entropy 0.0777 0.0901 0.2424 0.1873 0.0548
 Gemini Pro Resp. Count 162 161 160 160 159
@@
 M-Score 90.54% 97.09% 84.04% 55.21% 98.53%
-Entropy 0.0777 0.0901 0.2424 0.1873 0.0548
+Entropy 0.1757 0.0746 0.3370 0.6125 0.0404*
 Llama 2 7B Resp. Count 62 62 61 61 50
@@
 IH-Score 1.61% 1.61% 5.40% 8.74% 4.67%
-M-Score 87.10% 83.33% 56.78% 40.98% 33.33*%
+M-Score 87.10% 83.33% 56.78% 40.98% 33.33%
 Entropy 0.2625 0.3070 0.4766 0.4351 0.5521
@@
 Entropy 0.4470 0.3474 0.3514 0.3832 0.5762
+DeepSeek-R1 Resp. Count 234 234 234 234 234 234
+I-Score 5.41% 6.27% 4.70% 7.26%* 4.56% 5.84%
+IH-Score 7.83% 7.55% 6.27% 9.26% 6.41% 8.69%
+M-Score 36.04% 32.05% 29.34%* 30.77% 46.87% 35.69%
+Entropy 0.5440 0.5273 0.5098 0.5444 0.5683 0.5526
 Baselines SAST SonarQube CodeQL Bandit Semgrep
@@
 C/C++
-Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
+Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot-guardrail}$
 GPT-3.5 Resp. Count 141 141 141 139 137
@@
 M-Score 98.02% 98.64% 91.08% 93.43% 96.28%
-Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
+Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
 Gemini Pro Resp. Count 223 221 221 221 220
@@
 M-Score 92.97% 98.19% 89.44% 61.54% 99.09%
-Entropy 0.0512 0.0332 0.1401 0.1576 0.1024
+Entropy 0.1430 0.0457 0.2493 0.5413 0.0250
 Llama 2 7B Resp. Count 57 57 57 51 33
@@
 IH-Score 2.34% 3.51% 5.85% 4.58% 7.07%
-M-Score 90.06% 87.13% 81.29% 52.94% 27.27*%
+M-Score 90.06% 87.13% 81.29% 52.94% 27.27%
 Entropy 0.2650 0.2167 0.2167 0.4092 0.5954
 Llama 2 70B Resp. Count 57 57 57 51 37
-I-Score 1.17% 1.17% 1.75% 6.54*1.80%
-IH-Score 4.68% 2.92% 3.51% 15.03*6.31%
+I-Score 1.17% 1.17% 1.75% 6.54% 1.80%
+IH-Score 4.68% 2.92% 3.51% 15.03% 6.31%
 M-Score 67.25% 80.12% 66.08% 40.52% 54.05%
 Entropy 0.5667 0.4145 0.3895 0.5515 0.6153
+DeepSeek-R1 Resp. Count 269 269 269 269 269 269
+I-Score 12.64% 8.43% 9.91% 10.61% 9.79% 12.89%*
+IH-Score 15.49%* 10.41% 11.15% 12.27% 12.64% 15.49%*
+M-Score 29.99% 20.94%* 21.93% 23.05% 34.32% 27.39
+Entropy 0.5411 0.4934 0.4884 0.5002 0.5654 0.5663
 Baselines SAST CppCheck Semgrep
@@
 Table 5: Performance scores for each LLM-Prompt combination on 57 instances that all LLM-prompt combinations could generate responses
-Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$
+Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot}$
 GPT-3.5 I-Score 3.51% 3.51% 0.58% 0.58% 2.34%
@@
 M-Score 81.87% 79.53% 94.74% 95.32% 77.19%
-GPT-4 I-Score 8.77% 8.77% 5.85% 12.87*9.94
-IH-Score 20.46% 19.30% 14.04% 23.98*19.88
+GPT-4 I-Score 8.77% 8.77% 5.85% 12.87% 9.94
+IH-Score 20.46% 19.30% 14.04% 23.98*% 19.88
 M-Score 43.86% 42.11% 45.61% 40.35% 41.52%
@@
 IH-Score 4.68% 2.34% 7.02% 17.54% 4.68%
-M-Score 70.18% 80.12% 42.69% 20.47*44.44%
+M-Score 70.18% 80.12% 42.69% 20.47*% 44.44%
+DeekSeek-R1 I-Score 16.96%* 16.37% 12.87% 16.37% 13.45% 12.28%
+IH-Score 23.98%* 19.88% 15.79% 21.64% 18.71% 18.13%
+M-Score 40.35% 37.43% 35.67% 30.99% 45.61% 32.75%
+Table 6: Performance scores and entropy for GPT-4 and DeepSeek-R1 under different prompts on the complete dataset
+Model & Prompt $\textbf{P}_{b}$ $\textbf{P}_{r}$ $\textbf{P}_{c}$ $\textbf{P}_{cid}$ $\textbf{P}_{cot}$ $\textbf{P}_{cot-guardrail}$
+GPT-4 Resp. Count 422 422 422 421 421
+I-Score 4.58% 3.71% 3.63% 5.30%* 4.35%
+IH-Score 8.53% 8.63% 8.92% 12.91%* 10.29%
+M-Score 57.89% 62.88% 45.33% 41.09%* 51.07%
+DeepSeek-R1 Resp. Count 503 503 503 503 503 503
+I-Score 9.28% 7.43% 7.49% 9.05% 7.36% 9.61%*
+IH-Score 11.93% 9.08% 8.88% 10.87% 9.74% 12.33%*
+M-Score 32.80% 26.11% 25.38%* 26.64% 40.16% 31.25
 3.1.1 RQ1.1
-When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 4 and Table 5), GPT-4 performs the best, obtaining the highest I-Score (Python: 4.76%, C/C++: 4.45%), IH-Score (Python: 9.34%, C/C++: 7.96%), and the lowest M-Score (Python: 61.31%, C/C++: 55.47%).
-Following closely behind is Llama 2 70B and GPT-3.5.
+When using basic prompts constructed by $\textbf{P}_{b}$ (see Table 3 and Table 4), DeepSeek-R1 performed the best, obtaining the highest I-Score (Python: 5.41%, C/C++: 12.64%), and the lowest M-Score (Python: 36.04%, C/C++: 29.99%).
+For the IH-Score, DeepSeek-R1 still achieved the highest value on the C/C++ dataset, reaching 15.49%, while on the Python dataset, GPT-4 outperformed DeepSeek-R1 (7.83%) with a score of 9.34%.
+GPT-4 was regarded as the second-best model, followed by Llama 2 70B and GPT-3.5.
 GPT-4 Turbo, Gemini Pro, and Llama 2 7B performed the worst, with their performance in the Python dataset even inferior to that of the baseline tool, Bandit.
 Similar results are observed in Table 5.
-Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, GPT-4 achieved I-Score = 8.77%, IH-Score = 20.46%, and M-Score = 43.86%.
-It is evident that GPT-4’s performance on the selected 57 instances exceeds its performance across the entire dataset.
+Specifically, in the selected 57 instances, under the basic prompt $\textbf{P}_{b}$, DeepSeek-R1 achieved I-Score = 16.96%, IH-Score = 23.98%, and M-Score = 40.35%.
+It is evident that DeepSeek-R1’s performance on the selected 57 instances exceeds its performance across the entire dataset.
 Given that these 57 code files are instances with relatively fewer tokens within our dataset, it could be speculated that the quantity of tokens input to LLM affects its performance in detecting security defects, which has been further explored and analyzed in RQ3.
 3.1.2 RQ1.2
-As shown in Table 4 and Table 5, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small.
-Specifically, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively.
+$\textbf{P}_{r}$: As shown in Table 3 and Table 4, the changes of the performance scores of each LLM under prompt $\textbf{P}_{r}$ compared with that under $\textbf{P}_{b}$ is generally small.
+Specifically, DeepSeek-R1, GPT-3.5, GPT-4, and GPT-4 Turbo show subtle increases and decreases in performance on the Python and C/C++ datasets, respectively.
 Gemini Pro and Llama 2 70B demonstrate minor performance decreases on both datasets, while Llama 2 7B shows a slight improvement.
@@
 We speculate that some LLMs may already understand the task described in the prompt quite well, and thus, introducing a persona could add unnecessary complexity to the prompt and lead to distractions.
-We can see from Table 4 and Table 5 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of GPT-3.5 significantly decreases under $\textbf{P}_{c}$ and $\textbf{P}_{cid}$ in both Python and C/C++ datasets.
+$\textbf{P}_{c}$: We can see from Table 3 and Table 4 that compared to using the basic prompt $\textbf{P}_{b}$, the performance of DeepSeek-R1 and GPT-3.5 significantly decreases under $\textbf{P}_{c}$ in both Python and C/C++ datasets.
 The performances of GPT-4 with $\textbf{P}_{c}$ exhibit different changes across different datasets, with an improvement in the C/C++ dataset but a decrease in the Python dataset.
 However, all other LLMs achieved higher performance scores across both datasets.
-Moreover, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B, and Llama 2 70B all achieved their best performance under $\textbf{P}_{cid}$.
-Therefore, it can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of security defect detection of various LLMs in code review.
-It is worth noting that under $\textbf{P}_{cid}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects.
+$\textbf{P}_{cid}$: Under $\textbf{P}_{cid}$, GPT-4, GPT-4 Turbo, Gemini Pro, Llama 2 7B and Llama 2 70B all achieved their best performance on both the Python and C/C++ datasets, while GPT-3.5 experienced a significant drop in scores.
+However, the impact of $\textbf{P}_{cid}$ on the performance of DeepSeek-R1 was language-specific: while it led to a decline in C/C++, it enhanced performance in Python.
+It can be inferred that adding information related to CWE to the prompt generally enhances the capabilities of general-purpose LLMs in security code review.
+Notably, under $\textbf{P}_{c}$, the decrease in GPT-4’s I-Score (Python: 4.76% →3.24%, C/C++: 4.45% →3.91%) and in M-Score (Python: 61.31% →46.47%, C/C++:55.47% →44.53%) reflects that guiding LLMs to reference CWE in the prompt indeed enables GPT-4 to detect more security defects.
 Still, it does not guarantee the accuracy of the detection results.
 Under $\textbf{P}_{cid}$, providing a specific CWE list as a reference in the prompt leads to significant improvements across all metrics and datasets for GPT-4.
-Hence, leveraging CWE information in the prompt is worthy of consideration.
+Hence, incorporating specific CWE entries in the prompt is worthy of consideration.
+For DeepSeek-R1, we analyzed why the impact of $\textbf{P}_{cid}$ on performance is programming language-specific.
+As mentioned in Section 2.2, Race Condition is the most prevalent security defect type in the Python dataset.
+The CWE list provided in $\textbf{P}_{cid}$ includes CWE-435, which is highly relevant to Race Condition, thereby enhancing model performance.
+In contrast, the C/C++ dataset features a more balanced distribution of security defect types, such as Integer overflow, Resource Leak and Crash.
+We included solely high-level CWE entries in the CWE list and some of them, such as CWE-682 (Incorrect Calculation), do not exactly match certain security defect types like Integer Overflow.
+These CWE entries failed to provide sufficient guidance and even introduced interference.
+More fine-grained subtypes may provide better guidance for specific security defect types.
+Therefore, determining the optimal granularity of CWE as auxiliary information for different LLMs in security defect detection warrants further investigation in future work.
 3.1.3 RQ1.3
 Table 3 and Table 4 illustrate that $\textbf{P}_{cot}$ significantly enhances the ability to detect security defects for Llama 2 7B and Llama 2 70B in both Python and C/C++ datasets.
-However, for Gemini Pro and GPT-4 Turbo, there is no improvement and even a detrimental effect to these two LLMs.
+However, for Gemini Pro, GPT-4 Turbo and DeepSeek-R1, there is no improvement and even a detrimental effect to these LLMs.
 As for GPT-3.5 and GPT-4, they respectively exhibit improvement and decrease in their performances across different datasets.
@@
 Consequently, when Prompt $\textbf{P}_{cot-2}$ was applied, the detection of security defects in the provided code was misled by the generated code context and yielded erroneous results.
+To isolate the impact of LLM-generated code context on model performance, we designed $\textbf{P}_{cot-guardrail}$ and conducted experiments under this prompt on the best-performing model, DeepSeek-R1.
+$\textbf{P}_{cot-guardrail}$ retains CoT guidance and commit message but does not include the fabricated code context.
+The results are also presented in Table 3, Table 4 and Table 5.
+We observed that, on both the Python and C/C++ datasets, $\textbf{P}_{cot-guardrail}$ consistently improved the performance of DeepSeek-R1 compared with $\textbf{P}_{b}$, whereas $\textbf{P}_{cot}$ led to a performance decline.
+This suggests that for DeepSeek-R1, LLM-generated code context introduced interference for security code review.
+However, in the selected 57 instances, DeepSeek-R1 performed worse with $\textbf{P}_{cot-guardrail}$ than with $\textbf{P}_{b}$.
+The 57 samples satisfy the token limits of all LLM-prompt combinations in our experiments, making them the samples with the fewest tokens in the dataset.
+Based on this, we hypothesize the following explanation: the commit message included in $\textbf{P}_{cot-guardrail}$ helps the model understand code semantics and the CoT instruction drives the model to perform step-by-step reasoning, making this prompt more suitable for assisting the model in analyzing security defects in complex and long code.
+For code files with fewer tokens, these additional guidance could be unnecessary or even introduce noise to the LLM.
+Missing code context can impact the reliability of model outcomes (li2025everything), which remains a challenge when applying LLMs to real-world security code review.
+We attempted to generate code context using the LLM itself, but did not directly measure how often the LLM hallucinated irrelevant code and how such hallucinations affected model performance.
+This is because evaluating the consistency between LLM-generated code and the actual source code is challenging, as it requires a comprehensive understanding of the codebase and its historical commits.
+We plan to address this in future work.
+Some works have performed context augmentation by incorporating call graphs extracted through static analysis tools (luo2024fellmvp) or by utilizing caller-callee relationships within the code (wen2024vuleval; gao2023far; sun2024llm4vuln).
+Given this, we will also compare these approaches with the LLM-driven code context generation in assisting LLMs with security code reviews, and explore how to better leverage the capabilities of LLMs to obtain more effective and precise code context.
 3.1.4 RQ1.4
-According to the mean entropies of various LLM-prompt combinations in Table 4 and Table 5, GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B and Llama 2 70B, with GPT-4 showing the least consistency.
-Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with GPT-4 Turbo, with a mean entropy of 0.0404.
-On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195.
+Figure 8: Response Entropy of each LLM-prompt combinations on the Python dataset
+Figure 9: Response Entropy of each LLM-prompt combinations on the C/C++ dataset
+According to the mean entropies and their 95% confidence intervals (CI) of various LLM-prompt combinations in Fig.
+8 and Fig.
+9, we found that GPT-3.5, GPT-4 Turbo, and Gemini Pro exhibit relatively higher consistency in their responses across three repetitive experiments, followed by Llama 2 7B, Llama 2 70B and DeepSeek-R1, with GPT-4 showing the least consistency.
+Specifically, on the Python dataset, the highest consistency of responses was observed under $\textbf{P}_{cot}$ with Gemini Pro, with a mean entropy of 0.0404 (95% CI: [0.02, 0.07]).
+On the C/C++ dataset, the combination of GPT-3.5 and $\textbf{P}_{c}$ demonstrates the lowest mean entropy of 0.0195 (95% CI: [0.00, 0.05]).
 Given these two LLM-prompt combinations’ poor security defect detection capabilities, their high consistency may be due to their tendency not to detect any security defects in the code.
-GPT-4 paired with $\textbf{P}_{cid}$ is considered the top-performing combination in security code review but shows high mean entropies on both Python and C/C++ datasets, with 0.9416 and 0.6955 respectively.
-These results suggest that LLMs with more robust security defect detection capabilities may also exhibit greater non-determinism.
+On the Python dataset, the highest average entropy of 0.9416 (95% CI: [0.89, 0.99]) was observed with the combination of GPT-4 and $\textbf{P}_{cot}$.
+On the C/C++ dataset, GPT-4 paired with $\textbf{P}_{cid}$ exhibited the highest average entropy of 0.7330, with a 95% CI from 0.67 to 0.79.
+Compared with other LLMs excluding DeepSeek-R1, GPT-4 not only demonstrated significantly superior performance, but also exhibited the highest uncertainty.
+In comparison with GPT-4, DeepSeek-R1 showed improved performance in the security code review task but produced more stable outputs.
+However, the average entropy of DeepSeek-R1 remained relatively high among all models evaluated in our experiments.
+As shown in Table 1, we kept the decoding parameters at their default settings for all models.
+Except for DeepSeek-R1, the temperature and top_p values of all LLMs fell within a similar range (0.9–1.0).
+Despite this, we observed substantial differences in response consistency across models.
+Furthermore, although DeepSeek-R1 was configured with a lower top_p value (0.7), its response consistency was still relatively poor among all models.
+This observation suggests that decoding parameters alone may have a limited effect on regulating output consistency, and the observed differences are more likely attributable to model-intrinsic randomness.
 3.2 Quality Problems in Responses (RQ2)
@@
 Figure 10: Average proportion of each problem type present in responses generated by GPT-4
-The average distribution of 12 problem types across four themes in responses to three repetitive experiments is illustrated in Fig.
-5, where the same color represents types within the same theme, with darker shades indicating the themes.
-From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%.
-Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met.
-Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution.
-Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%).
+[…]
+For GPT-4, most responses contained problems related to Conciseness, accounting for 90.18%.
+Nearly half of the responses (41.56%) involved Compliance-related problems.
+Furthermore, 27.08% of the responses showed problems with Understandability and problems related to Correctness were the least frequent, occurring in only 14.08% of the responses.
+In the responses of DeepSeek-R1, Conciseness problems were also the most prevalent, accounting for 84.29%.
+However, the frequencies of problems related to Compliance and Understandability were significantly lower than those of GPT-4, at 4.51% and 9.54%, respectively.
+Moreover, DeepSeek-R1 was much more prone to Correctness problems than GPT-4, with a rate of 33.66%.
 The typical types of problems within each theme are described in detail below.
-Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%).
+Correctness: The responses revealed two types of incorrectness: Incorrect fact (GPT-4: 6.25%, DeepSeek-R1: 31.81%) and Incorrect concept (GPT-4: 10.13%, DeepSeek-R1: 2.39%).
 The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario.
@@
 As shown in the example below (Example 1), the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
+Compared with GPT-4, DeepSeek-R1 exhibited a lower frequency of Incorrect concept problems, but its Incorrect fact rate was significantly higher.
+We observed that during security code reviews, GPT-4 tended to provide the range of code lines where defects were located, while DeepSeek-R1 often identified specific code lines and their corresponding code, making it more prone to errors in details.
+This may contribute to the higher Incorrect fact rate of DeepSeek-R1 compared to GPT-4.
+When applying LLMs to real-world security code reviews, we expect models to provide both specific and accurate factual information.
+Therefore, how to enhance the capability of LLMs to retain code details with precision requires further exploration.
 Example 1 - Incorrect concept
@@
 **Issue:** The method _inventory_has_changed uses a comparison logic that might not be entirely accurate in all scenarios...
-Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%).
-Among these, Vague statement is the most prevalent one.
+Understandability: Problems that reduce understandability of responses include Vague statement (GPT-4: 21.38%, DeepSeek-R1: 3.58%), Incoherent expression (GPT-4: 1.19%, DeepSeek-R1: 0.53%), Logical contradiction (GPT-4: 1.11%, DeepSeek-R1: 1.13%), and Groundless speculation (GPT-4: 5.62%, DeepSeek-R1: 4.11%).
+For GPT-4, Vague statement is the most prevalent one.
 It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing an overly general fix suggestion.
 An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
+While for DeepSeek-R1, the frequency of Vague statement was considerably lower than that of GPT-4.
+We speculated that its reasoning procedure under the ‘think’ mode somewhat contributed to generating specific and detailed responses.
 Example 2 - Vague statement
@@
 **General Concern**: Throughout the code, there are instances where error conditions might not be handled comprehensively...
-Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses.
-Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4.
+Conciseness: In all responses, some of them contain Redundant content (GPT-4: 1.14%, DeepSeek-R1: 39.76%), some include information Irrelevant to the task of security defect detection (GPT-4: 15.12%, DeepSeek-R1: 6.03%), and some contain Unnecessary elements (GPT-4: 86.38%, DeepSeek-R1: 63.29%), all of which are problems related to the conciseness of the response.
+We found that DeepSeek-R1 was more likely than GPT-4 to repeatedly mention the same information in its output, whereas GPT-4 produced content that is unrelated to the target task more frequently.
+Including Unnecessary content, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, is the most prevalent problem for both GPT-4 and DeepSeek-R1.
 Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes.
@@
 - There doesn’t seem to be any direct issue related to improper access control in the given code...
-Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts.
-Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform a security check in its response or misinterpreted requirements in the prompts for other tasks.
-Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found.
-Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt.
+Compliance: Four types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts.
+Among them, Detection failure was observed in very few responses (GPT-4: 0.15%, DeepSeek-R1: 0.13%), where LLM explicitly claimed that it could not perform a security check in its response or was stuck in an infinite output loop and failed to complete the security defect detection task.
+Task misunderstanding, which is similarly rare (GPT-4: 0.00%, DeepSeek-R1: 0.66%), indicates that LLM misinterprets the task details in the prompt.
+For instance, when using DeepSeek-R1 with $\textbf{P}_{cot-guardrail}$, if the commit message mentioned security defects addressed by the current commit, LLM may mistakenly interpret the task of security code review as explaining the security defects mentioned in the commit message.
+Non-compliant output was present in 20.43% of GPT-4 responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found.
+Incomplete answer is the most common problem (20.98%) under this theme for GPT-4, referring to cases in which the responses lack the line numbers or fix suggestions of detected security defects requested in the prompt.
+Although GPT-4 showed relatively high frequencies for these two problems, DeepSeek-R1 exhibited much lower rates for them, at 1.86% and 1.99%, respectively.
+Given this, we consider DeepSeek-R1 to exhibit stronger instruction-following capabilities than GPT-4 in security code review.
 The example below (Example 4) shows that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given.
@@
 **Solution:** Implement thorough checks and error handling...
-While Kabir et al.
-(2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours.
-Specifically, Kabir et al.
-(2024) requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues.
-In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%.
-This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific application areas.
+While kabir2023answers also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours.
+Specifically, kabir2023answers requested ChatGPT to answer programming questions on Stack Overflow and found that ChatGPT makes fewer factual errors compared to conceptual errors.
+This aligns with our findings for GPT-4.
+However, unlike their study, which identified correctness problems in over half (52%) of the responses, our study applied GPT-4 to security code review, revealing a much lower incidence of correctness issues at 14.81%.
+When applying DeepSeek-R1, the frequency of correctness problems increased to 33.66%.
+This suggests that the distribution of quality problems in LLM responses may vary significantly across different application domains and model architectures, necessitating distinct optimization directions and strategies in practical application.
 3.2.2 RQ2.2
@@
 Since a single response may exhibit multiple types of hallucination simultaneously, the sum of the proportions for each subtype exceeds the proportion of responses containing at least one hallucination.
-Overall, 20.19 % of the responses contain hallucinations.
-Among these, Factual Contradiction is the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency and Factual Fabrication, which appear in 6.25% and 5.62% of responses, respectively.
-To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (Tonmoy et al., 2024; Lewis et al., 2020), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (Kumar et al., 2025).
+Overall, 20.19% of GPT-4 responses contained hallucinations, whereas DeepSeek-R1 exhibited a higher rate of 36.45%.
+The former encountered more Factuality Hallucination, while the latter was predominantly affected by Faithfulness Hallucination.
+Specifically, for GPT-4, Factual Contradiction was the most frequent subtype, accounting for 10.13%, followed by Context Inconsistency (6.25%) and Factual Fabrication (5.62%).
+In contrast, for DeepSeek-R1, Context Inconsistency was the most prevalent subtype (31.81%), with Factual Fabrication and Factual Contradiction following at 4.11% and 2.39%, respectively.
+To alleviate these hallucinations, Retrieval-Augmented Generation (RAG) can be employed to incorporate external factual information (tonmoy2024comprehensive; lewis2020retrieval), and post-training can also be applied to enhance the model’s capability to capture and retain contextual knowledge (kumar2025llm).
 Table 7: Mapping of hallucination and quality problem types with corresponding proportions.
 Halluc. Type* Subtype QP Type* Explanation Prop.*
+GPT-4 DS-R1
 Factuality Hallucination Factual Contradiction Incorrect Concept In Incorrect Concept cases, the model’s output contradicts the authoritative definition of the corresponding CWE, thus categorized as Factual Contradiction.
-10.13%
+10.13% 2.39%
 Factual Fabrication Groundless Speculation Groundless speculation refers to subjective assumptions made without contextual evidence or knowledge support.
 As it involves fabricated factual content, it is categorized as Factual Fabrication.
-5.62%
-Faithfulness Hallucination Instruction Inconsistency Detection Failure Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction.
+5.62% 4.11%
+Faithfulness Hallucination Instruction Inconsistency Detection Failure, Task Misunderstanding Detection Failure refers to cases where the model fails to perform the security checks or misinterprets the task, leading to deviations from the instruction.
 Therefore, it is categorized as Instruction Inconsistency.
-0.15%
+0.15% 0.80%
 Context Inconsistency Incorrect Fact Incorrect Fact denotes cases where the line numbers, identifiers, or code content generated by the model do not align with the provided code context, thus categorized as Context Inconsistency.
-6.25%
+6.25% 31.81%
 Logical Inconsistency Logical Contradiction Logical Contradiction refers to cases where the model’s reasoning about a security defect is internally contradictory, thus classified as Logical Inconsistency.
-1.11%
-Any Hallucination 20.19%
+1.11% 1.13%
+Any Hallucination 20.19% 36.45%
     Abbreviations: Halluc. = Hallucination, QP = Quality Problem, Prop. = Proportion.
@@
 3.3.1 Model Stability
-To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991).
-This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022).
-A possible explanation is that although we have considered the non-determinism of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset.
-Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model.
-Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
+[…]
 Table 8: Explanatory powers of factors to the performance of GPT-4 under ${P}_{cid}$ in security code review.
-Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(>Chi)
-Token 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
-Community Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
-FileType Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
-AnnotationisSecurityRelated -1:0 0.97 2 9.95 0.0069**
-1:0 2.69
-AuthorExperience 0.87 1 4.89 0.0271*
-Commit 1.01 1 3.06 0.0801
-SecurityDefectType Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
-Crash:Race Condition 0.86
-Deadlock:Race Condition 0.92
-DoS:Race Condition 0.87
-Encryption:Race Condition 1.18
-Improper Access:Race Condition 0.68
-Integer Overflow:Race Condition 0.86
-Resource Leak:Race Condition 1.41
-Use After Free:Race Condition 0.37
-Extra:Race Condition 1.43
-Complexity 1.03 1 1.74 0.1868
-AnnotationRatio 0.56 1 1.55 0.2126
-AnnotationHasCode 1.11 1 1.09 0.2973
+Factor (:Reference) GPT-4 DeepSeek-R1
+OR D.F. $\chi^{2}$ Pr(>Chi) OR D.F. $\chi^{2}$ Pr(>Chi)
+Token 0.63 1 54.20 $0001^{***}$ 0.44 3 9.81 0.0017**
+Position 0.89 1 0.2685 0.6042 1.15 1 0.47 0.4909
+Community Qt:Openstack 1.54 1 8.14 0.0043** 1.41 1 6.03 0.0141*
+FileType Source:Auxiliary 2.81 1 19.04 $0001^{***}$ 0.57 1 5.57 0.0183*
+AnnotationisSecurityRelated -1:0 1.04 2 15.46 0.0004*** 1.28 2 14.02 0.0009***
+1:0 2.91 2.57
+AuthorExperience 0.93 1 1.29 0.2569 0.96 1 0.65 0.4203
+Commit 1.01 1 4.82 0.0281* 1.00 1 0.01 0.7379
+SecurityDefectType Buffer Overflow:Race Condition 0.84 10 11.29 0.3353 1.25 10 28.12 0.0017**
+Crash:Race Condition 0.94 1.01
+Deadlock:Race Condition 0.83 0.93
+DoS:Race Condition 0.98 1.99
+Encryption:Race Condition 0.83 4.43
+Improper Access:Race Condition 0.80 2.16
+Integer Overflow:Race Condition 0.85 1.16
+Resource Leak:Race Condition 1.49 1.14
+Use After Free:Race Condition 0.41 0.84
+Extra:Race Condition 0.74 1.99
+Complexity 1.03 1 2.60 0.1071 1.08 1 15.51 $0001^{***}$
+AnnotationRatio 0.60 1 1.30 0.2538 1.16 1 0.11 0.7379
+AnnotationHasCode 0.83 1 1.03 0.3094 1.46 1 4.37 0.0364*
   • Statistical significance of explanatory power according to Wald $\chi^{2}$ test:
@@
 We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact.
-As shown in Table 8, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (<.0001).
-Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$<.0001) and ($\chi^{2}$=18.28, $p$<.0001).
-Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively.
-However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold.
-Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom.
+As shown in Table 8, for GPT-4, ‘Token’ wielded the highest predictive power on the fitted model, with the highest $\chi^{2}$ (54.20) and the lowest p-value (<.0001).
+Trailing closely behind were ‘FileType’ and ‘AnnotationisSecurityRelated’ factors, with correspondingly significant contributions ($\chi^{2}$=19.04, pp<.0001) and ($\chi^{2}$=15.46, pp=0.0004).
+Adding to this, ‘Community’ and ‘Commit’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 8.14 and 4.82, respectively.
+For DeepSeek-R1, ‘Complexity’ and ‘AnnotationisSecurityRelated’ demonstrated the strongest explanatory power over the response variable, with $\chi^{2}$ of 15.51 and 14.02, and p-values both less than 0.001.
+Following closely were ‘SecurityDefectType’ and ‘Token’, with p-values of 0.0017 (less than 0.01), indicating that they also have a statistically significant impact on the model’s performance in security code review.
+Although the $\chi^{2}$ for ‘Community’ (6.03), ‘FileType’ (5.57) and ‘AnnotationHasCode’ (4.37) were relatively smaller, they still contributed predictive value in the fitted model.
+Factors not analyzed here were deemed not to largely influence the performance of GPT-4 and DeepSeek-R1 in detecting security defects in code reviews, as their p-values exceeding the 0.05 threshold and $\chi^{2}$ values were much lower than those discussed factors.
+Notably, ‘SecurityDefectType’ had a relatively high $\chi^{2}$ value (11.29 for GPT-4 and 28.12 for DeepSeek-R1), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom.
 As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
 3.3.3 Relationships between Factors and Responses
-The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016).
+The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (paul2021security; thongtanunam2017review; setia2016methodology).
 In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response.
@@
 For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable.
-We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 8).
-The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively.
-This suggests that the LLM is more likely to detect security defects in code files with fewer tokens, that contain functional logic, and were written by developers with less involvement in the project.
-Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt.
-In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect.
-Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97.
-This suggests that annotations may disrupt the LLM and are less likely to detect the target security defect.
+We utilized the summary function of the rms package (harrell2017package) to calculate the OR for each factor (see Table 8).
+As presented in Table  8, for both DeepSeek-R1 and GPT-4, the OR for the ‘token’ factor was less than 1, indicating that the models are more likely to detect security defects in code files with fewer tokens.
+The OR values for ‘AnnotationisSecurityRelated’ and ‘community’ were greater than 1, suggesting that the two models are more likely to detect security defects in code files that include security-related annotations and originate from the Qt community, rather than from OpenStack.
+However, regarding ‘FileType’, GPT-4 and DeepSeek-R1 showed opposing results.
+GPT-4 had an OR of 2.81 (>1>1), indicating a higher likelihood of detecting security defects in source code with functional logic, while DeepSeek-R1 had an OR of 0.57 (<1<1), suggesting better performance on header files utilized for declaration.
+Furthermore, DeepSeek-R1 exhibited varying detection capabilities for different types of security defects.
+According to the OR values of ‘SecurityDefectType’, DeepSeek-R1 demonstrated a greater tendency to detect various other security defect types compared to Race Condition, including but not limited to Encryption, Buffer overflow, Integer Overflow, and Resource leak.
+In addition, DeepSeek-R1 displayed stronger capabilities in security defect detection for files with more code complexity, as evidenced by its OR for ‘complexity’ being greater than 1 (1.08).
+This contrasts with the consensus in previous works that as complexity of software grows, security defects are becoming more challenging to detect sheng2024lprotector.
+A possible explanation is that we utilized McCabe’s Cyclomatic Complexity mccabe1976complexity to measure this factor, which is not always positively correlated with the actual difficulty of detection.
+Code with higher cyclomatic complexity may have a clearer code structure and retain more contextual information within the single code file, which could better support DeepSeek-R1’s reasoning procedure during its think mode.
+On the other hand, the analysis of security defects in code with lower cyclomatic complexity may require more cross-file code context and project-specific knowledge, leading to DeepSeek-R1 struggles to determine whether security defects exist.
+It is also worth noting that while previous studies have explored the relationship between the position of security defects and the performance of LLMs, our findings differ.
+According to the study by Sovrano_2025, the ‘position’ of security defect was determined by counting the characters preceding the security defect and they concluded that security defects located at the end of the code file are often harder to detect.
+However, when the file itself contains a large number of characters, even if the defect is located in the relatively middle or beginning part of the file, its ‘position’ can still be quite high.
+Therefore, the logistic regression analysis in Sovrano_2025 was to some extent confounded by the effect of character count (or token count).
+Our study, by measuring ‘position’ as the ratio of tokens preceding the security defect relative to the total token count, reduced the correlation between ‘Position’ and ‘Token’.
+It can be seen from Table 8 that in both of the two fitted models, the p-values of ‘Position’ were far greater than 0.05, indicating that the position of security defect has no significant effect on the performance of either GPT-4 or DeepSeek-R1.
+While the OR of ‘Position’ (0.89) in GPT-4 was less than 1, consistent with  Sovrano_2025, DeepSeek-R1 showed an OR of 1.15, which is greater than 1.
+Therefore, the impact of the position of security defect in code files on the detection performance of latest reasoning models warrants further analysis and validation.
 4 Discussion
@@
 However, LLMs can serve as auxiliary tools to assist reviewers.
-The findings of RQ3 suggest that LLMs are more adept at identifying relatively simple security defects in shorter, logic-centric code written by developers with less project experience.
-We recommend a multi-layered review strategy to fully leverage the capability of LLMs - reviewers can first use LLMs to conduct an initial automatic review of those short and simple commits for security defects, then review both the submitted code and the LLM’s detection results to improve efficiency.
-This would allow reviewers to concentrate on an in-depth review of those complex commits, including lengthy code written by more experienced developers, thereby reducing the risk of missing sophisticated security defects.
-Additionally, where LLMs currently show limitations, these complex commits should be more heavily incorporated into datasets used for fine-tuning LLMs to improve their effectiveness in security code review.
-Providing a task scope is crucial for enhancing the performance of LLMs.
-Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al.
-(2024).
-Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy.
+We recommend a multi-layered review strategy to fully leverage the capability of LLMs.
+The findings of RQ3 suggest that LLMs are more adept at identifying security defects in code file with fewer tokens and higher cyclomatic complexity.
+Utilizing these two factors, a lightweight model can be trained to classify the code file to be reviewed into the low-difficulty or high-difficulty groups.
+Firstly, LLMs are utilized to conduct an initial automatic review of these files.
+Then, for low-difficulty code, reviewers quickly confirm the security defects identified by LLMs.
+Code in which LLMs did not detect any security defects would then be resampled and rechecked to confirm its security.
+For high-difficulty code, reviewers utilized the detection results of LLMs as a reference, combined with their own domain knowledge, to perform a more thorough and in-depth analysis of the submitted code.
+This approach aims to improve the efficiency of security code review while minimizing the risk of missing security defects.
+Additionally, reviewers should provide feedback on each of the detection results generated by LLMs, which can then be used to train both the difficulty classification model and LLMs for security code review.
+When using a CWE list as auxiliary information for LLM in security code review, the granularity of CWE entries is crucial.
+Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of steenhoek2024comprehensive.
 However, in real-world security reviews, although with experience and domain knowledge, the reviewers lack prior knowledge of all potential types of security defects that may be present in the code.
-Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary.
 It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses.
-In Bakhshandeh et al.
-(2023), providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs.
+In bakhshandeh2023using, providing a list of all 75 CWEs in their prompts decreased the accuracy of security defect detection and location by the LLMs.
 This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects.
-Due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible.
+Therefore, constructing a concise yet comprehensive CWE list is necessary.
+In our experiments, due to the limited context length of the used LLMs, inputting the entire codebase during prompt design was not feasible.
 The lack of code context made it difficult for the LLM to fully understand the dependencies and interaction logic between files, which in turn affected its performance in identifying security defects, such as Race Condition and Insecure Library Version.
@@
 Ultimately, we identified and selected four prompts that outperformed others, which, together with the basic prompt, were selected as templates for our formal experiments.
-The results demonstrate that including a list of CWEs significantly enhances LLM performance, making it our study’s most effective prompting strategy.
-As the context window of LLMs continues to expand and prompt engineering techniques evolve, we plan to explore additional strategies in the future, such as incorporating RAG for external information and employing automated prompt optimization to dynamically adjust prompts during reasoning, thereby further improving the capability of LLMs in security defect detection.
-Using LLMs in security code review requires developers to standardize their code commenting practices.
-According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review.
-This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects.
-Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative.
-The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers.
-Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better supporting LLMs in code review tasks.
+The results in RQ1.3 demonstrate that for GPT-4, including a CWE list in the prompt significantly enhances LLM performance, while for DeepSeek-R1, the impact of the provided CWE list on model performance depends on the specific programming language.
+We therefore emphasize the need to explore CWE entries of different granularities for various types of security defects to determine the optimal CWE list.
+As the context window of LLMs continues to expand and prompt engineering techniques evolve, we also plan to explore additional strategies in the future, such as incorporating RAG for external information, employing automated prompt optimization to dynamically adjust prompts, and leveraging the agent architecture to dynamically invoke contextual information during reasoning, thereby further improving the capability of LLMs in security defect detection.
 With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs.
@@
 This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets.
-Additionally, according to RQ1, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2.
-Therefore, ensuring the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
+Additionally, according to RQ1, for general-purpose LLMs, despite GPT-4 Turbo having a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2.
+While for the reasoning-optimized LLM, although DeepSeek-R1 also has a 128k context window and is the best-performing model in this study, its rate of Incorrect fact reaches as high as 31.81%.
+Specifically, DeepSeek-R1 often presented problems with mismatched code line numbers and code snippets in its responses, which significantly undermines the understandability of its analysis.
+Therefore, while expanding the LLMs’ context window, it is also necessary to guarantee the model’s precise memory of detailed information such as key code segments, to enhance the accuracy of security defect detection and the clarity of defect descriptions.
 A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results.
-The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses.
-This may be because the outputs generated by LLMs with higher non-determinism are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower non-determinism.
-Both detection capability and the stability of detection results are crucial for applying LLMs in security code review.
+The results from RQ1.4 indicate that the two LLMs (DeepSeek-R1 and GPT-4) with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses.
+This may be because LLMs with better performance are more capable of reasoning from multiple perspectives.
+The inherent randomness of LLM leads to variations in reasoning paths, resulting in inconsistent outcomes.
+In comparison, the thought of less capable models is relatively shallow, and their strong consistency in responses may stem from failing to deeply analyze the code.
+The security code review task requires models to explore as many reasoning paths as possible in order to maximize the recall of security defects, making the inconsistency of responses inevitable.
+In real-world application of LLMs for security code review, both the accuracy and stability of detection results are crucial.
 However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results.
-Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the non-determinism inherent in LLMs represents a valuable direction for future improvements.
+Therefore, for researchers, exploring approaches to improve the detection sensitivity while maintaining result stability, such as through more precise contextual information or voting mechanisms, is a valuable direction for future improvement.
+Reasoning models exhibit limited sensitivity to logic-related security defects during security code review The results from RQ3 show that DeepSeek-R1 underperformed in detecting race conditions compared to most other security defect types.
+Previous studies have also demonstrated that the detection accuracy of LLMs suffers from an imbalance across different security defect types, with race conditions being relatively underexplored sheng2025llms.
+Given that Race Condition is a high-frequency type of security defect in Python projects (accounting for 65.1% on the Python dataset in this study, see fig.
+3), it is crucial to guide LLMs to pay more attention to logic-related security defects.
+Based on the results of RQ1, $\textbf{P}_{cid}$ and $\textbf{P}_{cot-guardrail}$ enhanced DeepSeek-R1’s capability of detecting security defect detection.
+To separately analyze their effect on race conditions, we further calculated the IH-Scores of DeepSeek-R1 on code files with race conditions from the complete dataset (Python and C/C++) under $\textbf{P}_{b}$, $\textbf{P}_{cid}$, and $\textbf{P}_{cot-guardrail}$, which were 5.09%, 7.56%, and 6.02%, respectively.
+This suggests two potentially effective approaches to improve the sensitivity of LLMs on logical-related security defects: first, explicitly guiding the model to concentrate on the corresponding areas through CWE entries; and second, leveraging the commit message to assist LLMs in better understanding code functionality and logic implementation.
 5 Threats to Validity
@@
 The data used in our study was obtained from two large-scale open-source projects, which may overlap with the data used for training LLMs, a common problem with LLM-based studies.
-Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time.
-Moreover, according to Cao et al.
-(2024), contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date.
+Since the training data cutoff dates for the models used (GPT-3.5 and GPT-4) are not disclosed and the dataset we used  (yu2023security) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time.
+Moreover, according to cao2024concerned, contaminated data does not always affect the results, and the model sometimes even performs better after the model’s cutoff date.
 Another threat is the design of prompts.
@@
 In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information.
-To specifically analyze its impact, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to generate code context according to the commit message before detecting security defects.
+To explore approaches for context augmentation, we designed a prompt $\textbf{P}_{cot}$, in which LLMs are instructed to reason step-by-step—generating code context according to the commit message before detecting security defects.
+Since multiple factors in $\textbf{P}_{cot}$ could influence model performance, we also designed $\textbf{P}_{cot-guardrail}$ as a control prompt that includes only the commit message and CoT instruction, in order to isolate the effect of the fabricated code context.
 External Validity concerns whether a causal link generalizes across contexts.
-In this study, the experiments were conducted on 6 LLMs based on 4 open-source projects across programming languages including Python and C/C++.
+In this study, the experiments were conducted on 7 LLMs based on 4 open-source projects across programming languages including Python and C/C++.
 We acknowledge that the dataset we utilized may not fully represent industrial projects across various programming languages, and the performance of the selected LLMs may not represent the capabilities of all existing LLMs.
 We implemented the following strategies to ensure the generalizability of the experimental results.
-For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular LLMs.
+For model diversity, our experiments incorporated both open-source and proprietary LLMs of varying parameter scales to encompass a diverse range of popular general-purpose and reasoning-optimized LLMs.
 For dataset construction, we collected code review data from real-world open-source projects in the OpenStack and Qt communities.
@@
 In addition, the data labeling in RQ1 and the analysis of quality problems in RQ2 were all carried out manually.
-In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (Glaser, 1965).
+In order to further reduce the impact of subjective judgments, we employed open coding and constant comparison (glaser1965constant).
 In the data labeling employed for RQ1, the first and third authors conducted a pilot data labeling and completed all the labeling work after reaching a consensus on the criteria through discussion.
@@
 Several studies have analyzed security defect detection during code reviews.
-Paul et al.
-(2021) built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects.
+paul2021security built a dataset from the Chromium OS project, consisting of 516 code reviews that successfully identified security defects and 374 code reviews that missed security defects.
 The study analyzed 18 code review attributes, including review duration, reviewer experience, and author experience, to explore factors influencing the identification of security defects.
 It was found that the duration of code reviews and the frequency of peer reviews positively affected the detection of security defects, while the number of directories under review and the number of prior commits had a negative impact.
-Alfadel et al.
-(2023) examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews.
+alfadel2023empirical examined pull requests (PR) from 10 popular JavaScript GitHub projects within the npm ecosystem, then analyzed the frequency, categories, and handling of security defects identified during code reviews.
 Their work indicated that only a small fraction of PRs raise security problems, and most of the identified security problems were resolved during the code review.
@@
 Many studies have focused on analyzing code for vulnerabilities and security defects in hopes of helping developers understand how security problems originate in a system, so that defects are detected early.
-Alves et al.
-(2016) constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others.
+alves2016software constructed a security defect dataset based on five widely used projects and 27 software metrics, such as lines of code, cyclomatic complexity and knots in the control flow graph, among others.
 This study demonstrated that most of these metrics are effective in distinguishing vulnerable from non-vulnerable code, thereby providing valuable support for security practitioners.
-Meneely et al.
-(2013) conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects.
+meneely2013patch conducted quantitative and qualitative analyses on 124 Vulnerability-Contributing Commits (VCC) in the Apache HTTP Server to identify characteristics of code changes with security defects.
 The findings suggest that longer lines of code and less experienced committers are more likely to introduce security defects.
-Bosu et al.
-(2014) identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with Meneely et al.
-(2013) that longer lines of code and less experienced committers are more likely to introduce security defects.
+bosu2014identifying identified 413 VCCs from code reviews in 10 open-source projects and reported consistent observations with meneely2013patch that longer lines of code and less experienced committers are more likely to introduce security defects.
 Additionally, they indicated that modified code files are more likely to contain security defects than new files.
@@
 Since our study focuses on using LLMs to assist reviewers in identifying security defects during code review, we limit our coverage to studies related to security defect detection, excluding those studies on security defect remediation or program repair.
-Charoenwet et al.
-(2024) conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs.
+charoenwet2024empirical conducted an empirical study on the application of static application security testing tools (SASTs) in security code review and found that a single SAST tool could produce warnings for 52% of VCCs.
 By combining several tools, warnings could be produced for 78% of the VCCs.
@@
 The majority of these studies conducted a coarse-grained assessment of LLMs in binary classification tasks that determine whether the code contains any security defects without specifying the corresponding defect type.
-Zhou et al.
-(2024b) utilized Common Weakness Enumeration (CWE), project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts.
+zhou2024largelanguage utilized CWE, project information, and a few-shot learning approach to design prompts and evaluated the capability of two popular LLMs, GPT-3.5 and GPT-4, in binary judgment under various prompts.
 They found that GPT-3.5 achieved competitive performance with the prior state-of-the-art detection approaches, while GPT-4 consistently outperformed the state-of-the-art.
-Purba et al.
-(2023) compared the performance of LLMs with popular static analysis tools in binary detection.
+purba2023software compared the performance of LLMs with popular static analysis tools in binary detection.
 Their work demonstrates that although LLMs achieved a good recall rate, their false positive rate was significantly higher than that of static analysis tools.
+jensen2024software evaluated the performance of three proprietary OpenAI LLMs and four small-scale models with fewer than 13B parameters to identify vulnerabilities in the SecurityEval dataset (siddiq2022securityeval).
+They adopted zero-shot and CoT prompting and found that the best‑performing model - GPT‑4 - achieved an accuracy of 95.6% in distinguishing between vulnerable and patched versions of code.
+Similarly,  steenhoek2024err assessed the capabilities of 14 LLMs on a real-world and function-level vulnerability dataset, SVEN, across seven different prompt strategies.
+They explored various approaches, including prompt optimization, scaling model size, increasing pre-training data, fine-tuning, and incorporating domain knowledge into prompts, yet none led to significant improvements in the performance of models in the binary vulnerability detection task.
 Unlike previous studies, our work provides a more fine-grained evaluation of the capability of LLMs in security code reviews by including additional insights such as defect localization, description, and suggestions for defect remediation.
 Previous studies also investigated whether LLMs could provide more information in security defect detection.
-Bakhshandeh et al.
-(2023) provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers.
+bakhshandeh2023using provided LLMs with a list of CWEs and the inspection results of static analysis tools on the prompts, requesting the LLMs to generate specific CWE names along with the corresponding code line numbers.
 Their study shows that LLMs effectively reduce false positives of static tools and demonstrate competitive or even superior performance in defect localization compared to these tools.
-Yin and Ni (2024) evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (Fu and Tantithamthavorn, 2022) and SVulD (Ni et al., 2023), generally outperformed LLMs.
+yin2024multitask evaluated the performance of LLMs across multiple tasks in security analysis, including defect detection, localization, and description, and found that the existing state-of-the-art approaches, such as LineVul (fu2022linevul) and SVulD (ni2023distinguishing), generally outperformed LLMs.
 Specifically, LLMs exhibited limited overall accuracy and varied performance across different CWE types in defect localization and description tasks.
+khare2025understanding comprehensively evaluated 16 pre-trained LLMs for detecting security defects on a function-level vulnerability dataset of 5,000 samples.
+Their study systematically compared model performance across three distinct prompting strategies and found that involving step-by-step analysis significantly improved LLM performance.
+This work also performed comparison of LLMs against leading static analysis tools and deep learning-based approaches, finding that LLMs reported higher detection accuracies than them.
 These studies have all been limited to quantitatively evaluating LLM performance.
 In comparison, our work performed an in-depth analysis of the quality problems in LLM-generated responses and factors affecting LLM performance to understand the potential and challenges LLMs face in security code review.
-Employing specific prompting frameworks to enhance the performance of LLMs in identifying security defects has also been explored.
-Yang et al.
-(2024) introduced a framework utilizing DL models, integrating context learning and Chain-of-Thought (CoT) to augment prompts in security defect detection.
+Fine-tuning has also been employed to improve the capability of LLMs in security defect detection.
+mechri2025secureqwen fine-tuned CodeQwen1.5-7B on a large Python dataset comprising both real-world code and synthetic code generated by LLMs, in order to guide the model to capture and classify security defects in code accurately.
+The model ultimately achieved high accuracy, with F1 scores ranging from 84% to 99%.
+Although fine-tuning can effectively improve the accuracy of model, fine-tuned LLMs often exhibit poor generalization.
+In practical application, adjustments to the model parameters and subsequent retraining are usually required.
+Consequently, some studies have shifted their focus toward approaches that avoid altering the core model, such as leveraging prompt engineering or incorporating external information to further enhance the precision of LLMs in security defect detection.
+yang2024dlap introduced a framework utilizing DL models, integrating context learning and CoT to augment prompts in security defect detection.
 Their framework yielded superior performance compared to state-of-the-art prompting techniques.
-Nong et al.
-(2024a) proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution).
-The research concerns and employed datasets of the above two studies (Yang et al., 2024; Nong et al., 2024a) differ significantly from ours.
+nong2024chain proposed a vulnerability-semantics-guided prompting (VSP) approach that combines CoT with various auxiliary information to evaluate the performance of LLMs in identifying security defects (including their classification and resolution).
+The research concerns and employed datasets of the above two studies (yang2024dlap; nong2024chain) differ significantly from ours.
 These studies focus on designing benchmarks to enhance LLM in vulnerability detection, utilizing synthetic datasets and CVE datasets composed of publicly disclosed vulnerabilities to validate their improvements.
 Compared with them, our work focuses on the empirical evaluation of LLMs in the context of security code review, employing a dataset of security defects collected from code reviews of real-world software projects.
+To further improve the accuracy of LLMs in security defect detection, some studies have adopted RAG to retrieve external knowledge and enhance the input prompts.
+du2024vul constructed multi-dimensional representations to retrieve knowledge most relevant to the target code from historical vulnerabilities and their corresponding fixes.
+This framework improved the accuracy of LLMs in distinguishing vulnerable code from patched code by 16%–24%.
+lu2024grace incorporated graph-structured information from the code and identified highly related code samples as external guidance into the prompt.
+Using this method, LLMs outperformed state-of-the-art deep learning-based tools in detecting vulnerabilities.
+With the emergency of reasoning-optimized LLMs, LLM-based agents have also shown promising applications in security defect detection.
+Inspired by courtroom, widyasari2025let constructed a framework consisting of four distinct agent roles, which are security researcher, code author, moderator, and review board.
+Their framework employs multiple rounds of interaction to arrive at a final determination on security defects in the code, which substantially surpassed the performance of approaches based on CoT and fine-tuning.
+tsai2025leveraging incorporated RAG into the multi-turn multi-role interaction agent framework, achieving a 29.9% improvement in F1 score compared to widyasari2025let.
 It is worth noting that the studies mentioned above focus on general vulnerability detection by LLMs, whereas our work targets the specific area of security code review by LLMs.
@@
 In our study, we comprehensively analyzed the practicality of current popular LLMs in the context of security code review.
-We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews.
-For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM.
-We performed a regression analysis to explore factors influencing its performance.
-Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to state-of-the-art static analysis tools.
-However, their application remains limited.
-(2) GPT-4 performs best among all LLMs when provided with a CWE list for reference.
-(3) During code review, GPT-4 is more effective in identifying security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project.
-(4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.
-We plan to extend this work in the next step by incorporating code reviews from proprietary software projects into our dataset and attempting enhancement strategies such as few-shot learning and fine-tuning.
-In addition to applying LLMs to help reviewers detect security defects during code review, we also plan to analyze the ability of LLMs to fix security defects in code so that developers can comprehensively evaluate the potential of LLMs in the task of secure code review.
+We investigated the capabilities of 7 LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews.
+For the two top-performing LLMs, each paired with their respective optimal prompts, we conducted a linguistic analysis of the quality problems in their generated responses.
+Furthermore, we performed separate regression analyses to explore the factors influencing the performance of each LLM-prompt combination.
+Our main findings are: (1) In security code review tasks, LLMs demonstrate stronger capabilities than state-of-the-art static analysis tools.
+Among these LLMs, the reasoning-optimized model outperforms established general-purpose models.
+However, their practical application remains limited.
+(2) DeepSeek-R1 is the top-performing LLM in security code review, followed by GPT-4.
+DeepSeek-R1 performs best when the prompt includes the commit message and CoT guidance, whereas GPT-4 performs best when provided with a CWE list in the prompt.
+(3) DeepSeek-R1 and GPT-4 are more effective at detecting security defects in files with fewer tokens and with security-related code annotations.
+Notably, DeepSeek-R1 was found to be less focused on Race Condition than other types of security defect.
+(4) The most prevalent quality problem for both DeepSeek-R1 and GPT-4 is verbosity.
+In comparison, GPT-4 tends to produce vague expressions and demonstrates poorer instruction-following ability, while DeepSeek-R1 is more prone to errors in recalling code details.
+In future work, we plan to explore several directions to further enhance the capabilities of LLMs in security code review: (1) Improving the sensitivity to security defects of LLMs, by investigating the optimal granularity of the CWE list provided in the prompt and incorporating external knowledge or project-specific context to ensure comprehensive consideration of frequent security defect types; (2) Enhancing contextual reasoning by developing approaches to provide more precise code context to inform the analysis procedure; (3) Increasing the stability of the LLMs’ outputs via techniques such as self-refinement and voting mechanisms, to generate more consistent and reliable results.
 Data availability
-We have shared the link to our dataset in the reference (Yu et al., 2025).
+We have shared the link to our dataset in the reference (replpack).
 Acknowledgments
-This work is supported by the National Natural Science Foundation of China under Grant No. 62172311 and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.
+This work is supported by the National Natural Science Foundation of China under Grant No.
+62172311 and the Major Science and Technology Project of Hubei Province under Grant No.
+2024BAA008.
+The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
 […]